### v1

# Robust Survival Analysis with Adversarial Regularization

Abstract
Survival Analysis (SA) is about modeling the time for an event of interest to occur, which has important applications in many fields, including medicine, defense, finance, and aerospace. Recent work has demonstrated the benefits of using Neural Networks (NNs) to capture complicated relationships in SA. However, the datasets used to train these models are often subject to uncertainty (e.g., noisy measurements, human error), which we show can substantially degrade the performance of existing techniques. To address this issue, this work leverages recent advances in NN verification to provide new algorithms for generating fully-parametric survival models that are robust to such uncertainties. In particular, we introduce a robust loss function for training the models and use CROWN-IBP regularization to address the computational challenges with solving the resulting Min-Max problem. To evaluate the proposed approach, we apply relevant perturbations to publicly available datasets in the SurvSet repository and compare survival models against several baselines. We empirically show that Survival Analysis with Adversarial Regularization (SAWAR) method on average ranks best for dataset perturbations of varying magnitudes on metrics such as Negative Log Likelihood (NegLL), Integrated Brier Score (IBS), and Concordance Index (CI), concluding that adversarial regularization enhances performance in SA.
Code: https://github.com/mlpotter/SAWAR
Index Terms:
Survival Analysis, Adversarial Robustness, Neural Networks, Calibration, CROWN-IBP
publicationid: pubid: $0$–$00$ 2024 IEEE
I Introduction
Survival Analysis (SA) is a highly active research field [1, 2] used to model the time before an event happens. For example, SA has been used extensively in medicine to quantify the effectiveness of various treatments as part of studies for cancer patients and drug trials [3, 4, 5]. While early approaches (dating back to the 1950s) predominantly used linear parametric models [6, 7], recent methods often leverage Neural Networks (NNs) to model complicated time-to-event data [8].
A key open challenge in this field is in developing models that are robust to uncertainty and variability in the datasets. Successful generalization of machine learning models over unseen datasets is crucial for clinical applications, as efficacy of developed drugs or treatments should be demonstrated over diverse patient demographics, institutions, and data acquisition settings [9, 10, 11]. Data uncertainty could be caused by human errors (e.g., notoriously poor handwriting of physicians [12], ineffective medical record keeping [13]), or inherent noise during transmissions or data collection [14]. However, many approaches assume perfect measurement of the covariates, ignoring aleatoric uncertainty [15]. In SA, the aleatoric uncertainty is only associated with the noise of the log of the measured time-to-event [16, 6]. Moreover, several studies have shown that adversaries can influence the final model to be over-/under-confident by purposefully poisoning studies or otherwise disrupting the data collection activities [17, 18]. Despite these known failure modes, and coupled with the sensitivity of NNs to input uncertainties, addressing the robustness of NNs against $\epsilon$-perturbations trained specifically for fully-parametric SA remains unexplored.
To address these issues, this paper’s contributions include:
  • identification of the sensitivity of existing SA methods, such as Deep Regularized Accelerated Failure Time (DRAFT) [19], to covariate perturbations via evaluating open-source datasets with a wide range of metrics related to predictive accuracy, calibration, and population curve estimation.
  • our approach, Survival Analysis with Adversarial Regularization (SAWAR), which is the first method for enabling adversarial robustness to covariate perturbations within a bounded uncertainty set for fully-parametric NN-based SA. We improve calibration and stability on unexpected or noisy data by introducing a loss function based on a Min-Max formulation and Convex Relaxation Based Backward Bounding Pass with Interval Bound Propagation (CROWN-IBP) [20, 21].
The paper is structured as follows: Section II describes the State of the Art; Section III provides background to SA terminology and dataset structure; Section IV provides background on the adversarial and perturbation attacks used; Section V discusses the custom loss employed to deal with those attacks; Section VI provides the results of the experimental activity; finally, Section VII derives conclusions on our findings.
II Related Work
This section focuses on the literature about NNs for SA, with further detail about NN implementations available in a recent survey paper [22]. The first use of NNs in this field was a Feedforward Neural Network (FFNN) to approximate the survival function [23], defined as the probability that no event occurs up to a specific time $t$. More recently, [24] proposed Cox-nnet, a NN that outputs a proportional-hazard [25, 6], which is the assumption that the difference in effect between different instances is constant as a function of time. Building on this idea, [26] developed DeepSurv, a Cox proportional-hazard NN for personalized treatment, which uses dropout layers to improve regularization and reduce overfitting. Instead of proportional hazard estimation, DeepWeiSurv [27] regresses the parameters of a Weibull mixture distribution with NNs.
Several recent works in SA aim to improve model calibration (e.g., by using new terms in the objective function [28, 29, 30]). A calibrated model provides a predicted confidence that is indicative of the true likelihood that the model is correct [31]. Meanwhile, few works have addressed adversarial robustness in SA. In general, a model is adversarially robust if small perturbations to the input do not cause the model to make incorrect predictions [32]. Adversarial robustness and model calibration are loosely coupled: while [33] showed that better calibration does not necessarily improve generalization, [34] showed that incorporating adversarial robustness during training improves trustworthiness in calibration and stability on unexpected data.
Existing adversarial training methods for SA face important tradeoffs between model complexity, performance, and ease of embedding of physical assumptions. Most of these techniques use Conditional Generative Adversarial Networks (CGANs) to generate samples of the time-to-event from a survival distribution in a non-parametric manner [19, 35, 36]. These models (e.g., CGANs) do not assume a functional form and define the survival distribution implicitly [37, 38, 39]. Therefore, there is no mathematical expression that describes the time-to-event Probability Density Function (PDF) value given covariates. One benefit of implicit models is that increased model complexity leads to increased flexibility, which is the ability to fit many other potential data realizations [40]. However, the higher model complexity generally violates the Bayesian Occam’s Razor, which penalizes models for having too much flexibility [40]. To satisfy the Bayesian Occam’s Razor’s “just right“ model complexity, modeling assumptions known as inductive biases are employed in the form of an explicit parametric model [39, 37]. For example, a common inductive bias for SA models is the memoryless property [38], which the exponential distribution satisfies. For this reason, this paper uses the exponential distribution, which is commonly used in SA [41, 42, 43] (and more generally in queuing theory [44, 45], reliability analysis [46, 8], and telecommunications [45]).
III Preliminaries on Survival Analysis
SA [25, 6] is a statistical approach used to analyze and model the time $t\in\mathcal{R_{++}}$ until an event $e\in\{0,1\}$ happens (e.g., a disease occurs, a device breaks). Typically, the study happens over a finite time period and instances/individuals may leave the study early before an event is recorded, with no follow up. In those cases, the dataset records that no event occurred up until the exit time of the study, which is called right censoring [47]. Such an SA dataset consists of $N$ independent instances/individuals, each denoted by index $i$, for which time between entry to a study and the subsequent event is recorded. Upon entry time to a study, the instance’s covariates $\boldsymbol{x}_{i}\in\mathcal{R}^{d}$ (i.e., variables that may affect the event of interest for the study) are measured. In summary, the dataset $\mathcal{D}$ has the following structure:
$\mathcal{D}=\{(\boldsymbol{x_{i}},t_{i},e_{i})\}_{i=1}^{N}.$
Being a statistical approach, we define the required probabilistic functions to model time-to-event data next.
III-A Proportional Hazard Model
Let us denote $T$ as a continuous, non-negative, random variable that represents the time-to-event, with PDF $f(t)$ and Cumulative Distribution Function (CDF) $F(t)=P[T\leq t]$. The survival function $S(t):=1-F(t)$ is the complement of this CDF [48, 6].
An exponential parametric survival model is expressed as
$S(t)=e^{-\Lambda(t)},$ (1)
[…]
$\lambda(t)=\frac{f(t)}{S(t)}=\lim_{\Delta\rightarrow 0}\frac{P[t<T<t+\Delta]}{\Delta}$ (2)
is the instantaneous failure rate at time $t$.
To incorporate the instance’s covariates, we use the proportional hazard model [49, 7], This model assumes that the covariates increase or decrease the hazard by a proportional amount at all times [48, 6],
$\lambda_{\boldsymbol{\theta}}(t,\boldsymbol{x})=\lambda_{0}(t)e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})},$ (3)
where $\lambda_{0}(t)$ is a baseline hazard, $e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}$ is the relative risk associated with covariates $\boldsymbol{x}$, and $G_{\boldsymbol{\theta}}(\boldsymbol{x})$ is an NN with parameters $\boldsymbol{\theta}$ in our setting.
The exponential proportional hazard model characterizes the baseline hazard rate as a constant function of time, and therefore $\lambda_{0}(t)=\lambda_{0}$. For simplicity, we assume an exponential baseline hazard, and we may “absorb” the $\lambda$ term into the bias of the NN:
$\lambda_{0}e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}=e^{\log\lambda_{0}+G_{\boldsymbol{\theta}}(\boldsymbol{x})}=e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}.$ (4)
The proportional hazard model PDF, CDF, and complement CDF for an exponential proportional hazard model are then given by:
$f_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ $=e^{G_{\theta}(\boldsymbol{x})}e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$
$F_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ $=1.0-e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$
$S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ $=e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}.$
$S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ is also known as the instance survival function.
The population survival curve is the marginalization of the instance survival function over the covariates,
[…]
Since the distribution of the covariates $p(\boldsymbol{x})$ is typically unknown [50], the integral in Eq. 5 is often replaced with a Monte Carlo estimate [28]:
$S(t)$ $\approx\frac{1}{N}\sum_{i=1}^{N}S_{\boldsymbol{\theta}}(t|\boldsymbol{x_{i}}).$ (6)
III-B Baseline Objective Functions
Given those definitions, the following loss functions enable finding model parameters $\boldsymbol{\theta}$ from dataset $\mathcal{D}$. In particular, we follow [51, 52, 19] to define two objective functions based on negative log likelihood (Section III-B1) and the rank correlation (Section III-B2).
III-B1 Right Censored Log Likelihood Objective
The $\mathcal{L}_{LL}$ objective is the log of the time-to-event likelihood over $\mathcal{D}$ specified by the proportional hazard model [6]:
$\mathcal{L}_{LL}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})=$ $\sum_{i=1}^{N}e_{i}\cdot\log f_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x_{i}})\ +$
$\sum_{i=1}^{N}(1-e_{i})\cdot\log S_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x_{i}}).$ (7)
The first summation in Eq. 7 is the log likelihood of the datapoints corresponding to when the event occurs and the exact time of occurrence is observed. The second summation in Eq. 7 is the log likelihood when right censoring occurs.
III-B2 Ranking Objective
The $\mathcal{L}_{rank}$ objective is a ranking loss that penalizes an incorrect ordering of pairs $\{F_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x}_{i}),F_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x}_{j})\}$, where instance $i$ with event time $t_{i}<t_{j}$ should have a higher probability of failure than instance $j$ at time $t_{i}$:
$\mathcal{L}_{rank}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})$ $=\sum_{i\neq j}A_{i,j}\eta\bigg{(}F_{\theta}(t_{i}|\boldsymbol{x}_{i}),F_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x}_{j})\bigg{)},$ (8)
where $\eta(\mathbf{x},\mathbf{y})=\exp(-\frac{(\mathbf{x}-\mathbf{y})}{\sigma})$ and $A_{i,j}$ indicates an “acceptable” comparison pair [51, 52, 53]. We set $\sigma=1$ for all experiments.
III-B3 Combined Objective
The combined right censored negative log likelihood objective and the risk ranking objective, with risk ranking objective weight trade-off $w=\frac{1}{\text{batch\_size}}$ is
$\mathcal{L(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})}=-\mathcal{L}_{LL}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})+w\cdot\mathcal{L}_{rank}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e}).$ (9)
This combined objective provides a trade-off between finding the NN parameters $\boldsymbol{\theta}$ that maximize the likelihood of observing $\mathcal{D}$ and maximize a rank correlation between the instance risk scores and the observed failure times. We will denote the combined loss as the baseline objective. Our baseline model’s objective is similar to DRAFT [54] but with a different ranking objective function.
Thus far, we have described an objective function that improves predictive performance:
$\boldsymbol{\theta}^{*}=\arg\min_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})$ (10)
In the next section, we address how to develop NN models that are robust against adversarial attacks.
IV Preliminaries on Adversarial Training
Adversarial training [32, 55, 21, 56] is a technique to improve robustness and generalization of NNs to uncertain data. Rather than training on pristine data, adversarial training perturbs the data throughout the training process to encourage the model to optimize for parameters that are resilient against such perturbations. For example, [56] formulated adversarial training as solving a Min-Max robust optimization problem:
$\min_{\boldsymbol{\theta}}\max_{\boldsymbol{\tilde{X}}}\ \mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}}),$ (11)
where $\tilde{\boldsymbol{X}}=\{\tilde{\boldsymbol{x}}_{i}\in\mathcal{B}(\boldsymbol{x_{i}},\epsilon)\}_{i=1}^{N}$, and $\mathcal{B}(\boldsymbol{x_{i}},\epsilon)$ is the $\ell_{\infty}$-ball of radius $\epsilon$ around $\boldsymbol{x_{i}}$.
In practice, the Min-Max problem posed in [56] is intractable to compute exactly. In particular, the inner maximization requires finding a global optimum over a high dimensional and non-convex loss function over NN parameters. Therefore, many recent adversarial training methods propose approximations to Eq. 11, leading to models with varying degrees of robustness to different data uncertainties. To understand the robustness properties of SA models to a variety of data uncertainties present in real applications, this paper investigates several perturbation methods, described next.
IV-A Certifiable Robustness
One such approximation is to use techniques from NN verification during training. For example, CROWN-IBP combines Convex Relaxation Based Backward Bounding Pass (CROWN) and Interval Bound Propogation (IBP) methods to obtain a convex relaxation for the lower bound and upper bound of each layer in the NN with respect to input data [55, 57]. CROWN-IBP is commonly used for verifying NN properties over a range of possible inputs. [58, 55] showed the linear relaxation may be extended to the entire objective function:
[…]
This relaxation has several benefits: it is convex with respect to the NN weights, which enables efficient optimization, and by optimizing over an upper bound on the inner maximization, this relaxation represents accounting for a strong adversary during training.
IV-B Adversarial Attacks
Another way to approximate Eq. 19 is to use adversarial attacks. Whereas CROWN-IBP provides an upper bound on the inner maximization (i.e., $\max_{\boldsymbol{\tilde{X}}}\mathcal{L}(\theta;\boldsymbol{\tilde{X}})$), adversarial attack algorithms typically provide a lower bound:
$\max_{\boldsymbol{\tilde{X}}}\ \mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}})\approx\mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}_{FGSM}})$ (13)
Fast Gradient Sign Method (FGSM) [59] uses the gradient of the objective function with respect to the input $\boldsymbol{x}$ to perturb $\boldsymbol{x}$ in the direction that maximizes the objective $\mathcal{L}$. Projected Gradient Descent (PGD) [56] iteratively applies that same type of first-order perturbation $K$ times (or until convergence is reached), but constrains the perturbation to $\boldsymbol{x}$ within a specified set (e.g., $\ell_{p}$-ball). FGSM is a special case of PGD, with $K=1$ and the $\ell_{\infty}$-ball.
Considering the structure of the datasets as described in Section III, we perturb only the covariates $\boldsymbol{x}$ and not the time-to-event $t$. The perturbed covariates are calculated as:
$\boldsymbol{x}^{(0)}$ $=\boldsymbol{x}$ (14)
$\boldsymbol{x}^{(k+1)}$ $=\Pi_{\mathcal{B}(\boldsymbol{x},\epsilon)}\left(\boldsymbol{x}^{(k)}+\alpha\big{(}\nabla_{\boldsymbol{x}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{x}^{(k)},\boldsymbol{t},\boldsymbol{e})\big{)}\right),$ (15)
$\boldsymbol{\tilde{x}}$ $=\boldsymbol{x}^{(K)},$ (16)
where $k$ is the iteration step and $\alpha$ is a step size parameter.
IV-C Random Noise
While robustness to a worst-case perturbation is important for many applications, adding random noise to training data has also been shown to improve generalization performance [60]. Therefore, this paper also considers Gaussian Noise perturbations with variance $\epsilon$ bounded within an $\ell_{\infty}$ ball:
$\boldsymbol{z}$ $\sim\mathcal{N}(0,\boldsymbol{I}),$ (17)
$\boldsymbol{\tilde{x}}$ $=\boldsymbol{x}+\Pi_{\mathcal{B}(0,\epsilon)}(\sqrt{\epsilon}\boldsymbol{I}\boldsymbol{z}).$ (18)
We next introduce our approach, SAWAR, to incorporate adversarial regularization with SA models for improved calibration, predictive generalization, and adversarial robustness.
V Survival Analysis with Adversarial Regularization
As discussed in Section I, a key open challenge in SA is in developing models that are robust to uncertainty and variability in the datasets. This paper aims to improve the adversarial robustness of an NN against covariate perturbations within the SA framework. In doing so, this paper proposes the first method for enabling adversarial robustness for fully-parametric NN-based SA, which improves calibration and stability on unexpected or noisy data.
The following sections detail our approach, which extends the baseline objective function in Section III-B3 to include adversarial regularization posed as a Min-Max optimization problem in Section Section IV [21, 55].
V-A Min-Max Formulation
When solving for $\boldsymbol{\theta}$ in SA, many approaches treat each measurement in dataset $\mathcal{D}$ as ground-truth, ignoring aleatoric uncertainty [15]. Typically in SA, the aleatoric uncertainty is only associated with the noise of the log of the measured time-to-event (e.g. the Gumbel distribution or Log Normal distribution) [16, 6]. Instead, SAWAR approach assumes that each covariate is subject to a perturbation within a bounded uncertainty set, which leads to a robust optimization formulation as in Section IV. Thus, we optimize the NN parameters for the worst-case realizations of the baseline objective in Section III-B3:
$\min_{\boldsymbol{\theta}}\max_{\boldsymbol{\tilde{X}}}\ \mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}},\boldsymbol{t},\boldsymbol{e}).$ (19)
As NNs are highly non-convex, solving this optimization problem is difficult [61]. Therefore, we use the adversarial training methods discussed in Section IV to approximate/relax the inner maximization term of the Min-Max formulation.
V-B Adversarial Robustness Regularization
[21] shows that tight relaxations to the inner maximization of a Min-Max problem may over-regularize the NN, leading to poor predictive performance. Therefore, we extend the baseline objective in Eq. 9 with the adversarial robustness objective in Eq. 19 [55, 57] to balance predictive performance and adversarial robustness as:
$\min_{\boldsymbol{\theta}}\left[\kappa\cdot\mathcal{L}+(1-\kappa)\cdot\max_{\boldsymbol{\tilde{X}}}\mathcal{L}\right],$ (20)
where $\kappa\in(0,1)$.
We propose to use the open-source library auto_LiRPA [58] to relax the inner maximization term of the adversarial regularization with CROWN-IBP. auto_LiRPA seamlessly integrates with PyTorch for automatic backpropagation, computing a fully differentiable upper bound on the inner maximization term of Eq. 20 in a computationally feasible manner.
V-C SAWAR Training
We show the pseudo-code in Algorithm 1. The NN parameters are updated using stochastic gradient descent with the ADAM optimizer [62] (12), where we use early stopping [63] with respect to the validation set to prevent overfitting (25, 26 and 27). We introduce the covariate perturbation attacks after 100 epochs (4, 5, 6 and 7). A smooth $\epsilon$-scheduler linearly increases the $\epsilon$-perturbation magnitude from $0.0$ to $0.5$ within a 30 epoch window (28) after 100 epochs. We do not allow early stopping until after the maximum $\epsilon$-perturbation magnitude is achieved.
Algorithm 1 Robust Training with CROWN-IBP
1:procedure train_step(ModelLoss, Data, $\epsilon$, $\kappa$, Train)
2:   $loss_{epoch}\leftarrow 0$
3:   for $x_{B},t_{B},y_{B}$ in Data do
4:      if $\epsilon>1e^{-20}$ then
5:         lossnatural $\leftarrow\text{ModelLoss}(x_{B},t_{B},e_{B})$
6:         $\tilde{x}_{B}\leftarrow\text{PerturbationLpNorm}(x_{B},\epsilon)$
[…]
8:         loss $\leftarrow\kappa\times loss_{natural}+(1-\kappa)\times loss_{robust}$
9:      else
10:         loss $\leftarrow\text{ModelLoss}(x_{B},t_{B},e_{B})$
11:      if Train then
12:         ADAM Optimizer Step with loss.backward()
13:      lossepoch $\leftarrow$ lossepoch + loss
14:   return lossepoch
15:
16:procedure train_robust(Model, Datatr, Dataval, $\kappa$)
17:   Init random seeds, ADAM optimizer, lr-scheduler, $\epsilon$-scheduler
18:   ModelLoss = Losswrapper(Model)
[…]
20:   for $t$ in Epochs do
21:      if $\epsilon$-scheduler reached $\epsilon_{max}$ then
22:         lr-scheduler Step
[…]
26:         Save Best Model
27:         lossbest $\leftarrow$ lossval
28:      $\epsilon$-scheduler Step
29:   return Model
We expect that the model’s adversarial robustness and predictive calibration should increase when incorporating data uncertainty in SAWAR objective [34] via the adversarial regularization term in Eq. 20 and illustrate SAWAR hypothesis via extensive experiments in the next section.
VI Experiments
We conduct experiments over 10 benchmark SA datasets (VI-B) to quantify the impact of perturbation after the robust training is complete. As evaluation metrics, we consider the Concordance Index (CI) [64], the Integrated Brier Score (IBS) [65] and the Negative Log Likelihood (NegLL) at varying $\epsilon$-perturbation magnitudes from 0 to 1. We show that SAWAR improves predictive accuracy, calibration, and population curve estimation.
VI-A Configuration
The experiments are conducted on a 12th Gen Intel(R) Core (TM) 16 GB RAM i9-12900 processor running a Windows operating system. Each dataset undergoes stratified partitioning into train, validation, test sets with proportions 60%, 20%, 20% respectively. The NN $G_{\boldsymbol{\theta}}$ has two hidden layers of 50 neurons each and Leaky ReLU activation functions. lifelines [66], and sksurv [53] software packages are employed to ensure standardization and reproducibility in computing the SA metrics.
VI-B Datasets
We use the open-source API SurvSet [67], to download the benchmark SA datasets shown in Table I.
Dataset
$N$
$n_{fac}$
$n_{ohe}$
$n_{num}$
TRACE
1878
4
4
2
stagec
146
4
15
3
flchain
7874
4
26
6
Aids2
2839
3
11
1
Framingham
4699
2
12
5
dataDIVAT1
5943
3
14
2
prostate
502
6
16
9
zinc
431
11
18
2
retinopathy
394
5
9
2
LeukSurv
1043
2
24
5
TABLE I: Studied datasets from SurvSet. $n_{fac}$ is the number of categorical features, $n_{ohe}$ is number of binary features (one-hot-encoded) and $n_{num}$ is number of numerical features.
  1. 1.
    The TRACE [68] dataset is from a study on a subset of patients admitted after myocardial infarction to examine various risk factors.
  2. 2.
    The stagec [69] dataset is from a study exploring the prognostic value of flow cytometry for patients with stage C prostate cancer.
  3. 3.
    The flchain [70] dataset is from a study on the relationship between serum free light chain (FLC) and mortality of Olmsted Country residents aged 50 years or more.
  4. 4.
    The Aids2 [71] dataset is from patients diagnosed with AIDS in Australia before 1 July 1991.
  5. 5.
    The Framingham [72] dataset is from the first prospective study of cardiovascular disease with identification of risk factors and joint effects.
  6. 6.
    The dataDIVAT1 [73] dataset is from the first sample from the DIVAT Data Bank for French kidney transplant recipients from DIVAT cohort.
  7. 7.
    The prostate [74] dataset is from a randomised clinical trial comparing treatment for patients with stage 3 and stage 4 prostate cancer.
  8. 8.
    The zinc [75] is from the first study to examine association between tissue elemental zinc levels and the esophageal squamous cell carcinoma.
  9. 9.
    The retinopathy [76] is from the trial of laser coagulation as treatment to delay diabetic retinopathy.
  10. 10.
    The LeukSurv [77] is from the study on survival of acute myeloid leukemia and connection to spatial variation in survival.
We preprocess each dataset by one-hot encoding categorical covariates, followed by standard normalizing all covariates.
Figure 1: Distribution of survival curves for baseline, noise, FGSM, PGD, and CROWN-IBP. As the strength of the adversarial training method increases, from left to right column, the 90% credible interval narrows. However, the credible interval of SAWAR does not drastically differ from the baseline.
VI-C Adversary Setup
A detailed insight into the robust evaluation procedure is shown in Algorithm 2. We use two adversarial attack methods when benchmarking the robustness of the adversarial regularization method (14, 15 and 19): FGSM attack (Section IV-B) and a novel worst-case attack. We define the worst-case attack as the maximum failure rate with respect to the covariate uncertainty:
$\tilde{\lambda}_{\theta}=\max_{\boldsymbol{\tilde{x}_{i}}\in\mathcal{B}(\boldsymbol{x_{i}},\epsilon)}e^{G_{\boldsymbol{\theta}}(\boldsymbol{\tilde{x}}_{i})},$ (21)
and define the $\epsilon$-worst-case survival curve as:
$\tilde{S}(t|\boldsymbol{x})=e^{-\tilde{\lambda}_{\theta}t},$ (22)
which is the lowest survival probability at a given time $t$ with respect to the $\epsilon$-perturbed input covariates. For the exponential distribution, as the failure rate increases with the $\epsilon$-perturbation, the probability of survival decreases monotonically. We use CROWN-IBP for a tight approximation to the upper bound on the maximization term in equation Eq. 21. We evaluate the SA metrics for $\epsilon\in[0,1]$ for the two attacks methods to study how sensitive each adversarial training method is to covariate perturbation (1).
Note that when the activation functions are piecewise linear, an exact solution can also be obtained by solving a Mixed Integer Linear Program (MILP). However, solving the MILP is computationally prohibitive: even for $\epsilon=0.1$ each MILP takes on average 9-10 seconds when tested on the Dialysis dataset, whereas CROWN-IBP takes milliseconds to calculate over all the data in the Dialysis dataset.
Algorithm 2 Adversarial Robustness Evaluation
1:procedure eval_robust(Model, Data, Metric, Method, $\epsilon$)
2:   $X,t,e$ = Data
3:   MetricValues=[]
4:   for $\epsilon$ in $0:0.1:1$ do
5:      if $\epsilon=0$ then
6:         $\lambda_{\theta}\leftarrow$ Model(X)
7:         MetricValues = [MetricValues, $\text{Metric}(\lambda_{\theta},t,e)$]
8:      else
9:         $\tilde{\lambda}_{\theta}\leftarrow$ ATTACK(Model,Data,Method,$\epsilon$)
10:         MetricValues = [MetricValues, $\text{Metric}(\tilde{\lambda}_{\theta},t,e)$]
11:   return MetricValues
12:
13:procedure attack(Model, Data, Method, $\epsilon$)
14:   $X,t,e$ = Data
15:   if Method = FGSM then
16:      ModelLoss = Losswrapper(Model)
17:      $\tilde{X}\leftarrow$ FGSM(ModelLoss, $X,t,e,\epsilon$)
18:      $\tilde{\lambda}_{\theta}\leftarrow$ Model($\tilde{X}$)
19:   if Method = worst-case then
20:      $\tilde{X}\leftarrow\text{PerturbationLpNorm}(X,\epsilon)$
21:      $\tilde{\lambda}_{\theta}\leftarrow\text{CROWN-IBP}(\text{Model})(\tilde{X},t,e)$
22:   return $\tilde{\lambda}_{\theta}$
VI-D Results
$\epsilon$ Concordance Index Integrated Brier Score Negative Log Likelihood
baseline noise fgsm pgd SAWAR baseline noise fgsm pgd SAWAR baseline noise fgsm pgd SAWAR
1.00 2.9 3.35 3.35 2.5 2.9 4.6 4.4 2.4 1.9 1.7 4.0 4.7 2.7 2.1 1.5
0.90 3.1 3.5 3.45 2.45 2.5 4.5 4.5 2.5 1.8 1.7 4.0 4.7 2.9 1.9 1.5
0.80 3.4 3.4 3.2 2.35 2.65 4.6 4.4 2.6 1.8 1.6 4.1 4.6 3.0 1.9 1.4
0.70 3.45 3.35 3.1 2.4 2.7 4.7 4.3 2.6 1.8 1.6 4.2 4.5 3.0 2.0 1.3
0.60 3.45 3.45 3.0 2.5 2.6 4.7 4.3 2.7 1.8 1.5 4.2 4.5 3.1 2.1 1.1
0.50 3.85 3.65 3.1 2.1 2.3 4.4 4.6 2.9 1.8 1.3 4.1 4.5 3.2 2.1 1.1
0.40 3.9 3.8 3.05 2.1 2.15 4.4 4.6 3.0 1.9 1.1 4.0 4.6 3.2 2.1 1.1
0.30 4.4 4.2 3.0 1.9 1.5 4.4 4.6 3.0 1.9 1.1 4.0 4.6 3.2 2.1 1.1
0.20 4.35 4.0 3.3 2.25 1.1 4.2 4.6 3.2 2.0 1.0 3.9 4.7 3.2 2.1 1.1
0.10 4.25 4.0 3.3 2.35 1.1 4.0 4.6 3.2 2.2 1.0 3.8 4.8 3.1 2.3 1.0
0.05 3.85 4.0 3.45 2.6 1.1 3.8 4.7 2.9 2.2 1.4 3.9 4.7 3.1 2.3 1.0
0.00 2.95 3.95 3.35 3.35 1.4 3.0 3.9 3.3 2.5 2.3 2.6 4.1 3.5 2.8 2.0
TABLE II: FGSM average ranks for CI, IBS, and NegLL metrics across all datasets. Lower number is better.
$\epsilon$ Concordance Index Integrated Brier Score Negative Log Likelihood
baseline noise fgsm pgd SAWAR baseline noise fgsm pgd SAWAR baseline noise fgsm pgd SAWAR
1.00 2.8 3.4 3.5 3.6 1.7 2.6 4.3 3.8 3.3 1.0 3.1 4.5 3.4 3.0 1.0
0.90 3.1 3.3 3.5 3.4 1.7 3.0 4.3 3.6 3.1 1.0 3.1 4.5 3.5 2.9 1.0
0.80 3.0 3.3 3.5 3.7 1.5 3.0 4.5 3.5 3.0 1.0 3.0 4.6 3.5 2.9 1.0
0.70 3.05 3.3 3.6 3.65 1.4 3.0 4.5 3.5 3.0 1.0 3.1 4.5 3.5 2.9 1.0
0.60 2.85 3.55 3.55 3.75 1.3 3.0 4.5 3.5 3.0 1.0 3.0 4.7 3.5 2.8 1.0
0.50 3.2 3.9 3.35 3.45 1.1 3.2 4.4 3.5 2.9 1.0 3.0 4.7 3.5 2.8 1.0
0.40 3.25 3.95 3.2 3.5 1.1 3.3 4.8 3.2 2.7 1.0 3.1 4.7 3.4 2.8 1.0
0.30 3.0 3.05 3.95 3.6 1.4 3.5 4.7 3.3 2.5 1.0 3.3 4.9 3.2 2.6 1.0
0.20 2.05 2.95 4.2 3.8 2.0 3.7 4.7 3.4 2.2 1.0 3.4 4.9 3.4 2.3 1.0
0.10 2.85 3.4 3.4 3.35 2.0 3.3 4.5 3.0 2.3 1.9 3.5 4.5 3.2 2.2 1.6
0.05 3.05 4.15 3.4 2.9 1.5 2.6 4.4 3.0 2.6 2.4 3.3 4.4 3.0 2.4 1.9
0.00 2.95 3.95 3.35 3.35 1.4 3.0 3.9 3.3 2.5 2.3 2.6 4.1 3.5 2.8 2.0
TABLE III: Worst-Case average ranks for CI, IBS, and NegLL metrics across all datasets. Lower number is better.
Furthermore, we analyze how these adversarial attacks impact the characteristics of the population survival curves.
VI-D1 Failure Rate Distribution
Let us assume that the input covariates are random variables. Then the failure rate becomes a random variable that occurs from the stochasticity of the input covariates. Moreover, we may may treat $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ as a stochastic process:
$\boldsymbol{x}\sim p(\boldsymbol{x});\lambda\sim\lambda_{\boldsymbol{\theta}}(\boldsymbol{x})$ (23)
$S_{\boldsymbol{\theta}}(t)\sim e^{-\lambda t}$ (24)
Furthermore, we compute statistical quantities such as expectation, lower 5th quantile, and upper 95th quantile:
$LB=Q_{05}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x})),$ (25)
$UB=Q_{95}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x})).$ (26)
Fig. 1 visualizes the credible interval and mean of the survival function stochastic process in Eq. 24. The 90% credible interval of the the stochastic process changes with respect to the strength of the adversarial regularization method (Fig. 1). We observe the trend that the credible interval slightly narrows down as the strength of the adversarial regularization method increases (Fig. 1). The variance of the failure rate decreasing leads to a more robust model against adversarial attacks. For example, when $\mathrm{Var}(\lambda_{\theta}(\boldsymbol{x}))=0$, then there is a mode collapse such that $\lambda(\boldsymbol{x}_{i})=\lambda\textrm{ }\forall i$. This is the case for point-estimates of the parameter $\lambda$ derived via maximum likelihood estimation. A point-estimate results in a population curve, but no instance-level survival curves. Therefore, by utilizing a weighted objective in Eq. 20, we trade-off individualized survival curves with vulnerability to adversarial attacks and a population survival curve which is completely resilient to adversarial attacks.
Figure 2: Adversarial robustness of survival curves for baseline, noise, FGSM, PGD, and CROWN-IBP training against $\epsilon$-worst-case attack. As the strength of the adversarial training method increases, from left to right column, the perturbed population survival curve for a given $\epsilon$ is closer to the KMC.
However, despite exhibiting significantly higher adversarial robustness, the credible interval remains relatively unchanged compared to the baseline method.
VI-D2 Population Survival Curve
The $\epsilon$-perturbation magnitude determines the severity of the worst-case survival curve. We observe the trend that the stronger the adversarial regularization method, the ”closer” the $\epsilon$-worst-case population curve becomes to the unperturbed population survival curve for a given $\epsilon\in[0,1]$ (Fig. 2). Moreover, we use the Kaplan Meier Curve (KMC), a commonly used frequentist approach for non-parametrically estimating the population survival curve using samples $\{t_{i},e_{i}\}_{i=1}^{N}$, to visually evaluate model calibration. The close alignment between the population survival curve and the KMC shown in Fig. 2 demonstrates the strong calibration of SAWAR model [29].
We calculate the average ranking of each adversarial regularization method across all datasets with respect to CI, IBS, NegLL in response to the FGSM and $\epsilon$-worst-case attacks (Tables III and II). We empirically showed that CROWN-IBP adversarial regularization leads to increased performance, consistently ranking the best across various $\epsilon$-perturbation magnitudes.
The exact corresponding metric values for each $\epsilon$ of the worst-case attack are provided in Appendix A.
VII Conclusions
This paper is the first to introduce CROWN-IBP regularization for adversarial robustness training in SA with fully-parametric NN models. We benchmarked SAWAR method on 10 time-to-event datasets with standard adversarial regularization methods such as Gaussian Noise, FGSM, and PGD against the FGSM and worst-case attack. SAWAR’s empirical results indicate that CROWN-IBP regularization consistently enhances performance in SA with respect to CI and IBS, and NegLL for all $\epsilon$-perturbation magnitudes from 0 to 1. Moreover, compared to conventional adversarial regularization methods, SAWAR empirical evidence demonstrates that CROWN-IBP regularization is more resilient to input covariate perturbation attacks, where we see relative improvements to the baseline in CI and IBS of 1%-150% across all $\epsilon$-perturbation magnitudes (Appendix A). Thus adversarial regularization improves predictive performance, adversarial robustness, and calibration.
Future work will include the use the Weibull Proportional Hazard model to enable time varying hazard rates. This will require extending auto_LiRPA to handle the power operator for operations such as $t^{k}$, where $k$ is a learnable parameter.
Appendix A Appendix
We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1, the relative percentage change from the baseline to the adversarial training methods becomes larger and then smaller. The relative percent changes from the baseline to SAWAR for $\epsilon$ equal to 1.0, 0.9, 0.8, 0.7, 0.6, 0.5, 0.4, 0.3, 0.2, 0.1, 0.05, 0 is 29.46, 48.18, 56.27, 72.23, 102.44, 133.35, 152.17, 140.33, 91.78, 21.17, 9.63, 1.07 respectively. We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
Aids2 baseline 0.499 0.497 0.501 0.505 0.507 0.509 0.517 0.535 0.565 0.573 0.573 0.572
noise 0.492 0.494 0.496 0.5 0.5 0.498 0.507 0.526 0.556 0.556 0.554 0.554
fgsm 0.498 0.495 0.498 0.497 0.5 0.502 0.508 0.529 0.554 0.557 0.555 0.553
pgd 0.498 0.497 0.497 0.5 0.501 0.501 0.507 0.531 0.557 0.558 0.556 0.554
SAWAR 0.567 0.575 0.577 0.579 0.58 0.58 0.579 0.579 0.579 0.579 0.579 0.579
Framingham baseline 0.612 0.623 0.635 0.646 0.656 0.664 0.672 0.678 0.682 0.686 0.687 0.688
noise 0.606 0.616 0.626 0.635 0.642 0.649 0.659 0.674 0.683 0.685 0.685 0.685
fgsm 0.607 0.611 0.617 0.622 0.626 0.635 0.652 0.669 0.682 0.686 0.687 0.686
pgd 0.567 0.578 0.588 0.599 0.61 0.624 0.642 0.659 0.68 0.688 0.688 0.687
SAWAR 0.694 0.697 0.699 0.701 0.702 0.704 0.704 0.705 0.705 0.704 0.704 0.704
LeukSurv baseline 0.611 0.616 0.623 0.623 0.627 0.631 0.63 0.631 0.63 0.624 0.624 0.626
noise 0.552 0.557 0.562 0.566 0.569 0.574 0.587 0.61 0.628 0.627 0.621 0.617
fgsm 0.546 0.549 0.554 0.563 0.561 0.566 0.583 0.602 0.624 0.624 0.621 0.62
pgd 0.537 0.536 0.538 0.549 0.546 0.558 0.577 0.598 0.62 0.622 0.619 0.617
SAWAR 0.483 0.498 0.51 0.532 0.553 0.583 0.589 0.607 0.617 0.623 0.632 0.636
TRACE baseline 0.564 0.586 0.612 0.637 0.665 0.69 0.714 0.73 0.734 0.733 0.735 0.735
noise 0.565 0.588 0.613 0.64 0.668 0.692 0.716 0.73 0.733 0.735 0.734 0.735
fgsm 0.553 0.577 0.603 0.63 0.658 0.688 0.713 0.722 0.727 0.73 0.733 0.734
pgd 0.557 0.582 0.609 0.636 0.664 0.696 0.723 0.731 0.733 0.732 0.733 0.733
SAWAR 0.73 0.733 0.736 0.738 0.739 0.739 0.74 0.74 0.739 0.738 0.738 0.737
dataDIVAT1 baseline 0.534 0.558 0.583 0.606 0.624 0.639 0.65 0.658 0.662 0.664 0.664 0.664
noise 0.534 0.539 0.551 0.571 0.587 0.612 0.636 0.656 0.654 0.64 0.635 0.631
fgsm 0.575 0.585 0.594 0.606 0.617 0.628 0.638 0.644 0.651 0.643 0.638 0.635
pgd 0.517 0.529 0.548 0.568 0.591 0.618 0.636 0.654 0.657 0.651 0.647 0.644
SAWAR 0.627 0.642 0.652 0.658 0.659 0.659 0.659 0.658 0.657 0.656 0.656 0.655
flchain baseline 0.096 0.098 0.102 0.107 0.114 0.127 0.156 0.461 0.9 0.916 0.922 0.924
noise 0.096 0.099 0.104 0.112 0.127 0.164 0.335 0.768 0.912 0.921 0.923 0.923
fgsm 0.102 0.116 0.159 0.288 0.599 0.852 0.91 0.917 0.923 0.927 0.927 0.925
pgd 0.102 0.127 0.235 0.538 0.824 0.904 0.917 0.921 0.925 0.929 0.929 0.929
SAWAR 0.84 0.925 0.927 0.929 0.93 0.93 0.929 0.929 0.93 0.93 0.931 0.931
prostate baseline 0.412 0.419 0.426 0.447 0.463 0.49 0.53 0.581 0.614 0.635 0.639 0.643
noise 0.49 0.503 0.514 0.524 0.533 0.541 0.553 0.566 0.576 0.572 0.577 0.576
fgsm 0.514 0.522 0.526 0.536 0.541 0.549 0.558 0.573 0.582 0.596 0.596 0.596
pgd 0.522 0.526 0.532 0.535 0.54 0.548 0.557 0.567 0.576 0.593 0.591 0.593
SAWAR 0.575 0.584 0.598 0.6 0.608 0.632 0.638 0.638 0.639 0.641 0.642 0.644
retinopathy baseline 0.525 0.539 0.562 0.569 0.587 0.592 0.592 0.599 0.605 0.601 0.604 0.605
noise 0.526 0.545 0.561 0.57 0.581 0.588 0.588 0.6 0.603 0.6 0.604 0.606
fgsm 0.47 0.485 0.51 0.536 0.569 0.578 0.591 0.6 0.599 0.602 0.606 0.607
pgd 0.471 0.481 0.509 0.53 0.567 0.577 0.591 0.599 0.598 0.601 0.606 0.607
SAWAR 0.588 0.597 0.611 0.616 0.619 0.615 0.621 0.62 0.62 0.619 0.617 0.617
stagec baseline 0.442 0.457 0.476 0.495 0.514 0.548 0.582 0.668 0.721 0.697 0.688 0.688
noise 0.471 0.481 0.505 0.49 0.514 0.534 0.582 0.697 0.721 0.712 0.688 0.697
fgsm 0.423 0.438 0.447 0.471 0.51 0.567 0.62 0.663 0.688 0.678 0.688 0.688
pgd 0.462 0.486 0.471 0.495 0.51 0.567 0.611 0.668 0.692 0.688 0.697 0.697
SAWAR 0.428 0.447 0.481 0.514 0.567 0.639 0.673 0.692 0.702 0.692 0.692 0.707
zinc baseline 0.339 0.349 0.356 0.376 0.427 0.507 0.657 0.758 0.796 0.796 0.792 0.793
noise 0.315 0.329 0.339 0.376 0.424 0.545 0.704 0.778 0.791 0.805 0.806 0.81
fgsm 0.339 0.357 0.384 0.439 0.55 0.657 0.738 0.776 0.787 0.804 0.809 0.812
pgd 0.333 0.351 0.384 0.439 0.55 0.671 0.738 0.776 0.79 0.8 0.809 0.808
SAWAR 0.383 0.464 0.533 0.649 0.707 0.746 0.766 0.779 0.789 0.787 0.795 0.8
TABLE IV: Concordance Index metric for SurvSet datasets (higher is better) for each adversarial training method against the worst-case adversarial attack.
We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1, the relative percentage change from the baseline to the adversarial training methods becomes larger and then smaller. The relative percent changes from the baseline to SAWAR for $\epsilon$ equal to 1.0, 0.9, 0.8, 0.7, 0.6, 0.5, 0.4, 0.3, 0.2, 0.1, 0.05, 0 is -44.13, -45.46, -46.68, -47.57, -47.86, -47.10, -45.03, -40.89, -33.06, -20.29, -11.34, -1.68 respectively (where lower percentage change is better). We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
Aids2 baseline 0.24 0.233 0.224 0.212 0.197 0.178 0.157 0.137 0.124 0.12 0.121 0.122
noise 0.254 0.253 0.25 0.245 0.236 0.22 0.196 0.161 0.13 0.121 0.121 0.122
fgsm 0.252 0.249 0.244 0.236 0.223 0.203 0.176 0.145 0.125 0.121 0.121 0.121
pgd 0.252 0.248 0.243 0.234 0.221 0.2 0.172 0.143 0.124 0.121 0.121 0.121
SAWAR 0.126 0.124 0.124 0.124 0.123 0.123 0.123 0.123 0.123 0.123 0.123 0.124
Framingham baseline 0.81 0.8 0.778 0.733 0.65 0.52 0.369 0.243 0.165 0.128 0.119 0.115
noise 0.818 0.818 0.816 0.809 0.778 0.681 0.495 0.294 0.174 0.128 0.119 0.116
fgsm 0.818 0.816 0.81 0.791 0.728 0.58 0.369 0.208 0.14 0.12 0.116 0.115
pgd 0.818 0.818 0.817 0.811 0.784 0.683 0.461 0.235 0.138 0.117 0.115 0.115
SAWAR 0.165 0.152 0.141 0.133 0.127 0.122 0.118 0.116 0.114 0.114 0.113 0.113
LeukSurv baseline 0.163 0.163 0.163 0.163 0.162 0.161 0.156 0.144 0.126 0.109 0.105 0.103
noise 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.157 0.134 0.12 0.114
fgsm 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.161 0.148 0.122 0.114 0.111
pgd 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.16 0.145 0.12 0.114 0.111
SAWAR 0.154 0.15 0.144 0.136 0.128 0.12 0.115 0.112 0.11 0.107 0.107 0.106
TRACE baseline 0.634 0.634 0.633 0.626 0.603 0.549 0.449 0.329 0.234 0.183 0.172 0.167
noise 0.634 0.634 0.632 0.622 0.596 0.536 0.435 0.319 0.23 0.183 0.172 0.167
fgsm 0.634 0.634 0.634 0.631 0.613 0.555 0.434 0.294 0.212 0.177 0.171 0.168
pgd 0.634 0.634 0.632 0.621 0.58 0.491 0.362 0.245 0.189 0.171 0.167 0.166
SAWAR 0.282 0.254 0.231 0.213 0.198 0.186 0.177 0.17 0.166 0.163 0.163 0.163
dataDIVAT1 baseline 0.721 0.702 0.67 0.616 0.531 0.417 0.308 0.234 0.195 0.18 0.178 0.178
noise 0.749 0.749 0.748 0.748 0.743 0.72 0.626 0.413 0.248 0.198 0.192 0.192
fgsm 0.749 0.749 0.748 0.747 0.74 0.699 0.547 0.322 0.215 0.193 0.192 0.192
pgd 0.749 0.749 0.749 0.748 0.746 0.724 0.582 0.305 0.192 0.184 0.186 0.187
SAWAR 0.266 0.241 0.224 0.212 0.202 0.193 0.187 0.183 0.18 0.178 0.178 0.178
flchain baseline 0.831 0.831 0.831 0.831 0.831 0.83 0.829 0.818 0.486 0.096 0.07 0.052
noise 0.831 0.831 0.831 0.831 0.831 0.83 0.829 0.807 0.25 0.087 0.063 0.052
fgsm 0.831 0.831 0.831 0.83 0.83 0.828 0.707 0.146 0.094 0.06 0.054 0.051
pgd 0.831 0.831 0.831 0.83 0.83 0.797 0.298 0.114 0.079 0.055 0.051 0.05
SAWAR 0.522 0.161 0.076 0.063 0.057 0.054 0.052 0.051 0.05 0.05 0.049 0.049
prostate baseline 0.53 0.529 0.528 0.526 0.519 0.505 0.475 0.418 0.335 0.256 0.231 0.218
noise 0.53 0.53 0.53 0.53 0.53 0.53 0.53 0.528 0.477 0.371 0.325 0.296
fgsm 0.53 0.53 0.53 0.53 0.53 0.53 0.529 0.509 0.413 0.29 0.257 0.246
pgd 0.53 0.53 0.53 0.53 0.53 0.53 0.528 0.502 0.396 0.28 0.254 0.244
SAWAR 0.508 0.482 0.441 0.393 0.348 0.309 0.278 0.251 0.229 0.214 0.209 0.205
retinopathy baseline 0.662 0.657 0.645 0.621 0.582 0.522 0.43 0.324 0.244 0.203 0.196 0.195
noise 0.662 0.657 0.645 0.622 0.583 0.523 0.432 0.326 0.245 0.204 0.196 0.195
fgsm 0.656 0.646 0.626 0.597 0.556 0.493 0.401 0.301 0.232 0.201 0.196 0.197
pgd 0.656 0.645 0.625 0.596 0.554 0.49 0.397 0.299 0.231 0.201 0.197 0.198
SAWAR 0.526 0.484 0.436 0.385 0.334 0.29 0.254 0.228 0.211 0.203 0.202 0.202
stagec baseline 0.505 0.505 0.505 0.504 0.502 0.486 0.431 0.329 0.211 0.175 0.181 0.192
noise 0.505 0.505 0.505 0.504 0.502 0.485 0.432 0.338 0.216 0.175 0.181 0.192
fgsm 0.505 0.505 0.505 0.503 0.494 0.458 0.382 0.268 0.174 0.161 0.169 0.178
pgd 0.505 0.505 0.505 0.503 0.493 0.454 0.377 0.261 0.169 0.159 0.167 0.176
SAWAR 0.491 0.476 0.45 0.408 0.354 0.296 0.24 0.195 0.168 0.16 0.162 0.167
zinc baseline 0.907 0.907 0.905 0.899 0.882 0.823 0.652 0.368 0.18 0.109 0.096 0.09
noise 0.908 0.907 0.907 0.904 0.893 0.838 0.625 0.32 0.168 0.109 0.096 0.089
fgsm 0.907 0.907 0.904 0.892 0.839 0.659 0.365 0.194 0.124 0.095 0.089 0.086
pgd 0.907 0.907 0.903 0.89 0.832 0.64 0.348 0.187 0.122 0.094 0.088 0.087
SAWAR 0.72 0.609 0.476 0.352 0.26 0.198 0.159 0.132 0.113 0.1 0.096 0.094
TABLE V: Integrated Brier Score metric for SurvSet datasets (lower is better) for each adversarial training method against the worst-case adversarial attack.
$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
Aids2 baseline 1.14e+04 7.22e+03 4.60e+03 2.87e+03 1.80e+03 1.16e+03 7.95e+02 6.06e+02 5.32e+02 5.16e+02 5.17e+02 5.19e+02
noise 3.84e+05 1.61e+05 6.68e+04 2.78e+04 1.14e+04 4.42e+03 1.83e+03 8.54e+02 5.64e+02 5.20e+02 5.19e+02 5.21e+02
fgsm 6.77e+04 3.50e+04 1.78e+04 8.89e+03 4.35e+03 2.14e+03 1.10e+03 6.66e+02 5.35e+02 5.19e+02 5.19e+02 5.20e+02
pgd 5.49e+04 2.86e+04 1.49e+04 7.58e+03 3.79e+03 1.92e+03 1.02e+03 6.42e+02 5.31e+02 5.19e+02 5.19e+02 5.19e+02
SAWAR 5.29e+02 5.24e+02 5.22e+02 5.22e+02 5.21e+02 5.21e+02 5.21e+02 5.21e+02 5.21e+02 5.22e+02 5.22e+02 5.22e+02
Framingham baseline 1.47e+05 7.07e+04 3.42e+04 1.68e+04 8.47e+03 4.57e+03 2.79e+03 2.01e+03 1.68e+03 1.54e+03 1.51e+03 1.50e+03
noise 1.27e+07 3.17e+06 8.03e+05 2.06e+05 5.38e+04 1.48e+04 4.92e+03 2.39e+03 1.72e+03 1.54e+03 1.51e+03 1.50e+03
fgsm 1.95e+06 5.96e+05 1.84e+05 5.70e+04 1.81e+04 6.33e+03 2.84e+03 1.85e+03 1.58e+03 1.51e+03 1.50e+03 1.49e+03
pgd 2.01e+07 4.36e+06 9.57e+05 2.13e+05 4.88e+04 1.21e+04 3.85e+03 1.97e+03 1.57e+03 1.50e+03 1.49e+03 1.49e+03
SAWAR 1.66e+03 1.62e+03 1.58e+03 1.56e+03 1.53e+03 1.52e+03 1.50e+03 1.50e+03 1.49e+03 1.49e+03 1.49e+03 1.49e+03
LeukSurv baseline 1.79e+07 4.80e+06 1.29e+06 3.48e+05 9.29e+04 2.53e+04 6.59e+03 1.77e+03 5.76e+02 3.10e+02 2.74e+02 2.58e+02
noise 5.46e+23 1.79e+21 7.98e+18 2.52e+16 8.53e+13 1.88e+11 5.25e+08 2.35e+06 3.37e+04 2.60e+03 1.21e+03 7.21e+02
fgsm 8.34e+16 1.79e+15 3.10e+13 4.86e+11 8.13e+09 1.49e+08 2.72e+06 7.03e+04 3.45e+03 6.20e+02 4.31e+02 3.59e+02
pgd 1.28e+15 3.89e+13 1.03e+12 2.67e+10 7.96e+08 2.30e+07 7.01e+05 2.80e+04 1.98e+03 4.65e+02 3.58e+02 3.17e+02
SAWAR 2.07e+04 8.38e+03 3.50e+03 1.45e+03 6.62e+02 3.45e+02 2.91e+02 2.70e+02 2.60e+02 2.54e+02 2.52e+02 2.50e+02
TRACE baseline 4.56e+09 4.41e+08 4.44e+07 4.61e+06 5.31e+05 6.71e+04 9.88e+03 2.20e+03 9.05e+02 6.03e+02 5.53e+02 5.32e+02
noise 9.11e+08 1.07e+08 1.34e+07 1.74e+06 2.43e+05 3.71e+04 6.63e+03 1.78e+03 8.35e+02 5.93e+02 5.51e+02 5.32e+02
fgsm 2.01e+11 9.10e+09 4.31e+08 2.24e+07 1.26e+06 8.26e+04 7.79e+03 1.61e+03 7.45e+02 5.74e+02 5.47e+02 5.35e+02
pgd 4.53e+08 4.81e+07 5.35e+06 6.12e+05 7.63e+04 1.07e+04 2.09e+03 8.37e+02 5.94e+02 5.40e+02 5.31e+02 5.26e+02
SAWAR 8.79e+02 7.60e+02 6.82e+02 6.29e+02 5.94e+02 5.69e+02 5.53e+02 5.41e+02 5.33e+02 5.28e+02 5.27e+02 5.27e+02
dataDIVAT1 baseline 1.43e+04 8.36e+03 4.95e+03 3.01e+03 1.91e+03 1.31e+03 9.99e+02 8.51e+02 7.84e+02 7.56e+02 7.51e+02 7.49e+02
noise 5.69e+08 7.64e+07 1.05e+07 1.44e+06 2.07e+05 3.17e+04 5.75e+03 1.61e+03 9.08e+02 7.81e+02 7.67e+02 7.64e+02
fgsm 2.48e+08 3.36e+07 4.65e+06 6.56e+05 9.64e+04 1.54e+04 3.12e+03 1.14e+03 8.16e+02 7.65e+02 7.61e+02 7.60e+02
pgd 1.85e+11 8.02e+09 3.47e+08 1.75e+07 9.39e+05 6.14e+04 5.80e+03 1.27e+03 8.03e+02 7.57e+02 7.56e+02 7.57e+02
SAWAR 9.18e+02 8.55e+02 8.21e+02 8.00e+02 7.86e+02 7.75e+02 7.66e+02 7.59e+02 7.55e+02 7.52e+02 7.51e+02 7.51e+02
flchain baseline 3.35e+23 9.34e+20 2.56e+18 7.37e+15 2.23e+13 6.82e+10 2.29e+08 9.33e+05 1.67e+04 2.03e+03 1.25e+03 1.09e+03
noise 2.18e+35 1.20e+31 7.58e+26 4.68e+22 3.48e+18 2.65e+14 2.21e+10 6.77e+06 4.45e+04 1.96e+03 1.22e+03 1.11e+03
fgsm 1.61e+20 3.47e+17 7.79e+14 1.86e+12 5.49e+09 4.02e+07 6.37e+05 1.86e+04 2.15e+03 1.15e+03 1.09e+03 1.08e+03
pgd 1.30e+16 7.91e+13 4.93e+11 3.47e+09 4.60e+07 1.33e+06 4.93e+04 4.99e+03 1.45e+03 1.10e+03 1.08e+03 1.07e+03
SAWAR 6.31e+03 2.22e+03 1.43e+03 1.19e+03 1.12e+03 1.09e+03 1.08e+03 1.07e+03 1.07e+03 1.07e+03 1.07e+03 1.07e+03
prostate baseline 9.74e+05 3.22e+05 1.06e+05 3.58e+04 1.21e+04 4.25e+03 1.64e+03 7.62e+02 4.66e+02 3.70e+02 3.51e+02 3.44e+02
noise 7.00e+20 5.05e+18 3.75e+16 2.85e+14 2.28e+12 2.03e+10 2.59e+08 9.04e+06 4.85e+05 6.66e+04 2.94e+04 1.37e+04
fgsm 3.16e+14 1.04e+13 3.40e+11 1.16e+10 4.07e+08 1.52e+07 6.47e+05 3.56e+04 3.75e+03 1.03e+03 7.20e+02 5.82e+02
pgd 8.84e+12 4.29e+11 2.18e+10 1.14e+09 5.94e+07 3.27e+06 1.92e+05 1.44e+04 1.91e+03 6.40e+02 5.00e+02 4.39e+02
SAWAR 9.65e+03 3.88e+03 1.66e+03 8.48e+02 5.43e+02 4.30e+02 3.83e+02 3.59e+02 3.45e+02 3.37e+02 3.35e+02 3.34e+02
retinopathy baseline 1.11e+04 5.64e+03 2.88e+03 1.51e+03 8.08e+02 4.64e+02 2.98e+02 2.18e+02 1.84e+02 1.70e+02 1.68e+02 1.67e+02
noise 1.11e+04 5.61e+03 2.88e+03 1.51e+03 8.13e+02 4.68e+02 2.99e+02 2.19e+02 1.84e+02 1.70e+02 1.68e+02 1.68e+02
fgsm 7.43e+03 3.89e+03 2.07e+03 1.12e+03 6.33e+02 3.86e+02 2.63e+02 2.03e+02 1.78e+02 1.68e+02 1.67e+02 1.68e+02
pgd 7.30e+03 3.83e+03 2.04e+03 1.11e+03 6.25e+02 3.81e+02 2.60e+02 2.02e+02 1.77e+02 1.68e+02 1.67e+02 1.68e+02
SAWAR 4.02e+02 3.31e+02 2.79e+02 2.41e+02 2.14e+02 1.95e+02 1.82e+02 1.74e+02 1.69e+02 1.67e+02 1.67e+02 1.67e+02
stagec baseline 1.16e+06 2.94e+05 7.50e+04 1.85e+04 4.81e+03 1.28e+03 3.57e+02 1.14e+02 5.35e+01 4.10e+01 3.98e+01 3.97e+01
noise 2.07e+06 4.93e+05 1.14e+05 2.73e+04 6.49e+03 1.61e+03 4.19e+02 1.25e+02 5.50e+01 4.10e+01 3.96e+01 3.95e+01
fgsm 1.80e+05 5.45e+04 1.67e+04 5.23e+03 1.64e+03 5.30e+02 1.77e+02 7.43e+01 4.54e+01 3.96e+01 3.91e+01 3.92e+01
pgd 1.54e+05 4.77e+04 1.47e+04 4.63e+03 1.47e+03 4.84e+02 1.69e+02 7.23e+01 4.48e+01 3.93e+01 3.89e+01 3.90e+01
SAWAR 1.03e+03 5.79e+02 3.30e+02 1.94e+02 1.21e+02 8.21e+01 6.10e+01 4.88e+01 4.22e+01 3.89e+01 3.83e+01 3.81e+01
zinc baseline 2.62e+06 5.85e+05 1.33e+05 3.13e+04 7.43e+03 1.87e+03 5.36e+02 2.03e+02 1.11e+02 8.26e+01 7.71e+01 7.46e+01
noise 2.93e+07 4.72e+06 7.76e+05 1.32e+05 2.34e+04 4.45e+03 9.48e+02 2.56e+02 1.14e+02 8.17e+01 7.69e+01 7.53e+01
fgsm 2.50e+06 4.44e+05 9.49e+04 2.11e+04 4.89e+03 1.24e+03 3.66e+02 1.44e+02 9.01e+01 7.66e+01 7.53e+01 7.57e+01
pgd 2.55e+06 4.38e+05 8.27e+04 1.87e+04 4.43e+03 1.13e+03 3.40e+02 1.39e+02 8.88e+01 7.66e+01 7.55e+01 7.60e+01
SAWAR 1.31e+03 7.77e+02 4.72e+02 2.98e+02 1.98e+02 1.42e+02 1.10e+02 9.21e+01 8.28e+01 7.87e+01 7.80e+01 7.78e+01
TABLE VI: Negative Log Likelihood metric for SurvSet datasets (lower is better) for each adversarial training method against the worst-case adversarial attack.

### v3

@@
-Robust Survival Analysis with Adversarial Regularization
+Survival Analysis with Adversarial Regularization
 Abstract
-Survival Analysis (SA) is about modeling the time for an event of interest to occur, which has important applications in many fields, including medicine, defense, finance, and aerospace.
-Recent work has demonstrated the benefits of using Neural Networks (NNs) to capture complicated relationships in SA.
-However, the datasets used to train these models are often subject to uncertainty (e.g., noisy measurements, human error), which we show can substantially degrade the performance of existing techniques.
-To address this issue, this work leverages recent advances in NN verification to provide new algorithms for generating fully-parametric survival models that are robust to such uncertainties.
-In particular, we introduce a robust loss function for training the models and use CROWN-IBP regularization to address the computational challenges with solving the resulting Min-Max problem.
-To evaluate the proposed approach, we apply relevant perturbations to publicly available datasets in the SurvSet repository and compare survival models against several baselines.
-We empirically show that Survival Analysis with Adversarial Regularization (SAWAR) method on average ranks best for dataset perturbations of varying magnitudes on metrics such as Negative Log Likelihood (NegLL), Integrated Brier Score (IBS), and Concordance Index (CI), concluding that adversarial regularization enhances performance in SA.
+Survival Analysis (SA) models the time until an event occurs, with applications in fields like medicine, defense, finance, and aerospace.
+Recent work shows that Neural Networks (NNs) can capture complex relationships in SA.
+However, dataset uncertainties (e.g., noisy measurements, human error) can degrade model performance.
+To address this, we leverage NN verification advances to create algorithms for robust, fully-parametric survival models.
+We introduce a robust loss function and use CROWN-IBP regularization to handle computational challenges in the Min-Max problem.
+Evaluating our approach on SurvSet datasets, we find that our Survival Analysis with Adversarial Regularization (SAWAR) method consistently outperforms baselines under various perturbations with respect to Negative Log Likelihood (NegLL), Integrated Brier Score (IBS), and Concordance Index (CI).
+This demonstrates that adversarial regularization enhances SA performance and calibration, mitigating data uncertainty and improving generalization across diverse datasets up to 150% across all perturbation magnitudes.
 Code: https://github.com/mlpotter/SAWAR
 Index Terms:
-Survival Analysis, Adversarial Robustness, Neural Networks, Calibration, CROWN-IBP
-publicationid: pubid: $0$–$00$ 2024 IEEE
+Survival Analysis, Adversarial Robustness, Neural Networks, Calibration, CROWN-IBP, Cox-Weibull
 I Introduction
-Survival Analysis (SA) is a highly active research field [1, 2] used to model the time before an event happens.
-For example, SA has been used extensively in medicine to quantify the effectiveness of various treatments as part of studies for cancer patients and drug trials [3, 4, 5].
-While early approaches (dating back to the 1950s) predominantly used linear parametric models [6, 7], recent methods often leverage Neural Networks (NNs) to model complicated time-to-event data [8].
-A key open challenge in this field is in developing models that are robust to uncertainty and variability in the datasets.
-Successful generalization of machine learning models over unseen datasets is crucial for clinical applications, as efficacy of developed drugs or treatments should be demonstrated over diverse patient demographics, institutions, and data acquisition settings [9, 10, 11].
-Data uncertainty could be caused by human errors (e.g., notoriously poor handwriting of physicians [12], ineffective medical record keeping [13]), or inherent noise during transmissions or data collection [14].
-However, many approaches assume perfect measurement of the covariates, ignoring aleatoric uncertainty [15].
-In SA, the aleatoric uncertainty is only associated with the noise of the log of the measured time-to-event [16, 6].
-Moreover, several studies have shown that adversaries can influence the final model to be over-/under-confident by purposefully poisoning studies or otherwise disrupting the data collection activities [17, 18].
-Despite these known failure modes, and coupled with the sensitivity of NNs to input uncertainties, addressing the robustness of NNs against $\epsilon$-perturbations trained specifically for fully-parametric SA remains unexplored.
+Survival Analysis (SA) models the time until an event occurs and has extensive applications.
+For example, SA has been used extensively in medicine for evaluating treatments in cancer patients and drug trials [1].
+Early methods relied on linear parametric models [2], but recent advancements utilize Neural Networks (NNs) to handle complex time-to-event data [3].
+A major challenge in Survival Analysis (SA) is developing models robust to dataset uncertainty and variability.
+This is particularly essential for clinical applications where treatment efficacy must be demonstrated across diverse patient demographics, institutions, and data acquisition settings [4, 5, 6].
+Data uncertainty can arise from human errors, such as poor handwriting by physicians [7] and ineffective medical record-keeping [8], as well as inherent noise during data collection [9].
+Many approaches assume perfect measurement of covariates, neglecting aleatoric uncertainty [10], which in SA is associated with noise in the log of measured time-to-event [11].
+Further, adversaries can manipulate data to make models over- or under-confident [12].
+Despite these issues and the sensitivity of NNs to input uncertainties, robustness techniques against data perturbations in SA remains underexplored.
 To address these issues, this paper’s contributions include:
-  • identification of the sensitivity of existing SA methods, such as Deep Regularized Accelerated Failure Time (DRAFT) [19], to covariate perturbations via evaluating open-source datasets with a wide range of metrics related to predictive accuracy, calibration, and population curve estimation.
-  • our approach, Survival Analysis with Adversarial Regularization (SAWAR), which is the first method for enabling adversarial robustness to covariate perturbations within a bounded uncertainty set for fully-parametric NN-based SA.
-We improve calibration and stability on unexpected or noisy data by introducing a loss function based on a Min-Max formulation and Convex Relaxation Based Backward Bounding Pass with Interval Bound Propagation (CROWN-IBP) [20, 21].
-The paper is structured as follows: Section II describes the State of the Art; Section III provides background to SA terminology and dataset structure; Section IV provides background on the adversarial and perturbation attacks used; Section V discusses the custom loss employed to deal with those attacks; Section VI provides the results of the experimental activity; finally, Section VII derives conclusions on our findings.
+  • Survival Analysis with Adversarial Regularization (SAWAR) for training robust, fully-parametric, NN-based SA models using adversarial covariate perturbations and convex relaxations, enhancing model calibration and stability against noisy data.
+  • Experiments demonstrating the sensitivity of existing NN-based SA methods to covariate perturbations across 10 publicly available medical SA datasets, highlighting an area for improvement in the current state of the art.
 II Related Work
-This section focuses on the literature about NNs for SA, with further detail about NN implementations available in a recent survey paper [22].
-The first use of NNs in this field was a Feedforward Neural Network (FFNN) to approximate the survival function [23], defined as the probability that no event occurs up to a specific time $t$.
-More recently, [24] proposed Cox-nnet, a NN that outputs a proportional-hazard [25, 6], which is the assumption that the difference in effect between different instances is constant as a function of time.
-Building on this idea, [26] developed DeepSurv, a Cox proportional-hazard NN for personalized treatment, which uses dropout layers to improve regularization and reduce overfitting.
-Instead of proportional hazard estimation, DeepWeiSurv [27] regresses the parameters of a Weibull mixture distribution with NNs.
-Several recent works in SA aim to improve model calibration (e.g., by using new terms in the objective function [28, 29, 30]).
-A calibrated model provides a predicted confidence that is indicative of the true likelihood that the model is correct [31].
-Meanwhile, few works have addressed adversarial robustness in SA.
-In general, a model is adversarially robust if small perturbations to the input do not cause the model to make incorrect predictions [32].
-Adversarial robustness and model calibration are loosely coupled: while [33] showed that better calibration does not necessarily improve generalization, [34] showed that incorporating adversarial robustness during training improves trustworthiness in calibration and stability on unexpected data.
-Existing adversarial training methods for SA face important tradeoffs between model complexity, performance, and ease of embedding of physical assumptions.
-Most of these techniques use Conditional Generative Adversarial Networks (CGANs) to generate samples of the time-to-event from a survival distribution in a non-parametric manner [19, 35, 36].
-These models (e.g., CGANs) do not assume a functional form and define the survival distribution implicitly [37, 38, 39].
-Therefore, there is no mathematical expression that describes the time-to-event Probability Density Function (PDF) value given covariates.
-One benefit of implicit models is that increased model complexity leads to increased flexibility, which is the ability to fit many other potential data realizations [40].
-However, the higher model complexity generally violates the Bayesian Occam’s Razor, which penalizes models for having too much flexibility [40].
-To satisfy the Bayesian Occam’s Razor’s “just right“ model complexity, modeling assumptions known as inductive biases are employed in the form of an explicit parametric model [39, 37].
-For example, a common inductive bias for SA models is the memoryless property [38], which the exponential distribution satisfies.
-For this reason, this paper uses the exponential distribution, which is commonly used in SA [41, 42, 43] (and more generally in queuing theory [44, 45], reliability analysis [46, 8], and telecommunications [45]).
+This section reviews NN literature for SA, with more details in [13].
+The first use of NNs in SA was a shallow Feedforward Neural Network (FFNN) to approximate the survival function [14].
+Later, deeper models were developed, such as Cox-nnet [15], a NN for Cox Proportional Hazard (CPH) modeling [11].
+Building on this, DeepSurv [16] introduced a CPH NN with dropout layers for better generalization in personalized treatment.
+Then, mixture models such as DeepWeiSurv [17] further advanced the field by using NNs to regress Weibull mixture distribution parameters.
+These methods enhance modeling capability but do not always improve model calibration and robustness.
+Recent works focus on improving model calibration [18, 19, 20], which ensures predicted confidence matches true likelihood [21].
+However, few works address adversarial robustness in SA, where small input perturbations should not mislead the model [22].
+Although better calibration does not always enhance generalization [23], integrating adversarial robustness improves model trustworthiness in calibration and and stability on unexpected data [24].
+Existing adversarial training methods for SA involve tradeoffs between model complexity, performance, and incorporating physical assumptions.
+Most techniques use Conditional Generative Adversarial Networks (CGANs) to generate time-to-event samples from a survival distribution non-parametrically [25, 26, 27].
+These models implicitly define the survival distribution without assuming a functional/analytical form, offering increased flexibility to fit various data realizations [28, 29, 30].
+However, this flexibility can violate Bayesian Occam’s Razor, which penalizes overly complex models [30].
+To balance complexity and performance, inductive biases are used in explicit parametric models [29].
+One common bias in SA is the memoryless property, satisfied by the exponential distribution [28].
+Thus, this paper employs the exponential CPH model, widely used in SA [31, 32, 33], as well as in queuing theory [34], reliability analysis [3], and telecommunications [35].
 III Preliminaries on Survival Analysis
-SA [25, 6] is a statistical approach used to analyze and model the time $t\in\mathcal{R_{++}}$ until an event $e\in\{0,1\}$ happens (e.g., a disease occurs, a device breaks).
+SA [11] is a statistical approach used to analyze and model the time $t\in\mathcal{R_{++}}$ until an event $e\in\{0,1\}$ happens (e.g., a disease occurs, a device breaks).
 Typically, the study happens over a finite time period and instances/individuals may leave the study early before an event is recorded, with no follow up.
@@
 Such an SA dataset consists of $N$ independent instances/individuals, each denoted by index $i$, for which time between entry to a study and the subsequent event is recorded.
-Upon entry time to a study, the instance’s covariates $\boldsymbol{x}_{i}\in\mathcal{R}^{d}$ (i.e., variables that may affect the event of interest for the study) are measured.
+Upon entry time to a study, each instance’s covariates $\boldsymbol{x}_{i}\in\mathcal{R}^{d}$ (i.e., variables that may affect the event of interest for the study) are measured.
 In summary, the dataset $\mathcal{D}$ has the following structure:
-$\mathcal{D}=\{(\boldsymbol{x_{i}},t_{i},e_{i})\}_{i=1}^{N}.$
+$\mathcal{D}=\{(\boldsymbol{x}_{i},t_{i},e_{i})\}_{i=1}^{N}.$ (1)
 Being a statistical approach, we define the required probabilistic functions to model time-to-event data next.
-III-A Proportional Hazard Model
-Let us denote $T$ as a continuous, non-negative, random variable that represents the time-to-event, with PDF $f(t)$ and Cumulative Distribution Function (CDF) $F(t)=P[T\leq t]$.
+III-A Cox Proportional Hazard Model
+Let us denote $T$ as a continuous, non-negative, random variable that represents the time-to-event, with Probability Density Function (PDF) $f(t)$ and Cumulative Distribution Function (CDF) $F(t)=P[T\leq t]$.
 The survival function $S(t):=1-F(t)$ is the complement of this CDF [37, 11].
@@
 is the instantaneous failure rate at time $t$.
-To incorporate the instance’s covariates, we use the proportional hazard model [49, 7], This model assumes that the covariates increase or decrease the hazard by a proportional amount at all times [48, 6],
+To incorporate the instance’s covariates, we use the prevalent CPH model [2]:
 $\lambda_{\boldsymbol{\theta}}(t,\boldsymbol{x})=\lambda_{0}(t)e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})},$ (4)
 where $\lambda_{0}(t)$ is a baseline hazard, $e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}$ is the relative risk associated with covariates $\boldsymbol{x}$, and $G_{\boldsymbol{\theta}}(\boldsymbol{x})$ is an NN with parameters $\boldsymbol{\theta}$ in our setting.
-The exponential proportional hazard model characterizes the baseline hazard rate as a constant function of time, and therefore $\lambda_{0}(t)=\lambda_{0}$.
-For simplicity, we assume an exponential baseline hazard, and we may “absorb” the $\lambda$ term into the bias of the NN:
+The exponential CPH model characterizes the baseline hazard rate as $\lambda_{0}(t)=\lambda_{0}$.
+For simplicity, we assume an exponential baseline hazard and “absorb” the $\lambda$ term into the bias of the NN:
 $\lambda_{0}e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}=e^{\log\lambda_{0}+G_{\boldsymbol{\theta}}(\boldsymbol{x})}=e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}.$ (5)
-The proportional hazard model PDF, CDF, and complement CDF for an exponential proportional hazard model are then given by:
-$f_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ $=e^{G_{\theta}(\boldsymbol{x})}e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$
-$F_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ $=1.0-e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$
-$S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ $=e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}.$
+The exponential CPH model’s PDF, CDF, and complement CDF are $f_{\boldsymbol{\theta}}(t|\boldsymbol{x})=e^{G_{\theta}(\boldsymbol{x})}e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$, $F_{\boldsymbol{\theta}}(t|\boldsymbol{x})=1.0-e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$ , $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})=e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$, respectively.
 $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ is also known as the instance survival function.
@@
 […]
-Since the distribution of the covariates $p(\boldsymbol{x})$ is typically unknown [50], the integral in Eq. 5 is often replaced with a Monte Carlo estimate [28]:
+Since the distribution of the covariates $p(\boldsymbol{x})$ is typically unknown, the integral in Eq. 6 is often replaced with a Monte Carlo estimate [18]:
 $S(t)$ $\approx\frac{1}{N}\sum_{i=1}^{N}S_{\boldsymbol{\theta}}(t|\boldsymbol{x}_{i}).$ (7)
+Given these definitions, the next section describes the baseline objective for finding model parameters $\boldsymbol{\theta}$ from dataset $\mathcal{D}$.
 III-B Baseline Objective Functions
-Given those definitions, the following loss functions enable finding model parameters $\boldsymbol{\theta}$ from dataset $\mathcal{D}$.
-In particular, we follow [51, 52, 19] to define two objective functions based on negative log likelihood (Section III-B1) and the rank correlation (Section III-B2).
+We follow [38, 25] to define two objective functions based on negative log likelihood (Section III-B1) and the rank correlation (Section III-B2).
 III-B1 Right Censored Log Likelihood Objective
-The $\mathcal{L}_{LL}$ objective is the log of the time-to-event likelihood over $\mathcal{D}$ specified by the proportional hazard model [6]:
+The $\mathcal{L}_{LL}$ objective is the log of the time-to-event likelihood over $\mathcal{D}$ specified by the CPH model [11]:
 $\mathcal{L}_{LL}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})=$ $\sum_{i=1}^{N}e_{i}\cdot\log f_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x_{i}})\ +$
@@
 $\mathcal{L}_{rank}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})$ $=\sum_{i\neq j}A_{i,j}\eta\bigg{(}F_{\theta}(t_{i}|\boldsymbol{x}_{i}),F_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x}_{j})\bigg{)},$ (9)
-where $\eta(\mathbf{x},\mathbf{y})=\exp(-\frac{(\mathbf{x}-\mathbf{y})}{\sigma})$ and $A_{i,j}$ indicates an “acceptable” comparison pair [51, 52, 53]. We set $\sigma=1$ for all experiments.
+where $\eta(\mathbf{x},\mathbf{y})=\exp(-\frac{(\mathbf{x}-\mathbf{y})}{\sigma})$ and $A_{i,j}$ indicates an “acceptable” comparison pair [38, 39]. We set $\sigma=1$ for all experiments.
 III-B3 Combined Objective
-The combined right censored negative log likelihood objective and the risk ranking objective, with risk ranking objective weight trade-off $w=\frac{1}{\text{batch\_size}}$ is
+The combined right censored negative log likelihood objective and the risk ranking objective, with risk ranking objective weight trade-off $w=\frac{1}{\text{batch\_size}}$, is
 $\mathcal{L(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})}=-\mathcal{L}_{LL}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})+w\cdot\mathcal{L}_{rank}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e}).$ (10)
@@
 We will denote the combined loss as the baseline objective.
-Our baseline model’s objective is similar to DRAFT [54] but with a different ranking objective function.
+Our baseline model’s objective is similar to Deep Regularized Accelerated Failure Time (DRAFT) [25] but with a different ranking objective function.
+Thus, we consider the baseline training method for robust SA baseline method as DRAFT.
 Thus far, we have described an objective function that improves predictive performance:
-$\boldsymbol{\theta}^{*}=\arg\min_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})$ (10)
-In the next section, we address how to develop NN models that are robust against adversarial attacks.
+$\boldsymbol{\theta}^{*}=\arg\min_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e}).$ (11)
+In the following sections, we address how to develop NN models that are robust against adversarial perturbations.
 IV Preliminaries on Adversarial Training
-Adversarial training [32, 55, 21, 56] is a technique to improve robustness and generalization of NNs to uncertain data.
-Rather than training on pristine data, adversarial training perturbs the data throughout the training process to encourage the model to optimize for parameters that are resilient against such perturbations.
+Adversarial training [22, 40, 41] improves robustness and generalization of NNs to uncertain data.
+Rather than training on pristine data, most adversarial training approaches perturb the data throughout the training process to encourage the model to optimize for parameters that are resilient against such perturbations.
 For example, [42] formulated adversarial training as solving a Min-Max robust optimization problem:
@@
 Therefore, many recent adversarial training methods propose approximations to Eq.
-11, leading to models with varying degrees of robustness to different data uncertainties.
+12, leading to models with varying degrees of robustness to different data uncertainties.
 To understand the robustness properties of SA models to a variety of data uncertainties present in real applications, this paper investigates several perturbation methods, described next.
@@
 One such approximation is to use techniques from NN verification during training.
-For example, CROWN-IBP combines Convex Relaxation Based Backward Bounding Pass (CROWN) and Interval Bound Propogation (IBP) methods to obtain a convex relaxation for the lower bound and upper bound of each layer in the NN with respect to input data [55, 57].
+For example, CROWN-Interval Bound Propagation (CROWN-IBP) combines CROWN and Interval Bound Propogation (IBP) methods to obtain a convex relaxation for the lower bound and upper bound of each layer in the NN with respect to input data [40, 43].
 CROWN-IBP is commonly used for verifying NN properties over a range of possible inputs.
-[58, 55] showed the linear relaxation may be extended to the entire objective function:
+[44, 40] showed the linear relaxation can be extended to the entire objective function:
 […]
-This relaxation has several benefits: it is convex with respect to the NN weights, which enables efficient optimization, and by optimizing over an upper bound on the inner maximization, this relaxation represents accounting for a strong adversary during training.
-IV-B Adversarial Attacks
+This relaxation is convex with respect to NN weights, enabling efficient optimization, and by optimizing over an upper bound on the inner maximization, it accounts for a strong adversary during training.
+IV-B Adversarial Perturbations
 Another way to approximate Eq.
-19 is to use adversarial attacks.
-Whereas CROWN-IBP provides an upper bound on the inner maximization (i.e., $\max_{\boldsymbol{\tilde{X}}}\mathcal{L}(\theta;\boldsymbol{\tilde{X}})$), adversarial attack algorithms typically provide a lower bound:
-$\max_{\boldsymbol{\tilde{X}}}\ \mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}})\approx\mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}_{FGSM}})$ (13)
+20 is to use adversarial perturbations.
+Whereas CROWN-IBP provides an upper bound on the inner maximization (i.e., $\max_{\boldsymbol{\tilde{X}}}\mathcal{L}(\theta;\boldsymbol{\tilde{X}})$), adversarial perturbation algorithms typically provide a lower bound:
+$\max_{\boldsymbol{\tilde{X}}}\ \mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}})\geq\mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}}_{FGSM}).$ (14)
 Fast Gradient Sign Method (FGSM) [45] uses the gradient of the objective function with respect to the input $\boldsymbol{x}$ to perturb $\boldsymbol{x}$ in the direction that maximizes the objective $\mathcal{L}$.
-Projected Gradient Descent (PGD) [56] iteratively applies that same type of first-order perturbation $K$ times (or until convergence is reached), but constrains the perturbation to $\boldsymbol{x}$ within a specified set (e.g., $\ell_{p}$-ball).
+Projected Gradient Descent (PGD) [42] iteratively applies that same type of first-order perturbation $K$ times (or until convergence is reached), but constrains the perturbation of $\boldsymbol{x}$ to remain within a specified set (e.g., $\ell_{p}$-ball).
 FGSM is a special case of PGD, with $K=1$ and the $\ell_{\infty}$-ball.
 Considering the structure of the datasets as described in Section III, we perturb only the covariates $\boldsymbol{x}$ and not the time-to-event $t$. The perturbed covariates are calculated as:
-$\boldsymbol{x}^{(0)}$ $=\boldsymbol{x}$ (14)
+$\boldsymbol{x}^{(0)}$ $=\boldsymbol{x},$ (15)
 $\boldsymbol{x}^{(k+1)}$ $=\Pi_{\mathcal{B}(\boldsymbol{x},\epsilon)}\left(\boldsymbol{x}^{(k)}+\alpha\big{(}\nabla_{\boldsymbol{x}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{x}^{(k)},\boldsymbol{t},\boldsymbol{e})\big{)}\right),$ (16)
@@
 While robustness to a worst-case perturbation is important for many applications, adding random noise to training data has also been shown to improve generalization performance [46].
-Therefore, this paper also considers Gaussian Noise perturbations with variance $\epsilon$ bounded within an $\ell_{\infty}$ ball:
+Therefore, this paper also considers Gaussian Noise perturbations with variance $\epsilon$ bounded within an $\ell_{\infty}$-ball:
 $\boldsymbol{z}$ $\sim\mathcal{N}(0,\boldsymbol{I}),$ (18)
 $\boldsymbol{\tilde{x}}$ $=\boldsymbol{x}+\Pi_{\mathcal{B}(0,\epsilon)}(\sqrt{\epsilon}\boldsymbol{I}\boldsymbol{z}).$ (19)
-We next introduce our approach, SAWAR, to incorporate adversarial regularization with SA models for improved calibration, predictive generalization, and adversarial robustness.
+Building on this foundation, we next introduce our approach, SAWAR, that develops robust SA models based on adversarial regularization.
 V Survival Analysis with Adversarial Regularization
-As discussed in Section I, a key open challenge in SA is in developing models that are robust to uncertainty and variability in the datasets.
-This paper aims to improve the adversarial robustness of an NN against covariate perturbations within the SA framework.
-In doing so, this paper proposes the first method for enabling adversarial robustness for fully-parametric NN-based SA, which improves calibration and stability on unexpected or noisy data.
-The following sections detail our approach, which extends the baseline objective function in Section III-B3 to include adversarial regularization posed as a Min-Max optimization problem in Section Section IV [21, 55].
+This paper addresses the open challenge in SA of developing models robust to dataset uncertainty and variability, as motivated in Section I.
+We propose a method for adversarial robustness in fully-parametric NN-based SA, enhancing calibration and stability on unexpected or noisy data.
+Our approach extends the objective function (Eq.
+10) with adversarial regularization as a Min-Max optimization problem (Eq.
+12) [41, 40], as detailed in the following sections.
 V-A Min-Max Formulation
@@
 Typically in SA, the aleatoric uncertainty is only associated with the noise of the log of the measured time-to-event (e.g.
-the Gumbel distribution or Log Normal distribution) [16, 6].
-Instead, SAWAR approach assumes that each covariate is subject to a perturbation within a bounded uncertainty set, which leads to a robust optimization formulation as in Section IV.
-Thus, we optimize the NN parameters for the worst-case realizations of the baseline objective in Section III-B3:
+the Gumbel distribution or Log Normal distribution) [11].
+Instead, the proposed SAWAR approach assumes that each covariate is subject to a perturbation within a bounded uncertainty set, which leads to a robust optimization formulation as in Section IV.
+Thus, we optimize the NN parameters for the worst-case realizations of Eq.
+10:
 $\min_{\boldsymbol{\theta}}\max_{\boldsymbol{\tilde{X}}}\ \mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}},\boldsymbol{t},\boldsymbol{e}).$ (20)
-As NNs are highly non-convex, solving this optimization problem is difficult [61].
-Therefore, we use the adversarial training methods discussed in Section IV to approximate/relax the inner maximization term of the Min-Max formulation.
+We use the open-source, PyTorch-based library auto_LiRPA [44] to compute the CROWN-IBP upper bound to the inner maximization term in Eq. 20, which is fully differentiable for backpropogation.
 V-B Adversarial Robustness Regularization
-[21] shows that tight relaxations to the inner maximization of a Min-Max problem may over-regularize the NN, leading to poor predictive performance.
-Therefore, we extend the baseline objective in Eq.
+[41] shows that tight relaxations to the inner maximization of a Min-Max problem can over-regularize the NN, leading to poor predictive performance.
+Therefore, we balance the objective in Eq.
 10 with the adversarial robustness objective in Eq.
-19 [55, 57] to balance predictive performance and adversarial robustness as:
-$\min_{\boldsymbol{\theta}}\left[\kappa\cdot\mathcal{L}+(1-\kappa)\cdot\max_{\boldsymbol{\tilde{X}}}\mathcal{L}\right],$ (20)
+20 [40, 43] to combine predictive performance and adversarial robustness as:
+$\min_{\boldsymbol{\theta}}\left[\kappa\cdot\mathcal{L}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})+(1-\kappa)\cdot\max_{\boldsymbol{\tilde{X}}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}},\boldsymbol{t},\boldsymbol{e})\right],$ (21)
 where $\kappa\in(0,1)$.
-We propose to use the open-source library auto_LiRPA [58] to relax the inner maximization term of the adversarial regularization with CROWN-IBP.
-auto_LiRPA seamlessly integrates with PyTorch for automatic backpropagation, computing a fully differentiable upper bound on the inner maximization term of Eq.
-20 in a computationally feasible manner.
 V-C SAWAR Training
-We show the pseudo-code in Algorithm 1.
-The NN parameters are updated using stochastic gradient descent with the ADAM optimizer [62] (12), where we use early stopping [63] with respect to the validation set to prevent overfitting (25, 26 and 27).
-We introduce the covariate perturbation attacks after 100 epochs (4, 5, 6 and 7).
-A smooth $\epsilon$-scheduler linearly increases the $\epsilon$-perturbation magnitude from $0.0$ to $0.5$ within a 30 epoch window (28) after 100 epochs.
+The NN parameters are updated using stochastic gradient descent with the ADAM optimizer [47], where we use early stopping [48] with respect to the validation set to prevent overfitting We introduce the covariate perturbations after 100 epochs.
+A smooth $\epsilon$-scheduler linearly increases the $\epsilon$-perturbation magnitude from $0.0$ to $0.5$ within a 30 epoch window after 100 epochs.
 We do not allow early stopping until after the maximum $\epsilon$-perturbation magnitude is achieved.
-Algorithm 1 Robust Training with CROWN-IBP
-1:procedure train_step(ModelLoss, Data, $\epsilon$, $\kappa$, Train)
-2:   $loss_{epoch}\leftarrow 0$
-3:   for $x_{B},t_{B},y_{B}$ in Data do
-4:      if $\epsilon>1e^{-20}$ then
-5:         lossnatural $\leftarrow\text{ModelLoss}(x_{B},t_{B},e_{B})$
-6:         $\tilde{x}_{B}\leftarrow\text{PerturbationLpNorm}(x_{B},\epsilon)$
-[…]
-8:         loss $\leftarrow\kappa\times loss_{natural}+(1-\kappa)\times loss_{robust}$
-9:      else
-10:         loss $\leftarrow\text{ModelLoss}(x_{B},t_{B},e_{B})$
-11:      if Train then
-12:         ADAM Optimizer Step with loss.backward()
-13:      lossepoch $\leftarrow$ lossepoch + loss
-14:   return lossepoch
-15:
-16:procedure train_robust(Model, Datatr, Dataval, $\kappa$)
-17:   Init random seeds, ADAM optimizer, lr-scheduler, $\epsilon$-scheduler
-18:   ModelLoss = Losswrapper(Model)
-[…]
-20:   for $t$ in Epochs do
-21:      if $\epsilon$-scheduler reached $\epsilon_{max}$ then
-22:         lr-scheduler Step
-[…]
-[…]
-[…]
-26:         Save Best Model
-27:         lossbest $\leftarrow$ lossval
-28:      $\epsilon$-scheduler Step
-29:   return Model
 We expect that the model’s adversarial robustness and predictive calibration should increase when incorporating data uncertainty in SAWAR objective [24] via the adversarial regularization term in Eq.
-20 and illustrate SAWAR hypothesis via extensive experiments in the next section.
+21 and show SAWAR’s improved performance via extensive experiments in the next section.
 VI Experiments
-We conduct experiments over 10 benchmark SA datasets (VI-B) to quantify the impact of perturbation after the robust training is complete.
+We conduct experiments over 10 benchmark medical SA datasets (VI-B) to quantify the impact of perturbation after the robust training is complete.
 As evaluation metrics, we consider the Concordance Index (CI) [49], the Integrated Brier Score (IBS) [50] and the Negative Log Likelihood (NegLL) at varying $\epsilon$-perturbation magnitudes from 0 to 1.
-We show that SAWAR improves predictive accuracy, calibration, and population curve estimation.
+To demonstrate that SAWAR produces SA models that perform well and are robust to dataset uncertainty, this section quantifies the improvement in predictive accuracy, calibration, and population curve estimation.
 VI-A Configuration
-The experiments are conducted on a 12th Gen Intel(R) Core (TM) 16 GB RAM i9-12900 processor running a Windows operating system.
+The experiments are conducted on a Windows computer with a 12th Gen Intel(R) Core (TM) i9-12900 processor and 16 GB RAM.
 Each dataset undergoes stratified partitioning into train, validation, test sets with proportions 60%, 20%, 20% respectively.
-The NN $G_{\boldsymbol{\theta}}$ has two hidden layers of 50 neurons each and Leaky ReLU activation functions.
-lifelines [66], and sksurv [53] software packages are employed to ensure standardization and reproducibility in computing the SA metrics.
+The NN $G_{\boldsymbol{\theta}}$ has 2 hidden layers of 50 neurons each and Leaky ReLU activation functions.
+We use lifelines [51] and sksurv [39] for computing SA metrics.
 VI-B Datasets
-We use the open-source API SurvSet [67], to download the benchmark SA datasets shown in Table I.
-Dataset
-$N$
-$n_{fac}$
-$n_{ohe}$
-$n_{num}$
-TRACE
-1878
-4
-4
-2
-stagec
-146
-4
-15
-3
-flchain
-7874
-4
-26
-6
-Aids2
-2839
-3
-11
-1
-Framingham
-4699
-2
-12
-5
-dataDIVAT1
-5943
-3
-14
-2
-prostate
-502
-6
-16
-9
-zinc
-431
-11
-18
-2
-retinopathy
-394
-5
-9
-2
-LeukSurv
-1043
-2
-24
-5
+We use the API SurvSet [52] to download the benchmark medical SA datasets shown in Table I and described below:
+Dataset $N$ $n_{fac}$ $n_{ohe}$ $n_{num}$
+TRACE 1878 4 4 2
+stagec 146 4 15 3
+flchain 7874 4 26 6
+Aids2 2839 3 11 1
+Framingham 4699 2 12 5
+dataDIVAT1 5943 3 14 2
+prostate 502 6 16 9
+zinc 431 11 18 2
+retinopathy 394 5 9 2
+LeukSurv 1043 2 24 5
 TABLE I: Studied datasets from SurvSet. $n_{fac}$ is the number of categorical features, $n_{ohe}$ is number of binary features (one-hot-encoded) and $n_{num}$ is number of numerical features.
@@
     The LeukSurv [62] is from the study on survival of acute myeloid leukemia and connection to spatial variation in survival.
-We preprocess each dataset by one-hot encoding categorical covariates, followed by standard normalizing all covariates.
-Figure 1: Distribution of survival curves for baseline, noise, FGSM, PGD, and CROWN-IBP.
+We preprocess each dataset by one-hot encoding categorical covariates and standard normalization of all covariates.
+Figure 1: Distribution of survival curves for DRAFT, FGSM, PGD, and CROWN-IBP.
 As the strength of the adversarial training method increases, from left to right column, the 90% credible interval narrows.
-However, the credible interval of SAWAR does not drastically differ from the baseline.
+However, the credible interval of SAWAR does not drastically differ from the DRAFT.
+Figure 2: Adversarial robustness of survival curves for DRAFT, FGSM, PGD, and CROWN-IBP training against $\epsilon$-worst-case perturbation.
+As the strength of the adversarial training method increases, from left to right column, the perturbed population survival curve for a given $\epsilon$ is closer to the KMC.
 VI-C Adversary Setup
-A detailed insight into the robust evaluation procedure is shown in Algorithm 2.
-We use two adversarial attack methods when benchmarking the robustness of the adversarial regularization method (14, 15 and 19): FGSM attack (Section IV-B) and a novel worst-case attack.
-We define the worst-case attack as the maximum failure rate with respect to the covariate uncertainty:
-$\tilde{\lambda}_{\theta}=\max_{\boldsymbol{\tilde{x}_{i}}\in\mathcal{B}(\boldsymbol{x_{i}},\epsilon)}e^{G_{\boldsymbol{\theta}}(\boldsymbol{\tilde{x}}_{i})},$ (21)
-and define the $\epsilon$-worst-case survival curve as:
-$\tilde{S}(t|\boldsymbol{x})=e^{-\tilde{\lambda}_{\theta}t},$ (22)
-which is the lowest survival probability at a given time $t$ with respect to the $\epsilon$-perturbed input covariates.
-For the exponential distribution, as the failure rate increases with the $\epsilon$-perturbation, the probability of survival decreases monotonically.
-We use CROWN-IBP for a tight approximation to the upper bound on the maximization term in equation Eq.
-21.
-We evaluate the SA metrics for $\epsilon\in[0,1]$ for the two attacks methods to study how sensitive each adversarial training method is to covariate perturbation (1).
-Note that when the activation functions are piecewise linear, an exact solution can also be obtained by solving a Mixed Integer Linear Program (MILP).
-However, solving the MILP is computationally prohibitive: even for $\epsilon=0.1$ each MILP takes on average 9-10 seconds when tested on the Dialysis dataset, whereas CROWN-IBP takes milliseconds to calculate over all the data in the Dialysis dataset.
-Algorithm 2 Adversarial Robustness Evaluation
-1:procedure eval_robust(Model, Data, Metric, Method, $\epsilon$)
-2:   $X,t,e$ = Data
-3:   MetricValues=[]
-4:   for $\epsilon$ in $0:0.1:1$ do
-5:      if $\epsilon=0$ then
-6:         $\lambda_{\theta}\leftarrow$ Model(X)
-7:         MetricValues = [MetricValues, $\text{Metric}(\lambda_{\theta},t,e)$]
-8:      else
-9:         $\tilde{\lambda}_{\theta}\leftarrow$ ATTACK(Model,Data,Method,$\epsilon$)
-10:         MetricValues = [MetricValues, $\text{Metric}(\tilde{\lambda}_{\theta},t,e)$]
-11:   return MetricValues
-12:
-13:procedure attack(Model, Data, Method, $\epsilon$)
-14:   $X,t,e$ = Data
-15:   if Method = FGSM then
-16:      ModelLoss = Losswrapper(Model)
-17:      $\tilde{X}\leftarrow$ FGSM(ModelLoss, $X,t,e,\epsilon$)
-18:      $\tilde{\lambda}_{\theta}\leftarrow$ Model($\tilde{X}$)
-19:   if Method = worst-case then
-20:      $\tilde{X}\leftarrow\text{PerturbationLpNorm}(X,\epsilon)$
-21:      $\tilde{\lambda}_{\theta}\leftarrow\text{CROWN-IBP}(\text{Model})(\tilde{X},t,e)$
-22:   return $\tilde{\lambda}_{\theta}$
-VI-D Results
+We apply two adversarial perturbation settings on the test set to assess the robustness of the proposed adversarial regularization method: FGSM perturbation (Section IV-B) and a novel worst-case perturbation.
+FGSM is applied by using the gradient of the objective function with respect to the input $\boldsymbol{x}$ to perturb $\boldsymbol{x}$ in the direction that maximizes the objective Eq.
+20, such that:
 $\epsilon$ Concordance Index Integrated Brier Score Negative Log Likelihood
-baseline noise fgsm pgd SAWAR baseline noise fgsm pgd SAWAR baseline noise fgsm pgd SAWAR
+DRAFT Noise FGSM PGD SAWAR DRAFT Noise FGSM PGD SAWAR DRAFT Noise FGSM PGD SAWAR
 1.00 2.9 3.35 3.35 2.5 2.9 4.6 4.4 2.4 1.9 1.7 4.0 4.7 2.7 2.1 1.5
@@
 $\epsilon$ Concordance Index Integrated Brier Score Negative Log Likelihood
-baseline noise fgsm pgd SAWAR baseline noise fgsm pgd SAWAR baseline noise fgsm pgd SAWAR
+DRAFT Noise FGSM PGD SAWAR DRAFT Noise FGSM PGD SAWAR DRAFT Noise FGSM PGD SAWAR
 1.00 2.8 3.4 3.5 3.6 1.7 2.6 4.3 3.8 3.3 1.0 3.1 4.5 3.4 3.0 1.0
@@
 TABLE III: Worst-Case average ranks for CI, IBS, and NegLL metrics across all datasets. Lower number is better.
-Furthermore, we analyze how these adversarial attacks impact the characteristics of the population survival curves.
+$\tilde{\lambda}_{\boldsymbol{\theta}}=e^{G_{\boldsymbol{\theta}}\left(\text{FGSM}(\mathcal{L},\boldsymbol{x})\right)}.$ (22)
+We define the worst-case perturbation as the maximum failure rate with respect to the covariate uncertainty:
+$\tilde{\lambda}_{\boldsymbol{\theta}}=\max_{\boldsymbol{\tilde{x}_{i}}\in\mathcal{B}(\boldsymbol{x_{i}},\epsilon)}e^{G_{\boldsymbol{\theta}}(\boldsymbol{\tilde{x}}_{i})},$ (23)
+and the $\epsilon$-worst-case survival curve as the lowest survival probability at a given time $t$ with respect to the $\epsilon$-perturbed input covariates:
+$\tilde{S}(t|\boldsymbol{x})=e^{-\tilde{\lambda}_{\boldsymbol{\theta}}t}.$ (24)
+For the exponential CPH distribution, as the failure rate increases with the $\epsilon$-perturbation, the probability of survival decreases monotonically.
+We use CROWN-IBP for a tight approximation to the upper bound on the maximization term in equation Eq.
+23.
+We evaluate the SA metrics for $\epsilon\in[0,1]$, where $\epsilon=0$ is no perturbation to input covariates, for the two perturbations methods to study how sensitive each adversarial training method is to covariate perturbation.
+Note that when the activation functions are piecewise linear, an exact solution can also be obtained by solving a Mixed Integer Linear Program (MILP).
+However, solving the MILP is computationally prohibitive: even for $\epsilon=0.1$ each MILP takes on average 9-10 seconds when tested on the Dialysis dataset, whereas CROWN-IBP takes milliseconds to calculate over all the data in the Dialysis dataset.
+VI-D Results
+Furthermore, we analyze how adversarial perturbations impact the characteristics of the population survival curves.
 VI-D1 Failure Rate Distribution
-Let us assume that the input covariates are random variables.
-Then the failure rate becomes a random variable that occurs from the stochasticity of the input covariates.
-Moreover, we may may treat $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ as a stochastic process:
-$\boldsymbol{x}\sim p(\boldsymbol{x});\lambda\sim\lambda_{\boldsymbol{\theta}}(\boldsymbol{x})$ (23)
+Let us assume that the input covariates are random variables $\boldsymbol{x}\sim p(\boldsymbol{x})$.
+Then the failure rate becomes a random variable that occurs from the stochasticity of the input covariates $\lambda\sim\lambda_{\boldsymbol{\theta}}(\boldsymbol{x})$.
+Moreover, we treat $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ as a stochastic process :
 $S_{\boldsymbol{\theta}}(t)\sim e^{-\lambda t}$ (25)
-Furthermore, we compute statistical quantities such as expectation, lower 5th quantile, and upper 95th quantile:
-$LB=Q_{05}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x})),$ (25)
-$UB=Q_{95}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x})).$ (26)
+Furthermore, we compute statistical quantities such as expectation, lower 5th quantile $LB=Q_{05}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x}))$, and upper 95th quantile $UB=Q_{95}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x}))$.
 Fig.
 1 visualizes the credible interval and mean of the survival function stochastic process in Eq.
-24.
+25.
 The 90% credible interval of the the stochastic process changes with respect to the strength of the adversarial regularization method (Fig.
@@
 1).
-The variance of the failure rate decreasing leads to a more robust model against adversarial attacks.
+The variance of the failure rate decreasing leads to a more robust model against adversarial perturbations.
 For example, when $\mathrm{Var}(\lambda_{\theta}(\boldsymbol{x}))=0$, then there is a mode collapse such that $\lambda(\boldsymbol{x}_{i})=\lambda\textrm{ }\forall i$.
@@
 Therefore, by utilizing a weighted objective in Eq.
-20, we trade-off individualized survival curves with vulnerability to adversarial attacks and a population survival curve which is completely resilient to adversarial attacks.
-Figure 2: Adversarial robustness of survival curves for baseline, noise, FGSM, PGD, and CROWN-IBP training against $\epsilon$-worst-case attack.
-As the strength of the adversarial training method increases, from left to right column, the perturbed population survival curve for a given $\epsilon$ is closer to the KMC.
-However, despite exhibiting significantly higher adversarial robustness, the credible interval remains relatively unchanged compared to the baseline method.
+21, we trade-off individualized survival curves with vulnerability to adversarial perturbations and a population survival curve which is completely resilient to adversarial perturbations.
+However, despite exhibiting significantly higher adversarial robustness, the credible interval usually remains relatively unchanged compared to the DRAFT method.
 VI-D2 Population Survival Curve
 The $\epsilon$-perturbation magnitude determines the severity of the worst-case survival curve.
-We observe the trend that the stronger the adversarial regularization method, the ”closer” the $\epsilon$-worst-case population curve becomes to the unperturbed population survival curve for a given $\epsilon\in[0,1]$ (Fig.
+We observe the trend that the stronger the adversarial regularization method, the “closer” the $\epsilon$-worst-case population curve becomes to the unperturbed population survival curve for a given $\epsilon\in[0,1]$ (Fig.
 2).
@@
 2 demonstrates the strong calibration of SAWAR model [19].
-We calculate the average ranking of each adversarial regularization method across all datasets with respect to CI, IBS, NegLL in response to the FGSM and $\epsilon$-worst-case attacks (Tables III and II).
+We calculate the average ranking of each adversarial regularization method across all datasets with respect to CI, IBS, NegLL in response to the FGSM and $\epsilon$-worst-case perturbations (Tables III and II).
 We empirically showed that CROWN-IBP adversarial regularization leads to increased performance, consistently ranking the best across various $\epsilon$-perturbation magnitudes.
-The exact corresponding metric values for each $\epsilon$ of the worst-case attack are provided in Appendix A.
+More detailed results, along with the exact metric values for each $\epsilon$ of the worst-case perturbation for each dataset are available in the code repository linked in the abstract.
 VII Conclusions
-This paper is the first to introduce CROWN-IBP regularization for adversarial robustness training in SA with fully-parametric NN models.
-We benchmarked SAWAR method on 10 time-to-event datasets with standard adversarial regularization methods such as Gaussian Noise, FGSM, and PGD against the FGSM and worst-case attack.
-SAWAR’s empirical results indicate that CROWN-IBP regularization consistently enhances performance in SA with respect to CI and IBS, and NegLL for all $\epsilon$-perturbation magnitudes from 0 to 1.
-Moreover, compared to conventional adversarial regularization methods, SAWAR empirical evidence demonstrates that CROWN-IBP regularization is more resilient to input covariate perturbation attacks, where we see relative improvements to the baseline in CI and IBS of 1%-150% across all $\epsilon$-perturbation magnitudes (Appendix A).
-Thus adversarial regularization improves predictive performance, adversarial robustness, and calibration.
-Future work will include the use the Weibull Proportional Hazard model to enable time varying hazard rates.
+This paper introduces CROWN-IBP regularization for enhancing adversarial robustness in fully-parametric SA using NN models.
+We evaluated the SAWAR method on 10 time-to-event medical datasets, comparing it with standard adversarial regularization methods such as Gaussian noise, FGSM, and PGD, as well as the novel worst-case perturbation.
+Our empirical results show that CROWN-IBP regularization consistently improves performance in SA across various metrics, including CI, IBS, and NegLL, for all $\epsilon$-perturbation magnitudes from 0 to 1.
+Compared to conventional adversarial regularization techniques, SAWAR demonstrates superior resilience to input covariate perturbations, with relative improvements in CI and IBS ranging from 1% to 150% across all $\epsilon$-perturbation magnitudes.
+In summary, SAWAR significantly enhances predictive performance, adversarial robustness, and calibration.
+This approach effectively mitigates data uncertainty in SA and improves generalization across unseen datasets.
+Future work will include the use the Weibull CPH model to enable time varying hazard rates.
 This will require extending auto_LiRPA to handle the power operator for operations such as $t^{k}$, where $k$ is a learnable parameter.
 Appendix A Appendix
-We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1, the relative percentage change from the baseline to the adversarial training methods becomes larger and then smaller.
-The relative percent changes from the baseline to SAWAR for $\epsilon$ equal to 1.0, 0.9, 0.8, 0.7, 0.6, 0.5, 0.4, 0.3, 0.2, 0.1, 0.05, 0 is 29.46, 48.18, 56.27, 72.23, 102.44, 133.35, 152.17, 140.33, 91.78, 21.17, 9.63, 1.07 respectively.
+We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller.
+The relative percent changes in CI from the DRAFT training objective to SAWAR training objective is shown in Table IV (where higher percentage change is better).
 We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
 $\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
-Aids2 baseline 0.499 0.497 0.501 0.505 0.507 0.509 0.517 0.535 0.565 0.573 0.573 0.572
+%$\%\Delta$ 29.46 48.18 56.27 72.23 102.44 133.35 152.17 140.33 91.78 21.17 9.63 1.07
+TABLE IV: The relative percent change in Concordance Index metric from the DRAFT training objective to the SAWAR training objective averaged across the SurvSet datasets.
+A higher relative percent change is better.
+$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
+Aids2 DRAFT 0.499 0.497 0.501 0.505 0.507 0.509 0.517 0.535 0.565 0.573 0.573 0.572
 Noise 0.492 0.494 0.496 0.5 0.5 0.498 0.507 0.526 0.556 0.556 0.554 0.554
@@
 SAWAR 0.567 0.575 0.577 0.579 0.58 0.58 0.579 0.579 0.579 0.579 0.579 0.579
-Framingham baseline 0.612 0.623 0.635 0.646 0.656 0.664 0.672 0.678 0.682 0.686 0.687 0.688
+Framingham DRAFT 0.612 0.623 0.635 0.646 0.656 0.664 0.672 0.678 0.682 0.686 0.687 0.688
 Noise 0.606 0.616 0.626 0.635 0.642 0.649 0.659 0.674 0.683 0.685 0.685 0.685
@@
 SAWAR 0.694 0.697 0.699 0.701 0.702 0.704 0.704 0.705 0.705 0.704 0.704 0.704
-LeukSurv baseline 0.611 0.616 0.623 0.623 0.627 0.631 0.63 0.631 0.63 0.624 0.624 0.626
+LeukSurv DRAFT 0.611 0.616 0.623 0.623 0.627 0.631 0.63 0.631 0.63 0.624 0.624 0.626
 Noise 0.552 0.557 0.562 0.566 0.569 0.574 0.587 0.61 0.628 0.627 0.621 0.617
@@
 SAWAR 0.483 0.498 0.51 0.532 0.553 0.583 0.589 0.607 0.617 0.623 0.632 0.636
-TRACE baseline 0.564 0.586 0.612 0.637 0.665 0.69 0.714 0.73 0.734 0.733 0.735 0.735
+TRACE DRAFT 0.564 0.586 0.612 0.637 0.665 0.69 0.714 0.73 0.734 0.733 0.735 0.735
 Noise 0.565 0.588 0.613 0.64 0.668 0.692 0.716 0.73 0.733 0.735 0.734 0.735
@@
 SAWAR 0.73 0.733 0.736 0.738 0.739 0.739 0.74 0.74 0.739 0.738 0.738 0.737
-dataDIVAT1 baseline 0.534 0.558 0.583 0.606 0.624 0.639 0.65 0.658 0.662 0.664 0.664 0.664
+dataDIVAT1 DRAFT 0.534 0.558 0.583 0.606 0.624 0.639 0.65 0.658 0.662 0.664 0.664 0.664
 Noise 0.534 0.539 0.551 0.571 0.587 0.612 0.636 0.656 0.654 0.64 0.635 0.631
@@
 SAWAR 0.627 0.642 0.652 0.658 0.659 0.659 0.659 0.658 0.657 0.656 0.656 0.655
-flchain baseline 0.096 0.098 0.102 0.107 0.114 0.127 0.156 0.461 0.9 0.916 0.922 0.924
+flchain DRAFT 0.096 0.098 0.102 0.107 0.114 0.127 0.156 0.461 0.9 0.916 0.922 0.924
 Noise 0.096 0.099 0.104 0.112 0.127 0.164 0.335 0.768 0.912 0.921 0.923 0.923
@@
 SAWAR 0.84 0.925 0.927 0.929 0.93 0.93 0.929 0.929 0.93 0.93 0.931 0.931
-prostate baseline 0.412 0.419 0.426 0.447 0.463 0.49 0.53 0.581 0.614 0.635 0.639 0.643
+prostate DRAFT 0.412 0.419 0.426 0.447 0.463 0.49 0.53 0.581 0.614 0.635 0.639 0.643
 Noise 0.49 0.503 0.514 0.524 0.533 0.541 0.553 0.566 0.576 0.572 0.577 0.576
@@
 SAWAR 0.575 0.584 0.598 0.6 0.608 0.632 0.638 0.638 0.639 0.641 0.642 0.644
-retinopathy baseline 0.525 0.539 0.562 0.569 0.587 0.592 0.592 0.599 0.605 0.601 0.604 0.605
+retinopathy DRAFT 0.525 0.539 0.562 0.569 0.587 0.592 0.592 0.599 0.605 0.601 0.604 0.605
 Noise 0.526 0.545 0.561 0.57 0.581 0.588 0.588 0.6 0.603 0.6 0.604 0.606
@@
 SAWAR 0.588 0.597 0.611 0.616 0.619 0.615 0.621 0.62 0.62 0.619 0.617 0.617
-stagec baseline 0.442 0.457 0.476 0.495 0.514 0.548 0.582 0.668 0.721 0.697 0.688 0.688
+stagec DRAFT 0.442 0.457 0.476 0.495 0.514 0.548 0.582 0.668 0.721 0.697 0.688 0.688
 Noise 0.471 0.481 0.505 0.49 0.514 0.534 0.582 0.697 0.721 0.712 0.688 0.697
@@
 SAWAR 0.428 0.447 0.481 0.514 0.567 0.639 0.673 0.692 0.702 0.692 0.692 0.707
-zinc baseline 0.339 0.349 0.356 0.376 0.427 0.507 0.657 0.758 0.796 0.796 0.792 0.793
+zinc DRAFT 0.339 0.349 0.356 0.376 0.427 0.507 0.657 0.758 0.796 0.796 0.792 0.793
 Noise 0.315 0.329 0.339 0.376 0.424 0.545 0.704 0.778 0.791 0.805 0.806 0.81
@@
 SAWAR 0.383 0.464 0.533 0.649 0.707 0.746 0.766 0.779 0.789 0.787 0.795 0.8
-TABLE IV: Concordance Index metric for SurvSet datasets (higher is better) for each adversarial training method against the worst-case adversarial attack.
-We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1, the relative percentage change from the baseline to the adversarial training methods becomes larger and then smaller.
-The relative percent changes from the baseline to SAWAR for $\epsilon$ equal to 1.0, 0.9, 0.8, 0.7, 0.6, 0.5, 0.4, 0.3, 0.2, 0.1, 0.05, 0 is -44.13, -45.46, -46.68, -47.57, -47.86, -47.10, -45.03, -40.89, -33.06, -20.29, -11.34, -1.68 respectively (where lower percentage change is better).
+TABLE V: Concordance Index metric for SurvSet datasets (higher is better) for each adversarial training method against the worst-case adversarial attack.
+We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller.
+The relative percent changes in Integrated Brier Score metric from the DRAFT training objective to SAWAR training objective is shown in Table VI (where lower percentage change is better).
 We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
 $\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
-Aids2 baseline 0.24 0.233 0.224 0.212 0.197 0.178 0.157 0.137 0.124 0.12 0.121 0.122
+%$\%\Delta$44.13 -45.46 -46.68 -47.57 -47.86 -47.10 -45.03 -40.89 -33.06 -20.29 -11.34 -1.68
+TABLE VI: The relative percent change in Integrated Brier Score metric from the DRAFT training objective to the SAWAR training objective averaged across the SurvSet datasets.
+A lower relative percent change is better.
+$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
+Aids2 DRAFT 0.24 0.233 0.224 0.212 0.197 0.178 0.157 0.137 0.124 0.12 0.121 0.122
 Noise 0.254 0.253 0.25 0.245 0.236 0.22 0.196 0.161 0.13 0.121 0.121 0.122
@@
 SAWAR 0.126 0.124 0.124 0.124 0.123 0.123 0.123 0.123 0.123 0.123 0.123 0.124
-Framingham baseline 0.81 0.8 0.778 0.733 0.65 0.52 0.369 0.243 0.165 0.128 0.119 0.115
+Framingham DRAFT 0.81 0.8 0.778 0.733 0.65 0.52 0.369 0.243 0.165 0.128 0.119 0.115
 Noise 0.818 0.818 0.816 0.809 0.778 0.681 0.495 0.294 0.174 0.128 0.119 0.116
@@
 SAWAR 0.165 0.152 0.141 0.133 0.127 0.122 0.118 0.116 0.114 0.114 0.113 0.113
-LeukSurv baseline 0.163 0.163 0.163 0.163 0.162 0.161 0.156 0.144 0.126 0.109 0.105 0.103
+LeukSurv DRAFT 0.163 0.163 0.163 0.163 0.162 0.161 0.156 0.144 0.126 0.109 0.105 0.103
 Noise 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.157 0.134 0.12 0.114
@@
 SAWAR 0.154 0.15 0.144 0.136 0.128 0.12 0.115 0.112 0.11 0.107 0.107 0.106
-TRACE baseline 0.634 0.634 0.633 0.626 0.603 0.549 0.449 0.329 0.234 0.183 0.172 0.167
+TRACE DRAFT 0.634 0.634 0.633 0.626 0.603 0.549 0.449 0.329 0.234 0.183 0.172 0.167
 Noise 0.634 0.634 0.632 0.622 0.596 0.536 0.435 0.319 0.23 0.183 0.172 0.167
@@
 SAWAR 0.282 0.254 0.231 0.213 0.198 0.186 0.177 0.17 0.166 0.163 0.163 0.163
-dataDIVAT1 baseline 0.721 0.702 0.67 0.616 0.531 0.417 0.308 0.234 0.195 0.18 0.178 0.178
+dataDIVAT1 DRAFT 0.721 0.702 0.67 0.616 0.531 0.417 0.308 0.234 0.195 0.18 0.178 0.178
 Noise 0.749 0.749 0.748 0.748 0.743 0.72 0.626 0.413 0.248 0.198 0.192 0.192
@@
 SAWAR 0.266 0.241 0.224 0.212 0.202 0.193 0.187 0.183 0.18 0.178 0.178 0.178
-flchain baseline 0.831 0.831 0.831 0.831 0.831 0.83 0.829 0.818 0.486 0.096 0.07 0.052
+flchain DRAFT 0.831 0.831 0.831 0.831 0.831 0.83 0.829 0.818 0.486 0.096 0.07 0.052
 Noise 0.831 0.831 0.831 0.831 0.831 0.83 0.829 0.807 0.25 0.087 0.063 0.052
@@
 SAWAR 0.522 0.161 0.076 0.063 0.057 0.054 0.052 0.051 0.05 0.05 0.049 0.049
-prostate baseline 0.53 0.529 0.528 0.526 0.519 0.505 0.475 0.418 0.335 0.256 0.231 0.218
+prostate DRAFT 0.53 0.529 0.528 0.526 0.519 0.505 0.475 0.418 0.335 0.256 0.231 0.218
 Noise 0.53 0.53 0.53 0.53 0.53 0.53 0.53 0.528 0.477 0.371 0.325 0.296
@@
 SAWAR 0.508 0.482 0.441 0.393 0.348 0.309 0.278 0.251 0.229 0.214 0.209 0.205
-retinopathy baseline 0.662 0.657 0.645 0.621 0.582 0.522 0.43 0.324 0.244 0.203 0.196 0.195
+retinopathy DRAFT 0.662 0.657 0.645 0.621 0.582 0.522 0.43 0.324 0.244 0.203 0.196 0.195
 Noise 0.662 0.657 0.645 0.622 0.583 0.523 0.432 0.326 0.245 0.204 0.196 0.195
@@
 SAWAR 0.526 0.484 0.436 0.385 0.334 0.29 0.254 0.228 0.211 0.203 0.202 0.202
-stagec baseline 0.505 0.505 0.505 0.504 0.502 0.486 0.431 0.329 0.211 0.175 0.181 0.192
+stagec DRAFT 0.505 0.505 0.505 0.504 0.502 0.486 0.431 0.329 0.211 0.175 0.181 0.192
 Noise 0.505 0.505 0.505 0.504 0.502 0.485 0.432 0.338 0.216 0.175 0.181 0.192
@@
 SAWAR 0.491 0.476 0.45 0.408 0.354 0.296 0.24 0.195 0.168 0.16 0.162 0.167
-zinc baseline 0.907 0.907 0.905 0.899 0.882 0.823 0.652 0.368 0.18 0.109 0.096 0.09
+zinc DRAFT 0.907 0.907 0.905 0.899 0.882 0.823 0.652 0.368 0.18 0.109 0.096 0.09
 Noise 0.908 0.907 0.907 0.904 0.893 0.838 0.625 0.32 0.168 0.109 0.096 0.089
@@
 SAWAR 0.72 0.609 0.476 0.352 0.26 0.198 0.159 0.132 0.113 0.1 0.096 0.094
-TABLE V: Integrated Brier Score metric for SurvSet datasets (lower is better) for each adversarial training method against the worst-case adversarial attack.
+TABLE VII: Integrated Brier Score metric for SurvSet datasets (lower is better) for each adversarial training method against the worst-case adversarial attack.
 $\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
-Aids2 baseline 1.14e+04 7.22e+03 4.60e+03 2.87e+03 1.80e+03 1.16e+03 7.95e+02 6.06e+02 5.32e+02 5.16e+02 5.17e+02 5.19e+02
+Aids2 DRAFT 1.14e+04 7.22e+03 4.60e+03 2.87e+03 1.80e+03 1.16e+03 7.95e+02 6.06e+02 5.32e+02 5.16e+02 5.17e+02 5.19e+02
 Noise 3.84e+05 1.61e+05 6.68e+04 2.78e+04 1.14e+04 4.42e+03 1.83e+03 8.54e+02 5.64e+02 5.20e+02 5.19e+02 5.21e+02
@@
 SAWAR 5.29e+02 5.24e+02 5.22e+02 5.22e+02 5.21e+02 5.21e+02 5.21e+02 5.21e+02 5.21e+02 5.22e+02 5.22e+02 5.22e+02
-Framingham baseline 1.47e+05 7.07e+04 3.42e+04 1.68e+04 8.47e+03 4.57e+03 2.79e+03 2.01e+03 1.68e+03 1.54e+03 1.51e+03 1.50e+03
+Framingham DRAFT 1.47e+05 7.07e+04 3.42e+04 1.68e+04 8.47e+03 4.57e+03 2.79e+03 2.01e+03 1.68e+03 1.54e+03 1.51e+03 1.50e+03
 Noise 1.27e+07 3.17e+06 8.03e+05 2.06e+05 5.38e+04 1.48e+04 4.92e+03 2.39e+03 1.72e+03 1.54e+03 1.51e+03 1.50e+03
@@
 SAWAR 1.66e+03 1.62e+03 1.58e+03 1.56e+03 1.53e+03 1.52e+03 1.50e+03 1.50e+03 1.49e+03 1.49e+03 1.49e+03 1.49e+03
-LeukSurv baseline 1.79e+07 4.80e+06 1.29e+06 3.48e+05 9.29e+04 2.53e+04 6.59e+03 1.77e+03 5.76e+02 3.10e+02 2.74e+02 2.58e+02
+LeukSurv DRAFT 1.79e+07 4.80e+06 1.29e+06 3.48e+05 9.29e+04 2.53e+04 6.59e+03 1.77e+03 5.76e+02 3.10e+02 2.74e+02 2.58e+02
 Noise 5.46e+23 1.79e+21 7.98e+18 2.52e+16 8.53e+13 1.88e+11 5.25e+08 2.35e+06 3.37e+04 2.60e+03 1.21e+03 7.21e+02
@@
 SAWAR 2.07e+04 8.38e+03 3.50e+03 1.45e+03 6.62e+02 3.45e+02 2.91e+02 2.70e+02 2.60e+02 2.54e+02 2.52e+02 2.50e+02
-TRACE baseline 4.56e+09 4.41e+08 4.44e+07 4.61e+06 5.31e+05 6.71e+04 9.88e+03 2.20e+03 9.05e+02 6.03e+02 5.53e+02 5.32e+02
+TRACE DRAFT 4.56e+09 4.41e+08 4.44e+07 4.61e+06 5.31e+05 6.71e+04 9.88e+03 2.20e+03 9.05e+02 6.03e+02 5.53e+02 5.32e+02
 Noise 9.11e+08 1.07e+08 1.34e+07 1.74e+06 2.43e+05 3.71e+04 6.63e+03 1.78e+03 8.35e+02 5.93e+02 5.51e+02 5.32e+02
@@
 SAWAR 8.79e+02 7.60e+02 6.82e+02 6.29e+02 5.94e+02 5.69e+02 5.53e+02 5.41e+02 5.33e+02 5.28e+02 5.27e+02 5.27e+02
-dataDIVAT1 baseline 1.43e+04 8.36e+03 4.95e+03 3.01e+03 1.91e+03 1.31e+03 9.99e+02 8.51e+02 7.84e+02 7.56e+02 7.51e+02 7.49e+02
+dataDIVAT1 DRAFT 1.43e+04 8.36e+03 4.95e+03 3.01e+03 1.91e+03 1.31e+03 9.99e+02 8.51e+02 7.84e+02 7.56e+02 7.51e+02 7.49e+02
 Noise 5.69e+08 7.64e+07 1.05e+07 1.44e+06 2.07e+05 3.17e+04 5.75e+03 1.61e+03 9.08e+02 7.81e+02 7.67e+02 7.64e+02
@@
 SAWAR 9.18e+02 8.55e+02 8.21e+02 8.00e+02 7.86e+02 7.75e+02 7.66e+02 7.59e+02 7.55e+02 7.52e+02 7.51e+02 7.51e+02
-flchain baseline 3.35e+23 9.34e+20 2.56e+18 7.37e+15 2.23e+13 6.82e+10 2.29e+08 9.33e+05 1.67e+04 2.03e+03 1.25e+03 1.09e+03
+flchain DRAFT 3.35e+23 9.34e+20 2.56e+18 7.37e+15 2.23e+13 6.82e+10 2.29e+08 9.33e+05 1.67e+04 2.03e+03 1.25e+03 1.09e+03
 Noise 2.18e+35 1.20e+31 7.58e+26 4.68e+22 3.48e+18 2.65e+14 2.21e+10 6.77e+06 4.45e+04 1.96e+03 1.22e+03 1.11e+03
@@
 SAWAR 6.31e+03 2.22e+03 1.43e+03 1.19e+03 1.12e+03 1.09e+03 1.08e+03 1.07e+03 1.07e+03 1.07e+03 1.07e+03 1.07e+03
-prostate baseline 9.74e+05 3.22e+05 1.06e+05 3.58e+04 1.21e+04 4.25e+03 1.64e+03 7.62e+02 4.66e+02 3.70e+02 3.51e+02 3.44e+02
+prostate DRAFT 9.74e+05 3.22e+05 1.06e+05 3.58e+04 1.21e+04 4.25e+03 1.64e+03 7.62e+02 4.66e+02 3.70e+02 3.51e+02 3.44e+02
 Noise 7.00e+20 5.05e+18 3.75e+16 2.85e+14 2.28e+12 2.03e+10 2.59e+08 9.04e+06 4.85e+05 6.66e+04 2.94e+04 1.37e+04
@@
 SAWAR 9.65e+03 3.88e+03 1.66e+03 8.48e+02 5.43e+02 4.30e+02 3.83e+02 3.59e+02 3.45e+02 3.37e+02 3.35e+02 3.34e+02
-retinopathy baseline 1.11e+04 5.64e+03 2.88e+03 1.51e+03 8.08e+02 4.64e+02 2.98e+02 2.18e+02 1.84e+02 1.70e+02 1.68e+02 1.67e+02
+retinopathy DRAFT 1.11e+04 5.64e+03 2.88e+03 1.51e+03 8.08e+02 4.64e+02 2.98e+02 2.18e+02 1.84e+02 1.70e+02 1.68e+02 1.67e+02
 Noise 1.11e+04 5.61e+03 2.88e+03 1.51e+03 8.13e+02 4.68e+02 2.99e+02 2.19e+02 1.84e+02 1.70e+02 1.68e+02 1.68e+02
@@
 SAWAR 4.02e+02 3.31e+02 2.79e+02 2.41e+02 2.14e+02 1.95e+02 1.82e+02 1.74e+02 1.69e+02 1.67e+02 1.67e+02 1.67e+02
-stagec baseline 1.16e+06 2.94e+05 7.50e+04 1.85e+04 4.81e+03 1.28e+03 3.57e+02 1.14e+02 5.35e+01 4.10e+01 3.98e+01 3.97e+01
+stagec DRAFT 1.16e+06 2.94e+05 7.50e+04 1.85e+04 4.81e+03 1.28e+03 3.57e+02 1.14e+02 5.35e+01 4.10e+01 3.98e+01 3.97e+01
 Noise 2.07e+06 4.93e+05 1.14e+05 2.73e+04 6.49e+03 1.61e+03 4.19e+02 1.25e+02 5.50e+01 4.10e+01 3.96e+01 3.95e+01
@@
 SAWAR 1.03e+03 5.79e+02 3.30e+02 1.94e+02 1.21e+02 8.21e+01 6.10e+01 4.88e+01 4.22e+01 3.89e+01 3.83e+01 3.81e+01
-zinc baseline 2.62e+06 5.85e+05 1.33e+05 3.13e+04 7.43e+03 1.87e+03 5.36e+02 2.03e+02 1.11e+02 8.26e+01 7.71e+01 7.46e+01
+zinc DRAFT 2.62e+06 5.85e+05 1.33e+05 3.13e+04 7.43e+03 1.87e+03 5.36e+02 2.03e+02 1.11e+02 8.26e+01 7.71e+01 7.46e+01
 Noise 2.93e+07 4.72e+06 7.76e+05 1.32e+05 2.34e+04 4.45e+03 9.48e+02 2.56e+02 1.14e+02 8.17e+01 7.69e+01 7.53e+01
@@
 SAWAR 1.31e+03 7.77e+02 4.72e+02 2.98e+02 1.98e+02 1.42e+02 1.10e+02 9.21e+01 8.28e+01 7.87e+01 7.80e+01 7.78e+01
-TABLE VI: Negative Log Likelihood metric for SurvSet datasets (lower is better) for each adversarial training method against the worst-case adversarial attack.
+TABLE VIII: Negative Log Likelihood metric for SurvSet datasets (lower is better) for each adversarial training method against the worst-case adversarial attack.

### v4

@@
 Survival Analysis (SA) models the time until an event occurs, with applications in fields like medicine, defense, finance, and aerospace.
-Recent work shows that Neural Networks (NNs) can capture complex relationships in SA.
-However, dataset uncertainties (e.g., noisy measurements, human error) can degrade model performance.
-To address this, we leverage NN verification advances to create algorithms for robust, fully-parametric survival models.
-We introduce a robust loss function and use CROWN-IBP regularization to handle computational challenges in the Min-Max problem.
-Evaluating our approach on SurvSet datasets, we find that our Survival Analysis with Adversarial Regularization (SAWAR) method consistently outperforms baselines under various perturbations with respect to Negative Log Likelihood (NegLL), Integrated Brier Score (IBS), and Concordance Index (CI).
-This demonstrates that adversarial regularization enhances SA performance and calibration, mitigating data uncertainty and improving generalization across diverse datasets up to 150% across all perturbation magnitudes.
+Recent research indicates that Neural Networks (NNs) can effectively capture complex data patterns in SA, whereas simple generalized linear models often fall short in this regard.
+However, dataset uncertainties (e.g., noisy measurements, human error) can degrade NN model performance.
+To address this, we leverage advances in NN verification to develop training objectives for robust, fully-parametric SA models.
+Specifically, we propose an adversarially robust loss function based on a Min-Max optimization problem.
+We employ CROWN-Interval Bound Propagation (CROWN-IBP) to tackle the computational challenges inherent in solving this Min-Max problem.
+Evaluated over 10 SurvSet datasets, our method, Survival Analysis with Adversarial Regularization (SAWAR), consistently outperforms baseline adversarial training methods and state-of-the-art (SOTA) deep SA models across various covariate perturbations with respect to Negative Log Likelihood (NegLL), Integrated Brier Score (IBS), and Concordance Index (CI) metrics.
+Thus, we demonstrate that adversarial robustness enhances SA predictive performance and calibration, mitigating data uncertainty and improving generalization across diverse datasets by up to 150% compared to baselines.
 Code: https://github.com/mlpotter/SAWAR
@@
 I Introduction
-Survival Analysis (SA) models the time until an event occurs and has extensive applications.
-For example, SA has been used extensively in medicine for evaluating treatments in cancer patients and drug trials [1].
-Early methods relied on linear parametric models [2], but recent advancements utilize Neural Networks (NNs) to handle complex time-to-event data [3].
-A major challenge in Survival Analysis (SA) is developing models robust to dataset uncertainty and variability.
-This is particularly essential for clinical applications where treatment efficacy must be demonstrated across diverse patient demographics, institutions, and data acquisition settings [4, 5, 6].
-Data uncertainty can arise from human errors, such as poor handwriting by physicians [7] and ineffective medical record-keeping [8], as well as inherent noise during data collection [9].
-Many approaches assume perfect measurement of covariates, neglecting aleatoric uncertainty [10], which in SA is associated with noise in the log of measured time-to-event [11].
-Further, adversaries can manipulate data to make models over- or under-confident [12].
-Despite these issues and the sensitivity of NNs to input uncertainties, robustness techniques against data perturbations in SA remains underexplored.
-To address these issues, this paper’s contributions include:
-  • Survival Analysis with Adversarial Regularization (SAWAR) for training robust, fully-parametric, NN-based SA models using adversarial covariate perturbations and convex relaxations, enhancing model calibration and stability against noisy data.
-  • Experiments demonstrating the sensitivity of existing NN-based SA methods to covariate perturbations across 10 publicly available medical SA datasets, highlighting an area for improvement in the current state of the art.
+Survival Analysis (SA) models the time until an event occurs with extensive applications in various fields including medicine, system reliability, economics, and marketing.
+Medical research uses Survival Analysis (SA) to evaluate treatment efficacy and illness progression [1].
+Engineers assess a system’s reliability with SA to improve design and maintenance [2].
+Economists and social scientists use SA (also known as duration analysis) to investigate the length of unemployment spells and other major life events [3].
+Marketing uses SA to aid in the examination of customer retention and churn rates [4].
+While many of these domains have traditionally relied on linear parametric SA models [5], recent advancements have leveraged Neural Networks (NNs) to handle such complex time-to-event data [6]
+A major challenge in SA is developing models that effectively handle complex time-to-event data while also ensuring robustness against dataset uncertainty, accurate calibration, and strong generalization.
+For instance, data uncertainty in medical applications can arise from various sources, including human factors such as illegible handwriting by physicians [7] and poor medical record-keeping [8], as well as inherent noise from medical equipment during data collection [9], termed as aleatoric uncertainty.
+Most of the existing approaches address aleatoric uncertainty by assuming noise in the logarithm of measured time-to-event data [10].
+However, they often overlook uncertainty originating from the model’s input covariates.
+SA models ought to generalize over various input covariates to sustain high accuracy in real-life applications.
+In medical studies, this entails high accuracy over diverse patient demographics, institutions, and data acquisition settings.
+While traditional SA models, such as Linear Cox-Weibull model [5], offer inductive biases specific to time-to-event analysis that aid generalization, they struggle with complex data relationships.
+On the other hand, while NN models can manage complex data, they may not generalize well to unseen data [11].
+Combining the inductive biases of traditional SA models with NN techniques—such as modeling intensity or hazard functions with NNs —can both improve generalization and handle complex data [12].
+Nonetheless, NN in SA may still suffer from poor model calibration, even with good generalization.
+NNs are particularly sensitive to input perturbations [13].
+Adversaries can exploit this sensitivity to manipulate data, causing models to make over- or under-confident predictions [14, 15].
+Additionally, data perturbations may be intentionally introduced with techniques like differential privacy to protect sensitive information, such as personally identifiable information of patients [16].
+Although better calibration does not always enhance generalization [17], integrating adversarial robustness to input covariate perturbation can jointly improve model stability, trustworthiness, and calibration when encountering unseen data [18].
+Despite these challenges and the sensitivity of NNs to input uncertainties, robustness techniques for handling data perturbations in NN-based SA remain underexplored.
+To address these challenges, we make the following contributions:
+  • We propose a novel training objective, called Survival Analysis with Adversarial Regularization (SAWAR), designed to create robust, fully-parametric NN-based SA models by incorporating adversarial covariate perturbations via convex relaxations, and thereby improve model calibration and generalization against noisy data.
+  • Extensive experiments across 10 publicly-available SA datasets reveal the sensitivity of existing NN-based SA models to covariate perturbations with respect to Concordance Index (CI), Integrated Brier Score (IBS), and Negative Log Likelihood (NegLL) metrics, as well as survival curves.
+  • SAWAR enhances SA predictive performance and calibration, mitigating data uncertainty and improving generalization by up to 150% compared to baseline adversarial training methods and state-of-the-art (SOTA) NN-based SA models.
 II Related Work
-This section reviews NN literature for SA, with more details in [13].
-The first use of NNs in SA was a shallow Feedforward Neural Network (FFNN) to approximate the survival function [14].
-Later, deeper models were developed, such as Cox-nnet [15], a NN for Cox Proportional Hazard (CPH) modeling [11].
-Building on this, DeepSurv [16] introduced a CPH NN with dropout layers for better generalization in personalized treatment.
-Then, mixture models such as DeepWeiSurv [17] further advanced the field by using NNs to regress Weibull mixture distribution parameters.
-These methods enhance modeling capability but do not always improve model calibration and robustness.
-Recent works focus on improving model calibration [18, 19, 20], which ensures predicted confidence matches true likelihood [21].
-However, few works address adversarial robustness in SA, where small input perturbations should not mislead the model [22].
-Although better calibration does not always enhance generalization [23], integrating adversarial robustness improves model trustworthiness in calibration and and stability on unexpected data [24].
-Existing adversarial training methods for SA involve tradeoffs between model complexity, performance, and incorporating physical assumptions.
-Most techniques use Conditional Generative Adversarial Networks (CGANs) to generate time-to-event samples from a survival distribution non-parametrically [25, 26, 27].
-These models implicitly define the survival distribution without assuming a functional/analytical form, offering increased flexibility to fit various data realizations [28, 29, 30].
-However, this flexibility can violate Bayesian Occam’s Razor, which penalizes overly complex models [30].
-To balance complexity and performance, inductive biases are used in explicit parametric models [29].
-One common bias in SA is the memoryless property, satisfied by the exponential distribution [28].
-Thus, this paper employs the exponential CPH model, widely used in SA [31, 32, 33], as well as in queuing theory [34], reliability analysis [3], and telecommunications [35].
+The integration of NNs into SA began with initial models that replaced traditional linear approaches.
+Early works used a shallow Feedforward Neural Network (FFNN) to replace the linear estimate of the log-hazard in a Cox Proportional Hazard (CPH) model [19].
+Building on this, Cox-nnet [20] substituted the input covariates of the linear log-hazard function of a CPH model [21] with a deeper NN.
+Subsequently, DeepSurv [22] further advanced this approach by replacing the entire linear estimate with a deep NN and adding dropout to improve generalization.
+The trend continued with mixture models like DeepWeiSurv [23], which used NNs to regress Weibull mixture distribution parameters, offering increased modeling flexibility while maintaining traditional statistical model biases.
+As NN models evolved, the SA field saw the introduction of non-parametric approaches.
+For instance, DeepHit utilized NNs to learn the distribution of survival times directly without assumptions about the underlying stochastic process [24].
+Even more flexible techniques emerged, such as those employing Conditional Generative Adversarial Networks (CGANs) to generate time-to-event samples non-parametrically [25, 26, 27], with CGAN serving as an implicit likelihood [28].
+This approach defines the survival distribution without assuming a specific functional form [29, 30].
+For a more detailed overview of deep learning methods for SA, we refer the readers to [12].
+Despite these advances, integrating NNs with traditional SA models has been limited in addressing model calibration and robustness as discussed in Section I.
+Recent research has thus focused on improving model calibration [31, 32, 33] to ensure that predicted confidence levels reflect true likelihoods [34].
+However, better calibration does not necessarily enhance generalization [17].
+On the other hand, integrating adversarial robustness has been shown to improve both calibration and generalization when dealing with unexpected data [18].
+Nonetheless, there remains a gap in addressing adversarial robustness in SA to ensure that where small input perturbations do not mislead model predictions [15].
+Existing adversarial training methods for SA often face trade-offs between model complexity, prediction performance, and the incorporation of physical assumptions.
+While Generative Adversarial Network (GAN)-inspired methods incorporate adversarial components through a minimax game between the generator and discriminator [35, 36], they typically exhibit training instability and may violate Bayesian Occam’s Razor that advocates simpler models [30].
+To balance complexity, interpretability, and generalization, several methods adopt time-to-event inductive biases through specific functional forms of SA models with deep NN and/or non-parametric approaches [22, 20, 23, 37].
+One such inductive bias in SA is the memoryless property, satisfied by the exponential distribution [29].
+This paper employs the exponential CPH model (c.f.
+Section III), as it is widely used in SA [38, 39, 40], as well as in related fields like queuing theory [41], reliability analysis [6], and telecommunications [42].
+Crucially, this fully-parametric approach allows us to create an adversarially robust NN-based SA model that outperforms SOTA NN-based SA models including such as Deep Regularized Accelerated Failure Time (DRAFT) [25] and Adversarial Autoencoder with a Deep Cox Proportional Hazard (AAE-Cox) [37].
+AAE-Cox aims to balance modeling flexibility and inductive biases by combining adversarial GAN autoencoders with Deep Cox proportional networks.
+AAE-Cox has been shown to outperform methods like DeepSurv, TCAP [43], and DCAP [44].
 III Preliminaries on Survival Analysis
+Figure 1: Visual of time-to-event dataset with exact event times and right censoring. The solid circle on each patient line indicates when the event occurred for a given patient.
 SA [21] is a statistical approach used to analyze and model the time $t\in\mathcal{R_{++}}$ until an event $e\in\{0,1\}$ happens (e.g., a disease occurs, a device breaks).
-Typically, the study happens over a finite time period and instances/individuals may leave the study early before an event is recorded, with no follow up.
+Typically, a study of such events happens over a finite time period and instances/individuals may leave the study early before an event is recorded, with no follow-up.
 In those cases, the dataset records that no event occurred up until the exit time of the study, which is called right censoring  [45].
-Such an SA dataset consists of $N$ independent instances/individuals, each denoted by index $i$, for which time between entry to a study and the subsequent event is recorded.
-Upon entry time to a study, each instance’s covariates $\boldsymbol{x}_{i}\in\mathcal{R}^{d}$ (i.e., variables that may affect the event of interest for the study) are measured.
+Such an SA dataset consists of $N$ independent instances, each denoted by index $i\in[1,\ldots,N]$, for which the time between entry to a study and the subsequent event is recorded.
+Upon entry time to a study, each instance’s covariates $\boldsymbol{x}_{i}\in\mathcal{R}^{d}$ (i.e., variables that govern the event of interest) are measured.
+A visual of time-to-event data with right censoring is shown in Fig.
+1.
 In summary, the dataset $\mathcal{D}$ has the following structure:
@@
 The survival function $S(t):=1-F(t)$ is the complement of this CDF [46, 21].
-An exponential parametric survival model is expressed as
+An exponential SA model has the parametric form:
 $S(t)=e^{-\Lambda(t)},$ (2)
 […]
-$\lambda(t)=\frac{f(t)}{S(t)}=\lim_{\Delta\rightarrow 0}\frac{P[t<T<t+\Delta]}{\Delta}$ (3)
+$\lambda(t)=\frac{f(t)}{S(t)}=\lim_{\Delta\rightarrow 0}\frac{P[t<T<t+\Delta]}{\Delta},$ (3)
 is the instantaneous failure rate at time $t$.
-To incorporate the instance’s covariates, we use the prevalent CPH model [2]:
+To incorporate each instance’s covariates in predicting failure rate, we use the prevalent CPH model [5]:
 $\lambda_{\boldsymbol{\theta}}(t,\boldsymbol{x})=\lambda_{0}(t)e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})},$ (4)
-where $\lambda_{0}(t)$ is a baseline hazard, $e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}$ is the relative risk associated with covariates $\boldsymbol{x}$, and $G_{\boldsymbol{\theta}}(\boldsymbol{x})$ is an NN with parameters $\boldsymbol{\theta}$ in our setting.
-The exponential CPH model characterizes the baseline hazard rate as $\lambda_{0}(t)=\lambda_{0}$.
-For simplicity, we assume an exponential baseline hazard and “absorb” the $\lambda$ term into the bias of the NN:
-$\lambda_{0}e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}=e^{\log\lambda_{0}+G_{\boldsymbol{\theta}}(\boldsymbol{x})}=e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}.$ (5)
-The exponential CPH model’s PDF, CDF, and complement CDF are $f_{\boldsymbol{\theta}}(t|\boldsymbol{x})=e^{G_{\theta}(\boldsymbol{x})}e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$, $F_{\boldsymbol{\theta}}(t|\boldsymbol{x})=1.0-e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$ , $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})=e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$, respectively.
+where $\lambda_{0}(t)$ is a baseline hazard, $e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}$ is the relative hazard associated with covariates $\boldsymbol{x}$, and $G_{\boldsymbol{\theta}}(\boldsymbol{x})$ is an NN with parameters $\boldsymbol{\theta}$ in our setting.
+The exponential CPH model assumes a time-independent baseline hazard rate $\lambda_{0}(t)=\lambda_{0}$. This allows us to “absorb” the $\lambda_{0}$ term into the bias of the NN:
+$\lambda_{\boldsymbol{\theta}}(t,\boldsymbol{x})=e^{\log\lambda_{0}+G_{\boldsymbol{\theta}}(\boldsymbol{x})}=e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}.$ (5)
+Further, the exponential CPH model’s PDF, CDF, and complement CDF are given by $f_{\boldsymbol{\theta}}(t|\boldsymbol{x})=e^{G_{\theta}(\boldsymbol{x})}e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$, $F_{\boldsymbol{\theta}}(t|\boldsymbol{x})=1-e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$ , $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})=e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$, respectively.
 $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ is also known as the instance survival function.
-The population survival curve is the marginalization of the instance survival function over the covariates,
+The population survival curve is defined as the marginalization of the instance survival function over the covariates:
 […]
-Since the distribution of the covariates $p(\boldsymbol{x})$ is typically unknown, the integral in Eq. 6 is often replaced with a Monte Carlo estimate [18]:
+Since the distribution of covariates $p(\boldsymbol{x})$ is typically unknown, the integral in Eq. 6 is often replaced with a Monte Carlo estimate over a dataset $\mathcal{D}$ [31]:
 $S(t)$ $\approx\frac{1}{N}\sum_{i=1}^{N}S_{\boldsymbol{\theta}}(t|\boldsymbol{x}_{i}).$ (7)
-Given these definitions, the next section describes the baseline objective for finding model parameters $\boldsymbol{\theta}$ from dataset $\mathcal{D}$.
+Given these definitions, the next section describes the baseline objective for estimating model parameters $\boldsymbol{\theta}$ from dataset $\mathcal{D}$.
 III-B Baseline Objective Functions
-We follow [38, 25] to define two objective functions based on negative log likelihood (Section III-B1) and the rank correlation (Section III-B2).
+We employ an objective function that combines negative log likelihood (Section III-B1) and rank correlation (Section III-B2), following [47, 25].
 III-B1 Right Censored Log Likelihood Objective
-The $\mathcal{L}_{LL}$ objective is the log of the time-to-event likelihood over $\mathcal{D}$ specified by the CPH model [11]:
+The $\mathcal{L}_{LL}$ objective is the logarithm of the time-to-event likelihood over $\mathcal{D}$ as specified by the CPH model [21]:
 $\mathcal{L}_{LL}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})=$ $\sum_{i=1}^{N}e_{i}\cdot\log f_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x_{i}})\ +$
@@
 The first summation in Eq.
-8 is the log likelihood of the datapoints corresponding to when the event occurs and the exact time of occurrence is observed.
+8 is the log likelihood of the datapoints corresponding to when the event occurs and the exact time that occurrence is observed.
 The second summation in Eq.
-8 is the log likelihood when right censoring occurs.
+8 is the log likelihood when right censoring occurs, where it is known that no event occurred at least up until the censoring time.
 III-B2 Ranking Objective
@@
 $\mathcal{L}_{rank}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})$ $=\sum_{i\neq j}A_{i,j}\eta\bigg{(}F_{\theta}(t_{i}|\boldsymbol{x}_{i}),F_{\boldsymbol{\theta}}(t_{i}|\boldsymbol{x}_{j})\bigg{)},$ (9)
-where $\eta(\mathbf{x},\mathbf{y})=\exp(-\frac{(\mathbf{x}-\mathbf{y})}{\sigma})$ and $A_{i,j}$ indicates an “acceptable” comparison pair [38, 39]. We set $\sigma=1$ for all experiments.
+where $\eta(\mathbf{x},\mathbf{y})=\exp(-\frac{(\mathbf{x}-\mathbf{y})}{\sigma})$ and $A_{i,j}$ indicates an “acceptable” comparison pair [47, 48].
+Two instances are comparable if the instance with lower observed time experienced an event, i.e., if $A_{i,j}=\mathds{1}(t_{i}<t_{j}\cap e_{i}=1)$.
+We set $\sigma=1$ for all experiments.
+This objective is a continuous differentiable proxy for the non-continuous non-differentiable concordance index [49].
 III-B3 Combined Objective
-The combined right censored negative log likelihood objective and the risk ranking objective, with risk ranking objective weight trade-off $w=\frac{1}{\text{batch\_size}}$, is
-$\mathcal{L(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})}=-\mathcal{L}_{LL}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})+w\cdot\mathcal{L}_{rank}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e}).$ (10)
-This combined objective provides a trade-off between finding the NN parameters $\boldsymbol{\theta}$ that maximize the likelihood of observing $\mathcal{D}$ and maximize a rank correlation between the instance risk scores and the observed failure times.
-We will denote the combined loss as the baseline objective.
-Our baseline model’s objective is similar to Deep Regularized Accelerated Failure Time (DRAFT) [25] but with a different ranking objective function.
-Thus, we consider the baseline training method for robust SA baseline method as DRAFT.
-Thus far, we have described an objective function that improves predictive performance:
+The combined objective over the right censored dataset $\mathcal{D}$ is:
+$\mathcal{L(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})}=-\mathcal{L}_{LL}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})+w\cdot\mathcal{L}_{rank}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e}),$ (10)
+where $w$ is the hazard versus ranking objective weight trade-off. Accordingly, $\boldsymbol{\theta}$ is estimated by solving the following optimization problem:
 $\boldsymbol{\theta}^{*}=\arg\min_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e}).$ (11)
-In the following sections, we address how to develop NN models that are robust against adversarial perturbations.
+This combined objective provides a trade-off between maximizing the likelihood of observing $\mathcal{D}$ and maximizing a rank correlation between the instance hazard scores and the observed failure times.
+From here on, we term this combined loss as the baseline objective.
+We note that an NN-based SA model DRAFT [25] also employs a combined objective.
+Our baseline objective incorporates a different ranking objective than DRAFT and outperforms DRAFT as discussed in Section VI.
 IV Preliminaries on Adversarial Training
@@
 where $\tilde{\boldsymbol{X}}=\{\tilde{\boldsymbol{x}}_{i}\in\mathcal{B}(\boldsymbol{x}_{i},\epsilon)\}_{i=1}^{N}$, and $\mathcal{B}(\boldsymbol{x}_{i},\epsilon)$ is the $\ell_{\infty}$-ball of radius $\epsilon$ around $\boldsymbol{x}_{i}$.
-In practice, the Min-Max problem posed in [42] is intractable to compute exactly.
-In particular, the inner maximization requires finding a global optimum over a high dimensional and non-convex loss function over NN parameters.
+In practice, the Min-Max problem posed in [13] is intractable to solve exactly.
+In particular, the inner maximization requires finding a global optimum over a high dimensional and non-convex loss function over input covariates with the current NN parameters.
 Therefore, many recent adversarial training methods propose approximations to Eq.
@@
 To understand the robustness properties of SA models to a variety of data uncertainties present in real applications, this paper investigates several perturbation methods, described next.
-IV-A Certifiable Robustness
+IV-A Verifiable Robustness
 One such approximation is to use techniques from NN verification during training.
-For example, CROWN-Interval Bound Propagation (CROWN-IBP) combines CROWN and Interval Bound Propogation (IBP) methods to obtain a convex relaxation for the lower bound and upper bound of each layer in the NN with respect to input data [40, 43].
+Neural network robustness verification algorithms determine the output neuron’s upper and lower bounds for all inputs in a set $\mathcal{B}$, typically constrained by a norm-bounded perturbation [51].
+For example, CROWN-Interval Bound Propagation (CROWN-IBP) combines CROWN and Interval Bound Propogation (IBP) methods to obtain a convex relaxation for the lower bound and upper bound of each layer in the NN with respect to input data within the set $\mathcal{B}$.
+[50, 51, 52].
 CROWN-IBP is commonly used for verifying NN properties over a range of possible inputs.
@@
 […]
-This relaxation is convex with respect to NN weights, enabling efficient optimization, and by optimizing over an upper bound on the inner maximization, it accounts for a strong adversary during training.
+[…]
 IV-B Adversarial Perturbations
-Another way to approximate Eq.
-20 is to use adversarial perturbations.
-Whereas CROWN-IBP provides an upper bound on the inner maximization (i.e., $\max_{\boldsymbol{\tilde{X}}}\mathcal{L}(\theta;\boldsymbol{\tilde{X}})$), adversarial perturbation algorithms typically provide a lower bound:
+Another way to approximate Eq. 12 is to use adversarial perturbations.
+IV-B1 Fast Gradient Sign Method (FGSM)
+While CROWN-IBP provides an upper bound on the inner maximization (i.e., $\max_{\boldsymbol{\tilde{X}}}\mathcal{L}(\theta;\boldsymbol{\tilde{X}})$), adversarial perturbation algorithms typically provide a lower bound:
 $\max_{\boldsymbol{\tilde{X}}}\ \mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}})\geq\mathcal{L}(\boldsymbol{\theta};\boldsymbol{\tilde{X}}_{FGSM}).$ (14)
-Fast Gradient Sign Method (FGSM) [45] uses the gradient of the objective function with respect to the input $\boldsymbol{x}$ to perturb $\boldsymbol{x}$ in the direction that maximizes the objective $\mathcal{L}$.
+FGSM [54] uses the gradient of the objective function with respect to the input $\boldsymbol{x}$ to perturb $\boldsymbol{x}$ in the direction that maximizes the objective $\mathcal{L}$.
 Projected Gradient Descent (PGD) [13] iteratively applies that same type of first-order perturbation $K$ times (or until convergence is reached), but constrains the perturbation of $\boldsymbol{x}$ to remain within a specified set (e.g., $\ell_{p}$-ball).
@@
 $\boldsymbol{x}^{(0)}$ $=\boldsymbol{x},$ (15)
-$\boldsymbol{x}^{(k+1)}$ $=\Pi_{\mathcal{B}(\boldsymbol{x},\epsilon)}\left(\boldsymbol{x}^{(k)}+\alpha\big{(}\nabla_{\boldsymbol{x}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{x}^{(k)},\boldsymbol{t},\boldsymbol{e})\big{)}\right),$ (16)
-$\boldsymbol{\tilde{x}}$ $=\boldsymbol{x}^{(K)},$ (17)
-where $k$ is the iteration step and $\alpha$ is a step size parameter.
-IV-C Random Noise
+$\boldsymbol{x}^{(k+1)}$ $=\Pi_{\mathcal{B}(\boldsymbol{x},\epsilon)}\left(\boldsymbol{x}^{(k)}+\alpha\big{(}\nabla_{\boldsymbol{x}}\mathcal{L}(\boldsymbol{\theta};\boldsymbol{x}^{(k)},\boldsymbol{t},\boldsymbol{e})\big{)}\right),\forall k$
+$\boldsymbol{\tilde{x}}$ $=\boldsymbol{x}^{(K)}$
+where $k\in[1,\ldots,K-1]$ is the iteration step and $\alpha=\epsilon/K$ is a step size parameter.
+IV-B2 Random Noise
 While robustness to a worst-case perturbation is important for many applications, adding random noise to training data has also been shown to improve generalization performance [55].
@@
 V Survival Analysis with Adversarial Regularization
-This paper addresses the open challenge in SA of developing models robust to dataset uncertainty and variability, as motivated in Section I.
-We propose a method for adversarial robustness in fully-parametric NN-based SA, enhancing calibration and stability on unexpected or noisy data.
+This paper addresses the open challenge in SA of developing models robust to dataset uncertainty, generalizable, and calibrated.
+We propose a training objective for adversarial robustness in fully-parametric NN-based SA, enhancing calibration and generalization on unexpected or noisy data.
 Our approach extends the objective function (Eq.
@@
 When solving for $\boldsymbol{\theta}$ in SA, many approaches treat each measurement in dataset $\mathcal{D}$ as ground-truth, ignoring aleatoric uncertainty [10].
-Typically in SA, the aleatoric uncertainty is only associated with the noise of the log of the measured time-to-event (e.g.
+Typically in SA, the aleatoric uncertainty is only associated with the noise in the logarithm of measured time-to-event (e.g.
 the Gumbel distribution or Log Normal distribution) [21].
@@
 where $\kappa\in(0,1)$.
-V-C SAWAR Training
-The NN parameters are updated using stochastic gradient descent with the ADAM optimizer [47], where we use early stopping [48] with respect to the validation set to prevent overfitting We introduce the covariate perturbations after 100 epochs.
-A smooth $\epsilon$-scheduler linearly increases the $\epsilon$-perturbation magnitude from $0.0$ to $0.5$ within a 30 epoch window after 100 epochs.
-We do not allow early stopping until after the maximum $\epsilon$-perturbation magnitude is achieved.
 We expect that the model’s adversarial robustness and predictive calibration should increase when incorporating data uncertainty in SAWAR objective [18] via the adversarial regularization term in Eq.
@@
 VI Experiments
-We conduct experiments over 10 benchmark medical SA datasets (VI-B) to quantify the impact of perturbation after the robust training is complete.
-As evaluation metrics, we consider the Concordance Index (CI) [49], the Integrated Brier Score (IBS) [50] and the Negative Log Likelihood (NegLL) at varying $\epsilon$-perturbation magnitudes from 0 to 1.
-To demonstrate that SAWAR produces SA models that perform well and are robust to dataset uncertainty, this section quantifies the improvement in predictive accuracy, calibration, and population curve estimation.
-VI-A Configuration
-The experiments are conducted on a Windows computer with a 12th Gen Intel(R) Core (TM) i9-12900 processor and 16 GB RAM.
-Each dataset undergoes stratified partitioning into train, validation, test sets with proportions 60%, 20%, 20% respectively.
-The NN $G_{\boldsymbol{\theta}}$ has 2 hidden layers of 50 neurons each and Leaky ReLU activation functions.
-We use lifelines [51] and sksurv [39] for computing SA metrics.
-VI-B Datasets
+We conduct experiments over 10 benchmark medical SA datasets (Section VI-A) to quantify the impact of perturbation after the robust training is complete.
+To demonstrate that SAWAR produces SA models that are both high-performing and robust to dataset uncertainty, this section quantifies improvements in predictive accuracy, calibration, and population curve estimation compared to common adversarial training techniques (Section IV-B) including FGSM, Random Noise and PGD, as well as SOTA NN-based SA models, such as DRAFT and AAE-Cox.
+VI-A Datasets
 We use the API SurvSet [56] to download the benchmark medical SA datasets shown in Table I and described below:
@@
 We preprocess each dataset by one-hot encoding categorical covariates and standard normalization of all covariates.
+Each dataset undergoes stratified partitioning into train, validation, test sets with proportions 60%, 20%, 20% respectively.
+VI-B Experiment Setup
+The experiments are conducted on a Windows computer with a 12th Gen Intel(R) Core (TM) i9-12900 processor and 16 GB RAM.
+The NN $G_{\boldsymbol{\theta}}$ has 2 hidden layers of 50 neurons each and Leaky ReLU activation functions.
+For AAE-Cox, we adopt the publicly available code and hyperparameters [37].
+For DRAFT, we use the same hyperparameters as SAWAR as described below, excluding hyperparameters related to adversarial regularization.
+The NN parameters are optimized by solving Eq.
+(19) using stochastic gradient descent with the ADAM optimizer [67].
+We set $w=\frac{1}{\text{batch\_size}}$ and use early stopping [68] with respect to the validation set to prevent overfitting.
+We introduce the covariate perturbations after a set number of epochs that varies per dataset.
+A smooth $\epsilon$-scheduler linearly increases the $\epsilon$-perturbation magnitude from $0.0$ to $0.5$ within a 30 epoch window after the aforementioned set number of epochs.
+We do not allow early stopping until after the maximum $\epsilon$-perturbation magnitude of $0.5$ is achieved.
+As evaluation metrics, we consider the CI [49], the IBS [69] and the NegLL at varying $\epsilon$-perturbation magnitudes from 0 to 1. We use lifelines [70] and sksurv [48] for computing SA metrics.
 Figure 2: Distribution of survival curves for DRAFT, FGSM, PGD, and CROWN-IBP.
@@
 VI-C Adversary Setup
-We apply two adversarial perturbation settings on the test set to assess the robustness of the proposed adversarial regularization method: FGSM perturbation (Section IV-B) and a novel worst-case perturbation.
-FGSM is applied by using the gradient of the objective function with respect to the input $\boldsymbol{x}$ to perturb $\boldsymbol{x}$ in the direction that maximizes the objective Eq.
-20, such that:
-$\epsilon$ Concordance Index Integrated Brier Score Negative Log Likelihood
-DRAFT Noise FGSM PGD SAWAR DRAFT Noise FGSM PGD SAWAR DRAFT Noise FGSM PGD SAWAR
-1.00 2.9 3.35 3.35 2.5 2.9 4.6 4.4 2.4 1.9 1.7 4.0 4.7 2.7 2.1 1.5
-0.90 3.1 3.5 3.45 2.45 2.5 4.5 4.5 2.5 1.8 1.7 4.0 4.7 2.9 1.9 1.5
-0.80 3.4 3.4 3.2 2.35 2.65 4.6 4.4 2.6 1.8 1.6 4.1 4.6 3.0 1.9 1.4
-0.70 3.45 3.35 3.1 2.4 2.7 4.7 4.3 2.6 1.8 1.6 4.2 4.5 3.0 2.0 1.3
-0.60 3.45 3.45 3.0 2.5 2.6 4.7 4.3 2.7 1.8 1.5 4.2 4.5 3.1 2.1 1.1
-0.50 3.85 3.65 3.1 2.1 2.3 4.4 4.6 2.9 1.8 1.3 4.1 4.5 3.2 2.1 1.1
-0.40 3.9 3.8 3.05 2.1 2.15 4.4 4.6 3.0 1.9 1.1 4.0 4.6 3.2 2.1 1.1
-0.30 4.4 4.2 3.0 1.9 1.5 4.4 4.6 3.0 1.9 1.1 4.0 4.6 3.2 2.1 1.1
-0.20 4.35 4.0 3.3 2.25 1.1 4.2 4.6 3.2 2.0 1.0 3.9 4.7 3.2 2.1 1.1
-0.10 4.25 4.0 3.3 2.35 1.1 4.0 4.6 3.2 2.2 1.0 3.8 4.8 3.1 2.3 1.0
-0.05 3.85 4.0 3.45 2.6 1.1 3.8 4.7 2.9 2.2 1.4 3.9 4.7 3.1 2.3 1.0
-0.00 2.95 3.95 3.35 3.35 1.4 3.0 3.9 3.3 2.5 2.3 2.6 4.1 3.5 2.8 2.0
-TABLE II: FGSM average ranks for CI, IBS, and NegLL metrics across all datasets. Lower number is better.
-$\epsilon$ Concordance Index Integrated Brier Score Negative Log Likelihood
-DRAFT Noise FGSM PGD SAWAR DRAFT Noise FGSM PGD SAWAR DRAFT Noise FGSM PGD SAWAR
-1.00 2.8 3.4 3.5 3.6 1.7 2.6 4.3 3.8 3.3 1.0 3.1 4.5 3.4 3.0 1.0
-0.90 3.1 3.3 3.5 3.4 1.7 3.0 4.3 3.6 3.1 1.0 3.1 4.5 3.5 2.9 1.0
-0.80 3.0 3.3 3.5 3.7 1.5 3.0 4.5 3.5 3.0 1.0 3.0 4.6 3.5 2.9 1.0
-0.70 3.05 3.3 3.6 3.65 1.4 3.0 4.5 3.5 3.0 1.0 3.1 4.5 3.5 2.9 1.0
-0.60 2.85 3.55 3.55 3.75 1.3 3.0 4.5 3.5 3.0 1.0 3.0 4.7 3.5 2.8 1.0
-0.50 3.2 3.9 3.35 3.45 1.1 3.2 4.4 3.5 2.9 1.0 3.0 4.7 3.5 2.8 1.0
-0.40 3.25 3.95 3.2 3.5 1.1 3.3 4.8 3.2 2.7 1.0 3.1 4.7 3.4 2.8 1.0
-0.30 3.0 3.05 3.95 3.6 1.4 3.5 4.7 3.3 2.5 1.0 3.3 4.9 3.2 2.6 1.0
-0.20 2.05 2.95 4.2 3.8 2.0 3.7 4.7 3.4 2.2 1.0 3.4 4.9 3.4 2.3 1.0
-0.10 2.85 3.4 3.4 3.35 2.0 3.3 4.5 3.0 2.3 1.9 3.5 4.5 3.2 2.2 1.6
-0.05 3.05 4.15 3.4 2.9 1.5 2.6 4.4 3.0 2.6 2.4 3.3 4.4 3.0 2.4 1.9
-0.00 2.95 3.95 3.35 3.35 1.4 3.0 3.9 3.3 2.5 2.3 2.6 4.1 3.5 2.8 2.0
-TABLE III: Worst-Case average ranks for CI, IBS, and NegLL metrics across all datasets. Lower number is better.
-$\tilde{\lambda}_{\boldsymbol{\theta}}=e^{G_{\boldsymbol{\theta}}\left(\text{FGSM}(\mathcal{L},\boldsymbol{x})\right)}.$ (22)
+We apply two adversarial perturbation settings on the test set to assess the robustness of the proposed adversarial regularization method: FGSM perturbation (Section IV-B1) and a novel worst-case perturbation.
+The perturbed covariates $\boldsymbol{\tilde{x}}_{FGSM}$ with respect to the input $\boldsymbol{x}$ are found via Eq.
+15 and the corresponding failure rate becomes:
+$\tilde{\lambda}_{\boldsymbol{\theta}}=e^{G_{\boldsymbol{\theta}}\boldsymbol{\tilde{x}}_{FGSM}}.$ (20)
 We define the worst-case perturbation as the maximum failure rate with respect to the covariate uncertainty:
 $\tilde{\lambda}_{\boldsymbol{\theta}}=\max_{\boldsymbol{\tilde{x}_{i}}\in\mathcal{B}(\boldsymbol{x_{i}},\epsilon)}e^{G_{\boldsymbol{\theta}}(\boldsymbol{\tilde{x}}_{i})},$ (21)
-and the $\epsilon$-worst-case survival curve as the lowest survival probability at a given time $t$ with respect to the $\epsilon$-perturbed input covariates:
-$\tilde{S}(t|\boldsymbol{x})=e^{-\tilde{\lambda}_{\boldsymbol{\theta}}t}.$ (24)
+which we solve via CROWN-IBP (Section IV-B1) by optimizing a tight approximation of the upper bound of the maximization term.
+Accordingly, the $\epsilon$-worst-case survival curve has the lowest survival probability at a given time $t$ with respect to the $\epsilon$-perturbed input covariates:
+$\tilde{S}(t|\boldsymbol{\tilde{x}})=e^{-\tilde{\lambda}_{\boldsymbol{\theta}}t}.$ (22)
+$\epsilon$ Concordance Index Integrated Brier Score Negative Log Likelihood
+DRAFT Noise FGSM PGD AAE-Cox SAWAR DRAFT Noise FGSM PGD AAE-Cox SAWAR DRAFT Noise FGSM PGD AAE-Cox SAWAR
+1.00 3.8 4.4 3.1 2.7 3.8 3.2 5.4 5.6 3.2 2.3 2.6 1.9 5.1 5.8 3.6 2.6 2.5 1.4
+0.90 3.8 4.4 3.25 2.6 3.9 3.05 5.4 5.6 3.4 2.4 2.6 1.6 5.1 5.8 3.6 2.6 2.5 1.4
+0.80 3.9 4.4 3.65 2.9 3.6 2.55 5.4 5.5 3.6 2.5 2.5 1.5 5.1 5.8 3.6 2.6 2.5 1.4
+0.70 4.45 4.75 3.5 3.05 3.1 2.15 5.4 5.5 3.6 2.5 2.6 1.4 5.1 5.8 3.6 2.7 2.5 1.3
+0.60 4.9 5.0 3.5 3.0 2.8 1.8 5.3 5.6 3.6 2.7 2.5 1.3 5.2 5.7 3.6 2.7 2.5 1.3
+0.50 5.0 4.95 3.6 3.05 2.8 1.6 5.3 5.5 3.7 2.7 2.5 1.3 5.2 5.7 3.9 2.6 2.3 1.3
+0.40 5.3 4.7 3.7 3.3 2.6 1.4 5.3 5.5 3.9 2.7 2.3 1.3 5.3 5.6 3.9 2.6 2.3 1.3
+0.30 5.1 5.2 3.7 3.0 2.6 1.4 5.3 5.5 3.9 2.7 2.3 1.3 5.1 5.7 3.9 2.8 2.2 1.3
+0.20 5.5 5.1 3.9 2.7 2.3 1.5 5.3 5.5 3.9 2.9 2.1 1.3 4.9 5.8 4.0 2.9 2.1 1.3
+0.10 5.2 5.2 4.0 3.2 1.9 1.5 5.0 5.7 4.1 3.0 1.9 1.3 4.5 5.9 4.2 3.1 1.9 1.4
+0.05 5.4 5.0 3.9 3.3 2.0 1.4 4.6 5.6 4.3 3.4 1.6 1.5 4.5 5.9 4.3 3.3 1.7 1.3
+0.00 3.3 4.3 4.5 4.1 2.5 2.3 2.7 4.2 4.7 4.7 2.5 2.2 3.0 5.4 4.9 4.2 1.8 1.7
+TABLE II: FGSM average ranks for CI, IBS, and NegLL metrics across all datasets. Lower number is better. Bold and underline indicates the best and second best method, respectively.
+$\epsilon$ Concordance Index Integrated Brier Score Negative Log Likelihood
+DRAFT Noise FGSM PGD AAE-Cox SAWAR DRAFT Noise FGSM PGD AAE-Cox SAWAR DRAFT Noise FGSM PGD AAE-Cox SAWAR
+1.00 3.6 3.6 3.6 4.4 4.2 1.6 2.6 4.35 3.65 3.55 5.85 1.0 3.1 4.8 3.4 2.8 5.9 1.0
+0.90 3.4 3.9 3.6 4.3 4.4 1.4 2.9 4.25 3.65 3.35 5.85 1.0 3.1 4.8 3.4 2.8 5.9 1.0
+0.80 3.3 3.8 3.7 4.3 4.6 1.3 2.7 4.25 3.75 3.45 5.85 1.0 3.1 4.8 3.4 2.8 5.9 1.0
+0.70 3.4 3.8 3.8 4.1 4.6 1.3 2.8 4.65 3.65 3.15 5.75 1.0 3.1 4.8 3.5 2.7 5.9 1.0
+0.60 3.2 3.65 3.5 4.2 5.3 1.15 2.8 4.95 3.8 2.7 5.75 1.0 3.1 4.8 3.5 2.7 5.9 1.0
+0.50 3.0 3.6 3.5 4.1 5.6 1.2 3.1 4.8 3.8 2.7 5.6 1.0 3.1 4.8 3.5 2.7 5.9 1.0
+0.40 2.9 3.8 3.4 4.0 5.8 1.1 3.3 4.8 3.7 2.5 5.7 1.0 3.1 4.8 3.6 2.6 5.9 1.0
+0.30 2.8 3.9 3.6 3.3 6.0 1.4 3.6 4.6 3.5 2.4 5.9 1.0 3.1 4.9 3.6 2.5 5.9 1.0
+0.20 3.2 3.6 3.9 2.9 6.0 1.4 3.7 4.9 3.4 2.3 5.6 1.1 3.5 5.1 3.3 2.3 5.8 1.0
+0.10 3.2 4.0 4.2 3.6 4.0 2.0 3.9 5.5 3.6 2.9 3.8 1.3 3.9 5.8 4.1 3.1 3.0 1.1
+0.05 3.4 4.2 4.4 4.1 2.6 2.3 3.6 5.1 4.3 4.1 2.2 1.7 3.7 5.6 4.4 3.9 1.9 1.5
+0.00 3.3 4.3 4.5 4.1 2.5 2.3 2.7 4.2 4.7 4.7 2.5 2.2 3.0 5.4 4.9 4.2 1.8 1.7
+TABLE III: Worst-Case average ranks for CI, IBS, and NegLL metrics across all datasets. Lower number is better. Bold and underline indicates the best and second best method, respectively.
 For the exponential CPH distribution, as the failure rate increases with the $\epsilon$-perturbation, the probability of survival decreases monotonically.
-We use CROWN-IBP for a tight approximation to the upper bound on the maximization term in equation Eq.
-23.
-We evaluate the SA metrics for $\epsilon\in[0,1]$, where $\epsilon=0$ is no perturbation to input covariates, for the two perturbations methods to study how sensitive each adversarial training method is to covariate perturbation.
-Note that when the activation functions are piecewise linear, an exact solution can also be obtained by solving a Mixed Integer Linear Program (MILP).
-However, solving the MILP is computationally prohibitive: even for $\epsilon=0.1$ each MILP takes on average 9-10 seconds when tested on the Dialysis dataset, whereas CROWN-IBP takes milliseconds to calculate over all the data in the Dialysis dataset.
+We evaluate SA metrics for $\epsilon\in[0,1]$, where $\epsilon=0$ represents no perturbation to input covariates, in order to study how sensitive different adversarial training methods are to covariate perturbations.
+To simulate realistic conditions, we assume that the 10 benchmark survival analysis datasets are “clean” and contain various covariates (e.g., age, drug dosage in mg, tumor size) and survival outcomes (e.g., time until death, censored observations).
+By applying adversarial attack techniques, we perturb these input covariates to maximize the survival analysis error.
+The perturbations are designed to be subtle yet impactful, meaning they are small in magnitude but cause significant changes in the predicted survival times.
+For example, since the data is standard normalized, a perturbation of $\epsilon=0.1$ would shift an input covariate by 0.1 standard deviations from its current value.
+For instance, consider a scenario where one of the covariates is the body weight of a patient in kilograms.
+If patient $i$ has a body weight of 70 kg, an adversarial perturbation might increase this body weight to to 75 kg.
+While this change is minimal, it is selected because it may cause a drastic change in the predicted survival likelihood over the next 10 years for a non-robust NN-based SA model.
+In reality, a 5 kg change in weight should not have such a significant effect on survival, but the model may incorrectly interpret this small perturbation as a substantial change in survival hazard, highlighting its sensitivity to minor changes in input covariates.
 VI-D Results
-Furthermore, we analyze how adversarial perturbations impact the characteristics of the population survival curves.
-VI-D1 Failure Rate Distribution
-Let us assume that the input covariates are random variables $\boldsymbol{x}\sim p(\boldsymbol{x})$.
+We analyze how adversarial perturbations impact the characteristics of the population survival curves.
+VI-D1 Comparison to SOTA
+We show the average ranking of each adversarial regularization method across all datasets with respect to CI, IBS, NegLL in response to the FGSM and $\epsilon$-worst-case perturbations in Tables III and II, respectively.
+We empirically show that SAWAR through CROWN-IBP adversarial regularization leads to increased performance, consistently ranking the best across various $\epsilon$-perturbation magnitudes.
+Thus, SAWAR has better generalization performance, better calibration, and better adversarial robustness by up to 150% compared to baseline adversarial training methods and SOTA NN-based SA models.
+More detailed results, along with the exact metric values for each $\epsilon$ of the worst-case and the FGSM adversarial attacks for each dataset are available in the code repository linked in the abstract.
+VI-D2 Failure Rate Distribution
+Each instance $x$ in the dataset has its own survival curve $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$.
+All instance’s survival curves are visualized as the credible interval, where $x$ is sampled from the dataset distribution $p(x)$.
+Thus, let us assume that the input covariates are random variables $\boldsymbol{x}\sim p(\boldsymbol{x})$.
 Then the failure rate becomes a random variable that occurs from the stochasticity of the input covariates $\lambda\sim\lambda_{\boldsymbol{\theta}}(\boldsymbol{x})$.
 Moreover, we treat $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ as a stochastic process :
-$S_{\boldsymbol{\theta}}(t)\sim e^{-\lambda t}$ (25)
-Furthermore, we compute statistical quantities such as expectation, lower 5th quantile $LB=Q_{05}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x}))$, and upper 95th quantile $UB=Q_{95}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x}))$.
+$S_{\boldsymbol{\theta}}(t)\sim e^{-\lambda t}.$ (23)
+We compute the statistical quantities of $S_{\boldsymbol{\theta}}(t)$ such as expectation, lower 5th quantile $LB=Q_{05}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x}))$, and upper 95th quantile $UB=Q_{95}(\lambda_{\boldsymbol{\theta}}(\boldsymbol{x}))$.
 Fig.
 2 visualizes the credible interval and mean of the survival function stochastic process in Eq.
-25.
-The 90% credible interval of the the stochastic process changes with respect to the strength of the adversarial regularization method (Fig.
-1).
+23.
+The 90% credible interval of the stochastic process changes with respect to the strength of the adversarial regularization method (Fig.
+2).
 We observe the trend that the credible interval slightly narrows down as the strength of the adversarial regularization method increases (Fig.
-1).
+2).
+However, this trend has an exception with AAE-Cox.
+We hypothesize that since AAE-Cox is not truly an adversarial robust method (i.e., no perturbations to the input covariates during training), and by introducing more parameters via an encoder of a GAN, there are more model parameters to attack.
 The variance of the failure rate decreasing leads to a more robust model against adversarial perturbations.
@@
 Therefore, by utilizing a weighted objective in Eq.
-21, we trade-off individualized survival curves with vulnerability to adversarial perturbations and a population survival curve which is completely resilient to adversarial perturbations.
-However, despite exhibiting significantly higher adversarial robustness, the credible interval usually remains relatively unchanged compared to the DRAFT method.
-VI-D2 Population Survival Curve
+19, we introduce a trade-off between individualized survival curves with vulnerability to adversarial perturbations and a population survival curve that is completely resilient to adversarial perturbations.
+In doing so, while exhibiting significantly higher adversarial robustness, SAWAR’s credible interval remains relatively unchanged compared to the DRAFT method.
+VI-D3 Population Survival Curve
 The $\epsilon$-perturbation magnitude determines the severity of the worst-case survival curve.
 We observe the trend that the stronger the adversarial regularization method, the “closer” the $\epsilon$-worst-case population curve becomes to the unperturbed population survival curve for a given $\epsilon\in[0,1]$ (Fig.
-2).
+3).
 Moreover, we use the Kaplan Meier Curve (KMC), a commonly used frequentist approach for non-parametrically estimating the population survival curve using samples $\{t_{i},e_{i}\}_{i=1}^{N}$, to visually evaluate model calibration.
@@
 3 demonstrates the strong calibration of SAWAR model [32].
-We calculate the average ranking of each adversarial regularization method across all datasets with respect to CI, IBS, NegLL in response to the FGSM and $\epsilon$-worst-case perturbations (Tables III and II).
-We empirically showed that CROWN-IBP adversarial regularization leads to increased performance, consistently ranking the best across various $\epsilon$-perturbation magnitudes.
-More detailed results, along with the exact metric values for each $\epsilon$ of the worst-case perturbation for each dataset are available in the code repository linked in the abstract.
 VII Conclusions
-This paper introduces CROWN-IBP regularization for enhancing adversarial robustness in fully-parametric SA using NN models.
-We evaluated the SAWAR method on 10 time-to-event medical datasets, comparing it with standard adversarial regularization methods such as Gaussian noise, FGSM, and PGD, as well as the novel worst-case perturbation.
-Our empirical results show that CROWN-IBP regularization consistently improves performance in SA across various metrics, including CI, IBS, and NegLL, for all $\epsilon$-perturbation magnitudes from 0 to 1.
-Compared to conventional adversarial regularization techniques, SAWAR demonstrates superior resilience to input covariate perturbations, with relative improvements in CI and IBS ranging from 1% to 150% across all $\epsilon$-perturbation magnitudes.
-In summary, SAWAR significantly enhances predictive performance, adversarial robustness, and calibration.
-This approach effectively mitigates data uncertainty in SA and improves generalization across unseen datasets.
-Future work will include the use the Weibull CPH model to enable time varying hazard rates.
-This will require extending auto_LiRPA to handle the power operator for operations such as $t^{k}$, where $k$ is a learnable parameter.
+This paper introduces a robust training objective based on CROWN-IBP regularization to enhance adversarial robustness in fully-parametric NN-based SA models, improving both model calibration and generalization.
+We evaluated the SAWAR method on 10 time-to-event medical datasets, comparing it with standard adversarial regularization methods like Gaussian noise, FGSM, and PGD, as well as SOTA models such as DRAFT and AAE-Cox.
+Our empirical results demonstrate that SAWAR consistently improves SA performance across various metrics, including CI, IBS, and NegLL, for all $\epsilon$-perturbation magnitudes from 0 to 1.
+Compared to traditional adversarial regularization techniques, SAWAR shows superior resilience to input covariate perturbations, with relative improvements in CI and IBS ranging from 1
+In summary, SAWAR significantly enhances predictive performance, adversarial robustness, and calibration, effectively mitigating data uncertainty in SA and improving generalization on unseen datasets.
+Future work will involve extending the approach to the Weibull CPH model to enable time-varying hazard rates, which will require adapting auto_LiRPA to handle power operations such as $t^{k}$, where $k$ is a learnable parameter.
 Appendix A Appendix
-We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller.
+We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1 for the worst-case adversarial attack, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller.
 The relative percent changes in CI from the DRAFT training objective to SAWAR training objective is shown in Table IV (where higher percentage change is better).
@@
 $\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
-%$\%\Delta$ 29.46 48.18 56.27 72.23 102.44 133.35 152.17 140.33 91.78 21.17 9.63 1.07
-TABLE IV: The relative percent change in Concordance Index metric from the DRAFT training objective to the SAWAR training objective averaged across the SurvSet datasets.
+%$\%\Delta$ 72.8 79.69 92.54 90.82 81.56 66.21 41.01 12.83 3.19 1.88 1.72 1.6
+TABLE IV: The relative percent change in the Concordance Index metric from the DRAFT model to the SAWAR training objective averaged across the SurvSet datasets for the worst-case adversarial attack.
 A higher relative percent change is better.
 $\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
-Aids2 DRAFT 0.499 0.497 0.501 0.505 0.507 0.509 0.517 0.535 0.565 0.573 0.573 0.572
-Noise 0.492 0.494 0.496 0.5 0.5 0.498 0.507 0.526 0.556 0.556 0.554 0.554
-FGSM 0.498 0.495 0.498 0.497 0.5 0.502 0.508 0.529 0.554 0.557 0.555 0.553
-PGD 0.498 0.497 0.497 0.5 0.501 0.501 0.507 0.531 0.557 0.558 0.556 0.554
-SAWAR 0.567 0.575 0.577 0.579 0.58 0.58 0.579 0.579 0.579 0.579 0.579 0.579
-Framingham DRAFT 0.612 0.623 0.635 0.646 0.656 0.664 0.672 0.678 0.682 0.686 0.687 0.688
-Noise 0.606 0.616 0.626 0.635 0.642 0.649 0.659 0.674 0.683 0.685 0.685 0.685
-FGSM 0.607 0.611 0.617 0.622 0.626 0.635 0.652 0.669 0.682 0.686 0.687 0.686
-PGD 0.567 0.578 0.588 0.599 0.61 0.624 0.642 0.659 0.68 0.688 0.688 0.687
-SAWAR 0.694 0.697 0.699 0.701 0.702 0.704 0.704 0.705 0.705 0.704 0.704 0.704
-LeukSurv DRAFT 0.611 0.616 0.623 0.623 0.627 0.631 0.63 0.631 0.63 0.624 0.624 0.626
-Noise 0.552 0.557 0.562 0.566 0.569 0.574 0.587 0.61 0.628 0.627 0.621 0.617
-FGSM 0.546 0.549 0.554 0.563 0.561 0.566 0.583 0.602 0.624 0.624 0.621 0.62
-PGD 0.537 0.536 0.538 0.549 0.546 0.558 0.577 0.598 0.62 0.622 0.619 0.617
-SAWAR 0.483 0.498 0.51 0.532 0.553 0.583 0.589 0.607 0.617 0.623 0.632 0.636
-TRACE DRAFT 0.564 0.586 0.612 0.637 0.665 0.69 0.714 0.73 0.734 0.733 0.735 0.735
-Noise 0.565 0.588 0.613 0.64 0.668 0.692 0.716 0.73 0.733 0.735 0.734 0.735
-FGSM 0.553 0.577 0.603 0.63 0.658 0.688 0.713 0.722 0.727 0.73 0.733 0.734
-PGD 0.557 0.582 0.609 0.636 0.664 0.696 0.723 0.731 0.733 0.732 0.733 0.733
-SAWAR 0.73 0.733 0.736 0.738 0.739 0.739 0.74 0.74 0.739 0.738 0.738 0.737
-dataDIVAT1 DRAFT 0.534 0.558 0.583 0.606 0.624 0.639 0.65 0.658 0.662 0.664 0.664 0.664
-Noise 0.534 0.539 0.551 0.571 0.587 0.612 0.636 0.656 0.654 0.64 0.635 0.631
-FGSM 0.575 0.585 0.594 0.606 0.617 0.628 0.638 0.644 0.651 0.643 0.638 0.635
-PGD 0.517 0.529 0.548 0.568 0.591 0.618 0.636 0.654 0.657 0.651 0.647 0.644
-SAWAR 0.627 0.642 0.652 0.658 0.659 0.659 0.659 0.658 0.657 0.656 0.656 0.655
-flchain DRAFT 0.096 0.098 0.102 0.107 0.114 0.127 0.156 0.461 0.9 0.916 0.922 0.924
-Noise 0.096 0.099 0.104 0.112 0.127 0.164 0.335 0.768 0.912 0.921 0.923 0.923
-FGSM 0.102 0.116 0.159 0.288 0.599 0.852 0.91 0.917 0.923 0.927 0.927 0.925
-PGD 0.102 0.127 0.235 0.538 0.824 0.904 0.917 0.921 0.925 0.929 0.929 0.929
-SAWAR 0.84 0.925 0.927 0.929 0.93 0.93 0.929 0.929 0.93 0.93 0.931 0.931
-prostate DRAFT 0.412 0.419 0.426 0.447 0.463 0.49 0.53 0.581 0.614 0.635 0.639 0.643
-Noise 0.49 0.503 0.514 0.524 0.533 0.541 0.553 0.566 0.576 0.572 0.577 0.576
-FGSM 0.514 0.522 0.526 0.536 0.541 0.549 0.558 0.573 0.582 0.596 0.596 0.596
-PGD 0.522 0.526 0.532 0.535 0.54 0.548 0.557 0.567 0.576 0.593 0.591 0.593
-SAWAR 0.575 0.584 0.598 0.6 0.608 0.632 0.638 0.638 0.639 0.641 0.642 0.644
-retinopathy DRAFT 0.525 0.539 0.562 0.569 0.587 0.592 0.592 0.599 0.605 0.601 0.604 0.605
-Noise 0.526 0.545 0.561 0.57 0.581 0.588 0.588 0.6 0.603 0.6 0.604 0.606
-FGSM 0.47 0.485 0.51 0.536 0.569 0.578 0.591 0.6 0.599 0.602 0.606 0.607
-PGD 0.471 0.481 0.509 0.53 0.567 0.577 0.591 0.599 0.598 0.601 0.606 0.607
-SAWAR 0.588 0.597 0.611 0.616 0.619 0.615 0.621 0.62 0.62 0.619 0.617 0.617
-stagec DRAFT 0.442 0.457 0.476 0.495 0.514 0.548 0.582 0.668 0.721 0.697 0.688 0.688
-Noise 0.471 0.481 0.505 0.49 0.514 0.534 0.582 0.697 0.721 0.712 0.688 0.697
-FGSM 0.423 0.438 0.447 0.471 0.51 0.567 0.62 0.663 0.688 0.678 0.688 0.688
-PGD 0.462 0.486 0.471 0.495 0.51 0.567 0.611 0.668 0.692 0.688 0.697 0.697
-SAWAR 0.428 0.447 0.481 0.514 0.567 0.639 0.673 0.692 0.702 0.692 0.692 0.707
-zinc DRAFT 0.339 0.349 0.356 0.376 0.427 0.507 0.657 0.758 0.796 0.796 0.792 0.793
-Noise 0.315 0.329 0.339 0.376 0.424 0.545 0.704 0.778 0.791 0.805 0.806 0.81
-FGSM 0.339 0.357 0.384 0.439 0.55 0.657 0.738 0.776 0.787 0.804 0.809 0.812
-PGD 0.333 0.351 0.384 0.439 0.55 0.671 0.738 0.776 0.79 0.8 0.809 0.808
-SAWAR 0.383 0.464 0.533 0.649 0.707 0.746 0.766 0.779 0.789 0.787 0.795 0.8
+Dataset Algorithm
+Aids2 DRAFT 0.516 0.519 0.522 0.525 0.527 0.53 0.536 0.551 0.566 0.572 0.572 0.57
+Noise 0.513 0.515 0.518 0.521 0.525 0.529 0.535 0.548 0.568 0.571 0.57 0.569
+FGSM 0.512 0.516 0.52 0.523 0.526 0.532 0.537 0.55 0.565 0.569 0.568 0.566
+PGD 0.506 0.509 0.513 0.518 0.523 0.528 0.535 0.552 0.567 0.569 0.568 0.567
+AAE-Cox 0.504 0.508 0.507 0.506 0.505 0.501 0.499 0.504 0.51 0.561 0.573 0.573
+SAWAR 0.565 0.568 0.568 0.569 0.569 0.57 0.57 0.57 0.57 0.57 0.569 0.569
+Framingham DRAFT 0.618 0.634 0.649 0.663 0.676 0.688 0.699 0.708 0.716 0.72 0.721 0.722
+Noise 0.606 0.622 0.637 0.653 0.666 0.679 0.691 0.703 0.713 0.718 0.719 0.719
+FGSM 0.582 0.598 0.611 0.626 0.641 0.655 0.671 0.691 0.708 0.715 0.715 0.715
+PGD 0.539 0.559 0.582 0.605 0.627 0.647 0.669 0.693 0.71 0.717 0.717 0.717
+AAE-Cox 0.514 0.522 0.529 0.533 0.539 0.553 0.569 0.57 0.582 0.73 0.733 0.733
+SAWAR 0.721 0.725 0.729 0.731 0.733 0.734 0.735 0.736 0.736 0.737 0.737 0.737
+LeukSurv DRAFT 0.589 0.593 0.596 0.598 0.601 0.602 0.608 0.62 0.63 0.632 0.633 0.633
+Noise 0.545 0.548 0.545 0.547 0.551 0.56 0.581 0.609 0.63 0.628 0.626 0.624
+FGSM 0.511 0.514 0.52 0.525 0.535 0.549 0.579 0.612 0.637 0.638 0.636 0.635
+PGD 0.498 0.498 0.501 0.51 0.521 0.543 0.576 0.616 0.64 0.643 0.64 0.638
+AAE-Cox 0.559 0.552 0.542 0.555 0.548 0.55 0.548 0.537 0.545 0.631 0.658 0.656
+SAWAR 0.495 0.51 0.524 0.536 0.552 0.57 0.588 0.609 0.633 0.658 0.669 0.673
+TRACE DRAFT 0.581 0.605 0.633 0.66 0.685 0.708 0.726 0.736 0.741 0.743 0.744 0.745
+Noise 0.576 0.598 0.621 0.645 0.669 0.691 0.712 0.728 0.737 0.742 0.744 0.745
+FGSM 0.581 0.603 0.629 0.653 0.674 0.697 0.717 0.732 0.739 0.743 0.744 0.744
+PGD 0.571 0.595 0.621 0.646 0.668 0.691 0.714 0.73 0.739 0.743 0.744 0.745
+AAE-Cox 0.432 0.44 0.447 0.451 0.46 0.47 0.489 0.524 0.628 0.743 0.746 0.747
+SAWAR 0.714 0.722 0.728 0.733 0.735 0.737 0.739 0.742 0.744 0.746 0.747 0.747
+dataDIVAT1 DRAFT 0.573 0.585 0.598 0.611 0.625 0.636 0.646 0.653 0.657 0.657 0.656 0.655
+Noise 0.532 0.541 0.552 0.562 0.575 0.589 0.603 0.619 0.631 0.636 0.635 0.635
+FGSM 0.546 0.555 0.567 0.578 0.588 0.6 0.612 0.626 0.641 0.647 0.646 0.646
+PGD 0.523 0.534 0.545 0.557 0.569 0.581 0.599 0.621 0.641 0.648 0.648 0.648
+AAE-Cox 0.597 0.589 0.589 0.585 0.58 0.571 0.571 0.566 0.596 0.662 0.663 0.663
+SAWAR 0.64 0.645 0.649 0.654 0.658 0.661 0.664 0.665 0.667 0.668 0.669 0.67
+flchain DRAFT 0.109 0.111 0.115 0.122 0.133 0.158 0.239 0.566 0.905 0.917 0.92 0.921
+Noise 0.166 0.131 0.111 0.123 0.149 0.224 0.433 0.792 0.911 0.917 0.919 0.918
+FGSM 0.115 0.153 0.245 0.431 0.72 0.885 0.91 0.914 0.918 0.922 0.922 0.922
+PGD 0.165 0.268 0.457 0.744 0.876 0.904 0.912 0.917 0.92 0.923 0.923 0.923
+AAE-Cox 0.527 0.648 0.668 0.553 0.394 0.511 0.51 0.476 0.11 0.179 0.925 0.926
+SAWAR 0.593 0.684 0.866 0.918 0.922 0.925 0.926 0.927 0.927 0.927 0.927 0.927
+prostate DRAFT 0.402 0.416 0.428 0.444 0.469 0.505 0.542 0.585 0.627 0.653 0.661 0.668
+Noise 0.433 0.44 0.448 0.451 0.463 0.48 0.492 0.511 0.537 0.558 0.561 0.562
+FGSM 0.448 0.453 0.46 0.465 0.475 0.486 0.502 0.515 0.541 0.564 0.569 0.571
+PGD 0.447 0.455 0.459 0.467 0.474 0.487 0.509 0.53 0.564 0.585 0.588 0.588
+AAE-Cox 0.41 0.411 0.406 0.41 0.407 0.403 0.399 0.391 0.414 0.646 0.686 0.691
+SAWAR 0.608 0.623 0.637 0.652 0.665 0.671 0.672 0.673 0.667 0.663 0.66 0.657
+retinopathy DRAFT 0.553 0.568 0.578 0.597 0.616 0.632 0.645 0.653 0.663 0.666 0.667 0.669
+Noise 0.573 0.591 0.605 0.62 0.632 0.644 0.653 0.661 0.666 0.667 0.668 0.668
+FGSM 0.575 0.592 0.604 0.615 0.625 0.633 0.642 0.647 0.651 0.656 0.657 0.659
+PGD 0.571 0.589 0.599 0.61 0.621 0.63 0.64 0.647 0.651 0.655 0.656 0.657
+AAE-Cox 0.494 0.495 0.506 0.504 0.514 0.564 0.577 0.592 0.62 0.657 0.652 0.648
+SAWAR 0.668 0.669 0.67 0.666 0.662 0.66 0.656 0.654 0.653 0.65 0.648 0.647
+stagec DRAFT 0.358 0.378 0.382 0.406 0.425 0.449 0.454 0.475 0.489 0.504 0.507 0.512
+Noise 0.353 0.373 0.39 0.407 0.436 0.466 0.485 0.5 0.53 0.544 0.549 0.555
+FGSM 0.329 0.346 0.368 0.389 0.416 0.429 0.443 0.471 0.488 0.502 0.503 0.51
+PGD 0.341 0.352 0.381 0.399 0.419 0.431 0.442 0.47 0.49 0.496 0.498 0.505
+AAE-Cox 0.393 0.397 0.397 0.411 0.417 0.409 0.405 0.429 0.469 0.523 0.543 0.54
+SAWAR 0.393 0.401 0.413 0.425 0.436 0.461 0.488 0.506 0.534 0.543 0.548 0.558
+zinc DRAFT 0.262 0.27 0.284 0.296 0.317 0.355 0.44 0.579 0.712 0.755 0.765 0.77
+Noise 0.318 0.328 0.343 0.366 0.401 0.466 0.562 0.661 0.734 0.766 0.773 0.776
+FGSM 0.377 0.403 0.439 0.49 0.557 0.626 0.685 0.737 0.762 0.78 0.782 0.783
+PGD 0.384 0.406 0.443 0.495 0.559 0.632 0.695 0.741 0.77 0.778 0.78 0.781
+AAE-Cox 0.226 0.231 0.228 0.236 0.234 0.249 0.265 0.306 0.445 0.724 0.754 0.759
+SAWAR 0.667 0.725 0.763 0.785 0.785 0.785 0.781 0.778 0.778 0.776 0.776 0.774
 TABLE V: Concordance Index metric for SurvSet datasets (higher is better) for each adversarial training method against the worst-case adversarial attack.
-We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller.
+We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1 for the worst-case adversarial attack, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller.
 The relative percent changes in Integrated Brier Score metric from the DRAFT training objective to SAWAR training objective is shown in Table VI (where lower percentage change is better).
@@
 $\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
-%$\%\Delta$44.13 -45.46 -46.68 -47.57 -47.86 -47.10 -45.03 -40.89 -33.06 -20.29 -11.34 -1.68
-TABLE VI: The relative percent change in Integrated Brier Score metric from the DRAFT training objective to the SAWAR training objective averaged across the SurvSet datasets.
+%$\%\Delta$37.55 -42.51 -48.08 -53.15 -55.48 -55.34 -51.51 -42.94 -28.8 -10.87 -4.78 -0.65
+TABLE VI: The relative percent change in Integrated Brier Score metric from the DRAFT model to the SAWAR training objective averaged across the SurvSet datasets for the worst-case adversarial attack.
 A lower relative percent change is better.
 $\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
-Aids2 DRAFT 0.24 0.233 0.224 0.212 0.197 0.178 0.157 0.137 0.124 0.12 0.121 0.122
-Noise 0.254 0.253 0.25 0.245 0.236 0.22 0.196 0.161 0.13 0.121 0.121 0.122
-FGSM 0.252 0.249 0.244 0.236 0.223 0.203 0.176 0.145 0.125 0.121 0.121 0.121
-PGD 0.252 0.248 0.243 0.234 0.221 0.2 0.172 0.143 0.124 0.121 0.121 0.121
-SAWAR 0.126 0.124 0.124 0.124 0.123 0.123 0.123 0.123 0.123 0.123 0.123 0.124
-Framingham DRAFT 0.81 0.8 0.778 0.733 0.65 0.52 0.369 0.243 0.165 0.128 0.119 0.115
-Noise 0.818 0.818 0.816 0.809 0.778 0.681 0.495 0.294 0.174 0.128 0.119 0.116
-FGSM 0.818 0.816 0.81 0.791 0.728 0.58 0.369 0.208 0.14 0.12 0.116 0.115
-PGD 0.818 0.818 0.817 0.811 0.784 0.683 0.461 0.235 0.138 0.117 0.115 0.115
-SAWAR 0.165 0.152 0.141 0.133 0.127 0.122 0.118 0.116 0.114 0.114 0.113 0.113
-LeukSurv DRAFT 0.163 0.163 0.163 0.163 0.162 0.161 0.156 0.144 0.126 0.109 0.105 0.103
-Noise 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.157 0.134 0.12 0.114
-FGSM 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.161 0.148 0.122 0.114 0.111
-PGD 0.163 0.163 0.163 0.163 0.163 0.163 0.163 0.16 0.145 0.12 0.114 0.111
-SAWAR 0.154 0.15 0.144 0.136 0.128 0.12 0.115 0.112 0.11 0.107 0.107 0.106
-TRACE DRAFT 0.634 0.634 0.633 0.626 0.603 0.549 0.449 0.329 0.234 0.183 0.172 0.167
-Noise 0.634 0.634 0.632 0.622 0.596 0.536 0.435 0.319 0.23 0.183 0.172 0.167
-FGSM 0.634 0.634 0.634 0.631 0.613 0.555 0.434 0.294 0.212 0.177 0.171 0.168
-PGD 0.634 0.634 0.632 0.621 0.58 0.491 0.362 0.245 0.189 0.171 0.167 0.166
-SAWAR 0.282 0.254 0.231 0.213 0.198 0.186 0.177 0.17 0.166 0.163 0.163 0.163
-dataDIVAT1 DRAFT 0.721 0.702 0.67 0.616 0.531 0.417 0.308 0.234 0.195 0.18 0.178 0.178
-Noise 0.749 0.749 0.748 0.748 0.743 0.72 0.626 0.413 0.248 0.198 0.192 0.192
-FGSM 0.749 0.749 0.748 0.747 0.74 0.699 0.547 0.322 0.215 0.193 0.192 0.192
-PGD 0.749 0.749 0.749 0.748 0.746 0.724 0.582 0.305 0.192 0.184 0.186 0.187
-SAWAR 0.266 0.241 0.224 0.212 0.202 0.193 0.187 0.183 0.18 0.178 0.178 0.178
-flchain DRAFT 0.831 0.831 0.831 0.831 0.831 0.83 0.829 0.818 0.486 0.096 0.07 0.052
-Noise 0.831 0.831 0.831 0.831 0.831 0.83 0.829 0.807 0.25 0.087 0.063 0.052
-FGSM 0.831 0.831 0.831 0.83 0.83 0.828 0.707 0.146 0.094 0.06 0.054 0.051
-PGD 0.831 0.831 0.831 0.83 0.83 0.797 0.298 0.114 0.079 0.055 0.051 0.05
-SAWAR 0.522 0.161 0.076 0.063 0.057 0.054 0.052 0.051 0.05 0.05 0.049 0.049
-prostate DRAFT 0.53 0.529 0.528 0.526 0.519 0.505 0.475 0.418 0.335 0.256 0.231 0.218
-Noise 0.53 0.53 0.53 0.53 0.53 0.53 0.53 0.528 0.477 0.371 0.325 0.296
-FGSM 0.53 0.53 0.53 0.53 0.53 0.53 0.529 0.509 0.413 0.29 0.257 0.246
-PGD 0.53 0.53 0.53 0.53 0.53 0.53 0.528 0.502 0.396 0.28 0.254 0.244
-SAWAR 0.508 0.482 0.441 0.393 0.348 0.309 0.278 0.251 0.229 0.214 0.209 0.205
-retinopathy DRAFT 0.662 0.657 0.645 0.621 0.582 0.522 0.43 0.324 0.244 0.203 0.196 0.195
-Noise 0.662 0.657 0.645 0.622 0.583 0.523 0.432 0.326 0.245 0.204 0.196 0.195
-FGSM 0.656 0.646 0.626 0.597 0.556 0.493 0.401 0.301 0.232 0.201 0.196 0.197
-PGD 0.656 0.645 0.625 0.596 0.554 0.49 0.397 0.299 0.231 0.201 0.197 0.198
-SAWAR 0.526 0.484 0.436 0.385 0.334 0.29 0.254 0.228 0.211 0.203 0.202 0.202
-stagec DRAFT 0.505 0.505 0.505 0.504 0.502 0.486 0.431 0.329 0.211 0.175 0.181 0.192
-Noise 0.505 0.505 0.505 0.504 0.502 0.485 0.432 0.338 0.216 0.175 0.181 0.192
-FGSM 0.505 0.505 0.505 0.503 0.494 0.458 0.382 0.268 0.174 0.161 0.169 0.178
-PGD 0.505 0.505 0.505 0.503 0.493 0.454 0.377 0.261 0.169 0.159 0.167 0.176
-SAWAR 0.491 0.476 0.45 0.408 0.354 0.296 0.24 0.195 0.168 0.16 0.162 0.167
-zinc DRAFT 0.907 0.907 0.905 0.899 0.882 0.823 0.652 0.368 0.18 0.109 0.096 0.09
-Noise 0.908 0.907 0.907 0.904 0.893 0.838 0.625 0.32 0.168 0.109 0.096 0.089
-FGSM 0.907 0.907 0.904 0.892 0.839 0.659 0.365 0.194 0.124 0.095 0.089 0.086
-PGD 0.907 0.907 0.903 0.89 0.832 0.64 0.348 0.187 0.122 0.094 0.088 0.087
-SAWAR 0.72 0.609 0.476 0.352 0.26 0.198 0.159 0.132 0.113 0.1 0.096 0.094
+Dataset Algorithm
+Aids2 DRAFT 0.265 0.262 0.258 0.252 0.242 0.228 0.207 0.18 0.151 0.138 0.137 0.137
+Noise 0.266 0.264 0.261 0.255 0.246 0.231 0.209 0.179 0.15 0.138 0.137 0.138
+FGSM 0.265 0.263 0.258 0.251 0.239 0.221 0.197 0.167 0.144 0.137 0.137 0.138
+PGD 0.265 0.262 0.257 0.248 0.235 0.216 0.19 0.161 0.141 0.137 0.138 0.138
+AAE-Cox 0.269 0.269 0.269 0.268 0.265 0.259 0.25 0.231 0.183 0.138 0.137 0.137
+SAWAR 0.14 0.139 0.138 0.138 0.137 0.137 0.137 0.137 0.137 0.137 0.137 0.137
+Framingham DRAFT 0.831 0.825 0.811 0.783 0.724 0.618 0.459 0.289 0.174 0.124 0.114 0.11
+Noise 0.836 0.834 0.83 0.82 0.79 0.709 0.543 0.331 0.185 0.126 0.115 0.11
+FGSM 0.836 0.836 0.835 0.831 0.817 0.765 0.605 0.336 0.162 0.118 0.113 0.112
+PGD 0.836 0.836 0.836 0.831 0.812 0.74 0.555 0.295 0.146 0.115 0.113 0.112
+AAE-Cox 0.836 0.836 0.836 0.836 0.836 0.836 0.834 0.775 0.385 0.116 0.109 0.108
+SAWAR 0.17 0.155 0.144 0.134 0.127 0.122 0.117 0.114 0.112 0.111 0.111 0.111
+LeukSurv DRAFT 0.206 0.206 0.205 0.205 0.203 0.2 0.195 0.185 0.17 0.158 0.154 0.153
+Noise 0.206 0.206 0.206 0.206 0.206 0.206 0.206 0.206 0.2 0.171 0.166 0.173
+FGSM 0.206 0.206 0.206 0.206 0.206 0.206 0.206 0.203 0.187 0.161 0.158 0.159
+PGD 0.206 0.206 0.206 0.206 0.206 0.206 0.206 0.202 0.183 0.16 0.157 0.158
+AAE-Cox 0.206 0.206 0.206 0.206 0.206 0.206 0.206 0.206 0.198 0.162 0.154 0.152
+SAWAR 0.194 0.19 0.184 0.178 0.171 0.165 0.16 0.155 0.151 0.148 0.147 0.146
+TRACE DRAFT 0.62 0.611 0.593 0.558 0.504 0.432 0.347 0.265 0.204 0.172 0.165 0.162
+Noise 0.627 0.626 0.623 0.611 0.579 0.518 0.426 0.317 0.225 0.176 0.166 0.162
+FGSM 0.627 0.625 0.619 0.596 0.545 0.46 0.351 0.248 0.188 0.167 0.163 0.162
+PGD 0.627 0.625 0.617 0.593 0.54 0.455 0.347 0.247 0.188 0.167 0.164 0.162
+AAE-Cox 0.627 0.627 0.627 0.627 0.627 0.623 0.588 0.471 0.284 0.172 0.165 0.163
+SAWAR 0.324 0.281 0.246 0.22 0.202 0.188 0.179 0.171 0.166 0.163 0.162 0.161
+dataDIVAT1 DRAFT 0.702 0.689 0.663 0.614 0.527 0.405 0.285 0.21 0.177 0.17 0.172 0.177
+Noise 0.72 0.72 0.719 0.718 0.713 0.689 0.599 0.4 0.245 0.197 0.193 0.195
+FGSM 0.719 0.719 0.718 0.714 0.697 0.637 0.478 0.285 0.2 0.182 0.182 0.184
+PGD 0.719 0.719 0.716 0.71 0.687 0.613 0.442 0.261 0.191 0.18 0.181 0.182
+AAE-Cox 0.72 0.72 0.72 0.72 0.72 0.72 0.715 0.553 0.194 0.165 0.17 0.174
+SAWAR 0.189 0.184 0.181 0.179 0.177 0.177 0.177 0.177 0.177 0.178 0.179 0.179
+flchain DRAFT 0.816 0.816 0.816 0.816 0.816 0.816 0.814 0.797 0.46 0.093 0.069 0.054
+Noise 0.816 0.816 0.816 0.816 0.816 0.816 0.814 0.77 0.227 0.088 0.067 0.057
+FGSM 0.816 0.816 0.816 0.816 0.816 0.804 0.468 0.119 0.089 0.061 0.055 0.054
+PGD 0.817 0.817 0.817 0.816 0.81 0.638 0.162 0.107 0.076 0.057 0.054 0.053
+AAE-Cox nan nan nan nan nan nan nan nan 0.794 0.688 0.054 0.051
+SAWAR 0.445 0.363 0.241 0.094 0.067 0.059 0.057 0.055 0.054 0.053 0.053 0.053
+prostate DRAFT 0.517 0.517 0.516 0.513 0.508 0.495 0.466 0.408 0.322 0.234 0.202 0.181
+Noise 0.518 0.518 0.519 0.519 0.52 0.52 0.519 0.508 0.457 0.321 0.267 0.249
+FGSM 0.517 0.518 0.519 0.518 0.52 0.52 0.517 0.493 0.402 0.262 0.228 0.218
+PGD 0.518 0.518 0.518 0.518 0.517 0.518 0.515 0.486 0.383 0.248 0.219 0.209
+AAE-Cox 0.519 0.519 0.519 0.519 0.518 0.518 0.517 0.511 0.443 0.205 0.173 0.169
+SAWAR 0.37 0.334 0.291 0.262 0.236 0.218 0.204 0.192 0.184 0.178 0.176 0.175
+retinopathy DRAFT 0.728 0.722 0.71 0.687 0.647 0.579 0.48 0.364 0.266 0.204 0.188 0.179
+Noise 0.73 0.725 0.714 0.693 0.652 0.579 0.472 0.353 0.256 0.199 0.184 0.177
+FGSM 0.725 0.715 0.697 0.665 0.61 0.526 0.419 0.314 0.237 0.196 0.186 0.182
+PGD 0.724 0.714 0.695 0.662 0.606 0.521 0.413 0.309 0.235 0.195 0.186 0.182
+AAE-Cox 0.733 0.733 0.733 0.733 0.733 0.732 0.716 0.596 0.281 0.183 0.181 0.181
+SAWAR 0.588 0.548 0.499 0.44 0.378 0.317 0.265 0.224 0.197 0.184 0.181 0.181
+stagec DRAFT 0.556 0.549 0.544 0.543 0.539 0.517 0.468 0.404 0.332 0.274 0.256 0.245
+Noise 0.559 0.553 0.547 0.548 0.551 0.541 0.501 0.437 0.357 0.282 0.258 0.244
+FGSM 0.547 0.545 0.549 0.553 0.545 0.511 0.457 0.391 0.323 0.278 0.265 0.258
+PGD 0.547 0.546 0.549 0.553 0.542 0.507 0.453 0.388 0.322 0.278 0.267 0.26
+AAE-Cox 0.568 0.568 0.568 0.568 0.567 0.559 0.545 0.544 0.474 0.291 0.268 0.263
+SAWAR 0.521 0.497 0.469 0.436 0.401 0.365 0.334 0.303 0.276 0.256 0.249 0.243
+zinc DRAFT 0.847 0.847 0.846 0.843 0.831 0.787 0.641 0.376 0.184 0.119 0.109 0.107
+Noise 0.847 0.847 0.847 0.845 0.838 0.795 0.627 0.351 0.179 0.128 0.118 0.113
+FGSM 0.847 0.847 0.845 0.835 0.782 0.611 0.356 0.189 0.133 0.116 0.113 0.112
+PGD 0.847 0.847 0.845 0.834 0.777 0.6 0.341 0.182 0.132 0.116 0.113 0.112
+AAE-Cox 0.847 0.847 0.847 0.847 0.847 0.847 0.846 0.823 0.507 0.119 0.109 0.108
+SAWAR 0.654 0.532 0.401 0.289 0.21 0.161 0.133 0.118 0.111 0.109 0.109 0.109
 TABLE VII: Integrated Brier Score metric for SurvSet datasets (lower is better) for each adversarial training method against the worst-case adversarial attack.
 $\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
-Aids2 DRAFT 1.14e+04 7.22e+03 4.60e+03 2.87e+03 1.80e+03 1.16e+03 7.95e+02 6.06e+02 5.32e+02 5.16e+02 5.17e+02 5.19e+02
-Noise 3.84e+05 1.61e+05 6.68e+04 2.78e+04 1.14e+04 4.42e+03 1.83e+03 8.54e+02 5.64e+02 5.20e+02 5.19e+02 5.21e+02
-FGSM 6.77e+04 3.50e+04 1.78e+04 8.89e+03 4.35e+03 2.14e+03 1.10e+03 6.66e+02 5.35e+02 5.19e+02 5.19e+02 5.20e+02
-PGD 5.49e+04 2.86e+04 1.49e+04 7.58e+03 3.79e+03 1.92e+03 1.02e+03 6.42e+02 5.31e+02 5.19e+02 5.19e+02 5.19e+02
-SAWAR 5.29e+02 5.24e+02 5.22e+02 5.22e+02 5.21e+02 5.21e+02 5.21e+02 5.21e+02 5.21e+02 5.22e+02 5.22e+02 5.22e+02
-Framingham DRAFT 1.47e+05 7.07e+04 3.42e+04 1.68e+04 8.47e+03 4.57e+03 2.79e+03 2.01e+03 1.68e+03 1.54e+03 1.51e+03 1.50e+03
-Noise 1.27e+07 3.17e+06 8.03e+05 2.06e+05 5.38e+04 1.48e+04 4.92e+03 2.39e+03 1.72e+03 1.54e+03 1.51e+03 1.50e+03
-FGSM 1.95e+06 5.96e+05 1.84e+05 5.70e+04 1.81e+04 6.33e+03 2.84e+03 1.85e+03 1.58e+03 1.51e+03 1.50e+03 1.49e+03
-PGD 2.01e+07 4.36e+06 9.57e+05 2.13e+05 4.88e+04 1.21e+04 3.85e+03 1.97e+03 1.57e+03 1.50e+03 1.49e+03 1.49e+03
-SAWAR 1.66e+03 1.62e+03 1.58e+03 1.56e+03 1.53e+03 1.52e+03 1.50e+03 1.50e+03 1.49e+03 1.49e+03 1.49e+03 1.49e+03
-LeukSurv DRAFT 1.79e+07 4.80e+06 1.29e+06 3.48e+05 9.29e+04 2.53e+04 6.59e+03 1.77e+03 5.76e+02 3.10e+02 2.74e+02 2.58e+02
-Noise 5.46e+23 1.79e+21 7.98e+18 2.52e+16 8.53e+13 1.88e+11 5.25e+08 2.35e+06 3.37e+04 2.60e+03 1.21e+03 7.21e+02
-FGSM 8.34e+16 1.79e+15 3.10e+13 4.86e+11 8.13e+09 1.49e+08 2.72e+06 7.03e+04 3.45e+03 6.20e+02 4.31e+02 3.59e+02
-PGD 1.28e+15 3.89e+13 1.03e+12 2.67e+10 7.96e+08 2.30e+07 7.01e+05 2.80e+04 1.98e+03 4.65e+02 3.58e+02 3.17e+02
-SAWAR 2.07e+04 8.38e+03 3.50e+03 1.45e+03 6.62e+02 3.45e+02 2.91e+02 2.70e+02 2.60e+02 2.54e+02 2.52e+02 2.50e+02
-TRACE DRAFT 4.56e+09 4.41e+08 4.44e+07 4.61e+06 5.31e+05 6.71e+04 9.88e+03 2.20e+03 9.05e+02 6.03e+02 5.53e+02 5.32e+02
-Noise 9.11e+08 1.07e+08 1.34e+07 1.74e+06 2.43e+05 3.71e+04 6.63e+03 1.78e+03 8.35e+02 5.93e+02 5.51e+02 5.32e+02
-FGSM 2.01e+11 9.10e+09 4.31e+08 2.24e+07 1.26e+06 8.26e+04 7.79e+03 1.61e+03 7.45e+02 5.74e+02 5.47e+02 5.35e+02
-PGD 4.53e+08 4.81e+07 5.35e+06 6.12e+05 7.63e+04 1.07e+04 2.09e+03 8.37e+02 5.94e+02 5.40e+02 5.31e+02 5.26e+02
-SAWAR 8.79e+02 7.60e+02 6.82e+02 6.29e+02 5.94e+02 5.69e+02 5.53e+02 5.41e+02 5.33e+02 5.28e+02 5.27e+02 5.27e+02
-dataDIVAT1 DRAFT 1.43e+04 8.36e+03 4.95e+03 3.01e+03 1.91e+03 1.31e+03 9.99e+02 8.51e+02 7.84e+02 7.56e+02 7.51e+02 7.49e+02
-Noise 5.69e+08 7.64e+07 1.05e+07 1.44e+06 2.07e+05 3.17e+04 5.75e+03 1.61e+03 9.08e+02 7.81e+02 7.67e+02 7.64e+02
-FGSM 2.48e+08 3.36e+07 4.65e+06 6.56e+05 9.64e+04 1.54e+04 3.12e+03 1.14e+03 8.16e+02 7.65e+02 7.61e+02 7.60e+02
-PGD 1.85e+11 8.02e+09 3.47e+08 1.75e+07 9.39e+05 6.14e+04 5.80e+03 1.27e+03 8.03e+02 7.57e+02 7.56e+02 7.57e+02
-SAWAR 9.18e+02 8.55e+02 8.21e+02 8.00e+02 7.86e+02 7.75e+02 7.66e+02 7.59e+02 7.55e+02 7.52e+02 7.51e+02 7.51e+02
-flchain DRAFT 3.35e+23 9.34e+20 2.56e+18 7.37e+15 2.23e+13 6.82e+10 2.29e+08 9.33e+05 1.67e+04 2.03e+03 1.25e+03 1.09e+03
-Noise 2.18e+35 1.20e+31 7.58e+26 4.68e+22 3.48e+18 2.65e+14 2.21e+10 6.77e+06 4.45e+04 1.96e+03 1.22e+03 1.11e+03
-FGSM 1.61e+20 3.47e+17 7.79e+14 1.86e+12 5.49e+09 4.02e+07 6.37e+05 1.86e+04 2.15e+03 1.15e+03 1.09e+03 1.08e+03
-PGD 1.30e+16 7.91e+13 4.93e+11 3.47e+09 4.60e+07 1.33e+06 4.93e+04 4.99e+03 1.45e+03 1.10e+03 1.08e+03 1.07e+03
-SAWAR 6.31e+03 2.22e+03 1.43e+03 1.19e+03 1.12e+03 1.09e+03 1.08e+03 1.07e+03 1.07e+03 1.07e+03 1.07e+03 1.07e+03
-prostate DRAFT 9.74e+05 3.22e+05 1.06e+05 3.58e+04 1.21e+04 4.25e+03 1.64e+03 7.62e+02 4.66e+02 3.70e+02 3.51e+02 3.44e+02
-Noise 7.00e+20 5.05e+18 3.75e+16 2.85e+14 2.28e+12 2.03e+10 2.59e+08 9.04e+06 4.85e+05 6.66e+04 2.94e+04 1.37e+04
-FGSM 3.16e+14 1.04e+13 3.40e+11 1.16e+10 4.07e+08 1.52e+07 6.47e+05 3.56e+04 3.75e+03 1.03e+03 7.20e+02 5.82e+02
-PGD 8.84e+12 4.29e+11 2.18e+10 1.14e+09 5.94e+07 3.27e+06 1.92e+05 1.44e+04 1.91e+03 6.40e+02 5.00e+02 4.39e+02
-SAWAR 9.65e+03 3.88e+03 1.66e+03 8.48e+02 5.43e+02 4.30e+02 3.83e+02 3.59e+02 3.45e+02 3.37e+02 3.35e+02 3.34e+02
-retinopathy DRAFT 1.11e+04 5.64e+03 2.88e+03 1.51e+03 8.08e+02 4.64e+02 2.98e+02 2.18e+02 1.84e+02 1.70e+02 1.68e+02 1.67e+02
-Noise 1.11e+04 5.61e+03 2.88e+03 1.51e+03 8.13e+02 4.68e+02 2.99e+02 2.19e+02 1.84e+02 1.70e+02 1.68e+02 1.68e+02
-FGSM 7.43e+03 3.89e+03 2.07e+03 1.12e+03 6.33e+02 3.86e+02 2.63e+02 2.03e+02 1.78e+02 1.68e+02 1.67e+02 1.68e+02
-PGD 7.30e+03 3.83e+03 2.04e+03 1.11e+03 6.25e+02 3.81e+02 2.60e+02 2.02e+02 1.77e+02 1.68e+02 1.67e+02 1.68e+02
-SAWAR 4.02e+02 3.31e+02 2.79e+02 2.41e+02 2.14e+02 1.95e+02 1.82e+02 1.74e+02 1.69e+02 1.67e+02 1.67e+02 1.67e+02
-stagec DRAFT 1.16e+06 2.94e+05 7.50e+04 1.85e+04 4.81e+03 1.28e+03 3.57e+02 1.14e+02 5.35e+01 4.10e+01 3.98e+01 3.97e+01
-Noise 2.07e+06 4.93e+05 1.14e+05 2.73e+04 6.49e+03 1.61e+03 4.19e+02 1.25e+02 5.50e+01 4.10e+01 3.96e+01 3.95e+01
-FGSM 1.80e+05 5.45e+04 1.67e+04 5.23e+03 1.64e+03 5.30e+02 1.77e+02 7.43e+01 4.54e+01 3.96e+01 3.91e+01 3.92e+01
-PGD 1.54e+05 4.77e+04 1.47e+04 4.63e+03 1.47e+03 4.84e+02 1.69e+02 7.23e+01 4.48e+01 3.93e+01 3.89e+01 3.90e+01
-SAWAR 1.03e+03 5.79e+02 3.30e+02 1.94e+02 1.21e+02 8.21e+01 6.10e+01 4.88e+01 4.22e+01 3.89e+01 3.83e+01 3.81e+01
-zinc DRAFT 2.62e+06 5.85e+05 1.33e+05 3.13e+04 7.43e+03 1.87e+03 5.36e+02 2.03e+02 1.11e+02 8.26e+01 7.71e+01 7.46e+01
-Noise 2.93e+07 4.72e+06 7.76e+05 1.32e+05 2.34e+04 4.45e+03 9.48e+02 2.56e+02 1.14e+02 8.17e+01 7.69e+01 7.53e+01
-FGSM 2.50e+06 4.44e+05 9.49e+04 2.11e+04 4.89e+03 1.24e+03 3.66e+02 1.44e+02 9.01e+01 7.66e+01 7.53e+01 7.57e+01
-PGD 2.55e+06 4.38e+05 8.27e+04 1.87e+04 4.43e+03 1.13e+03 3.40e+02 1.39e+02 8.88e+01 7.66e+01 7.55e+01 7.60e+01
-SAWAR 1.31e+03 7.77e+02 4.72e+02 2.98e+02 1.98e+02 1.42e+02 1.10e+02 9.21e+01 8.28e+01 7.87e+01 7.80e+01 7.78e+01
+Dataset Algorithm
+Aids2 DRAFT 7.88e+05 2.75e+05 9.79e+04 3.51e+04 1.31e+04 5.06e+03 2.06e+03 9.59e+02 6.16e+02 5.45e+02 5.41e+02 5.41e+02
+Noise 3.15e+05 1.30e+05 5.45e+04 2.30e+04 9.53e+03 4.06e+03 1.79e+03 9.02e+02 6.06e+02 5.44e+02 5.41e+02 5.41e+02
+FGSM 1.29e+05 5.99e+04 2.79e+04 1.29e+04 5.92e+03 2.72e+03 1.32e+03 7.49e+02 5.69e+02 5.42e+02 5.41e+02 5.41e+02
+PGD 7.29e+04 3.62e+04 1.80e+04 8.83e+03 4.29e+03 2.10e+03 1.09e+03 6.82e+02 5.57e+02 5.42e+02 5.41e+02 5.41e+02
+AAE-Cox 6.04e+12 2.43e+11 1.30e+10 6.94e+08 4.08e+07 2.64e+06 1.76e+05 1.27e+04 1.18e+03 5.41e+02 5.38e+02 5.39e+02
+SAWAR 5.53e+02 5.48e+02 5.45e+02 5.43e+02 5.42e+02 5.42e+02 5.41e+02 5.41e+02 5.40e+02 5.40e+02 5.40e+02 5.40e+02
+Framingham DRAFT 2.24e+07 4.61e+06 9.70e+05 2.13e+05 4.98e+04 1.33e+04 4.62e+03 2.35e+03 1.71e+03 1.52e+03 1.49e+03 1.48e+03
+Noise 3.13e+07 6.71e+06 1.48e+06 3.35e+05 7.98e+04 2.05e+04 6.23e+03 2.67e+03 1.76e+03 1.53e+03 1.49e+03 1.48e+03
+FGSM 8.34e+09 7.44e+08 6.79e+07 6.46e+06 6.50e+05 7.31e+04 1.05e+04 2.77e+03 1.67e+03 1.50e+03 1.49e+03 1.48e+03
+PGD 1.53e+11 7.60e+09 3.88e+08 2.07e+07 1.20e+06 8.59e+04 9.43e+03 2.46e+03 1.61e+03 1.49e+03 1.48e+03 1.48e+03
+AAE-Cox 1.38e+26 1.22e+23 1.20e+20 1.10e+17 1.21e+14 1.26e+11 1.63e+08 3.73e+05 3.68e+03 1.49e+03 1.47e+03 1.47e+03
+SAWAR 1.67e+03 1.62e+03 1.58e+03 1.55e+03 1.53e+03 1.51e+03 1.50e+03 1.49e+03 1.48e+03 1.48e+03 1.48e+03 1.48e+03
+LeukSurv DRAFT 2.10e+08 3.80e+07 7.18e+06 1.29e+06 2.41e+05 4.59e+04 9.21e+03 2.13e+03 6.86e+02 3.75e+02 3.28e+02 3.07e+02
+Noise 1.40e+20 1.48e+18 1.45e+16 1.63e+14 1.29e+12 9.53e+09 8.92e+07 1.13e+06 2.90e+04 2.62e+03 1.25e+03 7.75e+02
+FGSM 6.24e+12 4.01e+11 2.30e+10 1.27e+09 7.06e+07 4.34e+06 2.73e+05 2.04e+04 2.18e+03 5.11e+02 3.72e+02 3.25e+02
+PGD 4.11e+11 3.30e+10 2.53e+09 1.99e+08 1.58e+07 1.30e+06 1.12e+05 1.10e+04 1.50e+03 4.34e+02 3.39e+02 3.07e+02
+AAE-Cox 4.24e+28 2.71e+25 1.61e+34 7.46e+29 2.45e+25 7.92e+19 7.26e+14 3.55e+09 1.15e+05 4.18e+02 2.99e+02 2.84e+02
+SAWAR 3.30e+03 2.00e+03 1.23e+03 7.95e+02 5.43e+02 3.97e+02 3.38e+02 3.06e+02 2.87e+02 2.75e+02 2.72e+02 2.70e+02
+TRACE DRAFT 5.09e+05 1.75e+05 6.15e+04 2.23e+04 8.34e+03 3.31e+03 1.48e+03 8.34e+02 6.09e+02 5.38e+02 5.27e+02 5.24e+02
+Noise 2.14e+09 2.12e+08 2.24e+07 2.58e+06 3.25e+05 4.62e+04 7.86e+03 1.89e+03 7.98e+02 5.76e+02 5.46e+02 5.35e+02
+FGSM 4.68e+08 4.01e+07 3.85e+06 4.31e+05 5.82e+04 9.91e+03 2.30e+03 8.91e+02 6.03e+02 5.39e+02 5.31e+02 5.28e+02
+PGD 4.42e+07 6.51e+06 1.04e+06 1.82e+05 3.49e+04 7.47e+03 2.01e+03 8.54e+02 5.99e+02 5.39e+02 5.31e+02 5.27e+02
+AAE-Cox 1.06e+24 1.53e+21 2.08e+18 2.78e+15 2.40e+12 2.92e+09 5.29e+06 2.86e+04 9.75e+02 5.41e+02 5.31e+02 5.29e+02
+SAWAR 1.49e+03 1.01e+03 7.74e+02 6.57e+02 5.98e+02 5.67e+02 5.48e+02 5.35e+02 5.28e+02 5.24e+02 5.24e+02 5.24e+02
+dataDIVAT1 DRAFT 2.97e+04 1.56e+04 8.36e+03 4.56e+03 2.59e+03 1.59e+03 1.11e+03 8.83e+02 7.91e+02 7.56e+02 7.51e+02 7.50e+02
+Noise 1.60e+10 1.12e+09 8.30e+07 6.62e+06 5.83e+05 5.99e+04 8.13e+03 1.87e+03 9.38e+02 7.78e+02 7.61e+02 7.59e+02
+FGSM 3.25e+08 3.59e+07 4.13e+06 5.10e+05 7.07e+04 1.17e+04 2.66e+03 1.10e+03 8.11e+02 7.58e+02 7.53e+02 7.52e+02
+PGD 1.22e+08 1.55e+07 2.06e+06 2.92e+05 4.56e+04 8.39e+03 2.15e+03 1.01e+03 7.94e+02 7.55e+02 7.52e+02 7.51e+02
+AAE-Cox 9.96e+18 6.79e+16 3.12e+14 1.88e+12 1.37e+10 1.14e+08 1.05e+06 1.57e+04 1.04e+03 7.50e+02 7.45e+02 7.45e+02
+SAWAR 7.89e+02 7.79e+02 7.70e+02 7.63e+02 7.58e+02 7.54e+02 7.51e+02 7.49e+02 7.47e+02 7.46e+02 7.46e+02 7.46e+02
+flchain DRAFT 1.57e+22 5.59e+19 2.03e+17 7.97e+14 3.37e+12 1.52e+10 7.49e+07 5.25e+05 1.45e+04 1.95e+03 1.25e+03 1.10e+03
+Noise 2.69e+32 2.90e+28 3.20e+24 5.94e+33 2.08e+27 7.50e+20 3.25e+14 2.31e+08 6.25e+04 2.40e+03 1.47e+03 1.24e+03
+FGSM 5.19e+19 7.59e+16 1.21e+14 2.16e+11 7.32e+08 9.31e+06 2.85e+05 1.58e+04 2.16e+03 1.18e+03 1.12e+03 1.10e+03
+PGD 3.10e+15 1.18e+13 6.11e+10 6.20e+08 1.63e+07 6.61e+05 3.81e+04 4.41e+03 1.43e+03 1.12e+03 1.09e+03 1.09e+03
+AAE-Cox nan nan nan nan nan nan nan nan 4.02e+33 2.23e+10 1.11e+03 1.08e+03
+SAWAR 5.84e+05 3.37e+04 4.04e+03 1.74e+03 1.31e+03 1.18e+03 1.13e+03 1.11e+03 1.10e+03 1.09e+03 1.09e+03 1.09e+03
+prostate DRAFT 5.92e+05 2.18e+05 8.11e+04 3.04e+04 1.16e+04 4.49e+03 1.83e+03 8.39e+02 4.86e+02 3.71e+02 3.48e+02 3.37e+02
+Noise 6.55e+23 1.63e+21 4.14e+18 1.02e+16 2.79e+13 8.22e+10 2.79e+08 1.36e+06 3.02e+04 3.29e+03 1.53e+03 9.01e+02
+FGSM 5.24e+15 9.63e+13 1.84e+12 3.69e+10 7.62e+08 1.69e+07 4.35e+05 1.56e+04 1.19e+03 4.22e+02 3.76e+02 3.63e+02
+PGD 1.20e+13 4.77e+11 1.96e+10 8.28e+08 3.66e+07 1.75e+06 9.45e+04 6.42e+03 8.39e+02 3.99e+02 3.67e+02 3.57e+02
+AAE-Cox 1.53e+20 9.61e+17 6.16e+15 4.00e+13 2.65e+11 1.87e+09 1.33e+07 1.16e+05 1.81e+03 3.47e+02 3.33e+02 3.31e+02
+SAWAR 6.21e+02 5.06e+02 4.28e+02 3.91e+02 3.67e+02 3.55e+02 3.47e+02 3.41e+02 3.37e+02 3.35e+02 3.34e+02 3.33e+02
+retinopathy DRAFT 1.21e+04 6.19e+03 3.21e+03 1.68e+03 9.15e+02 5.24e+02 3.31e+02 2.39e+02 1.96e+02 1.76e+02 1.71e+02 1.69e+02
+Noise 1.65e+04 8.10e+03 4.01e+03 2.03e+03 1.05e+03 5.83e+02 3.54e+02 2.47e+02 1.99e+02 1.77e+02 1.72e+02 1.69e+02
+FGSM 7.62e+03 4.03e+03 2.15e+03 1.18e+03 6.72e+02 4.10e+02 2.80e+02 2.16e+02 1.86e+02 1.73e+02 1.70e+02 1.69e+02
+PGD 7.32e+03 3.87e+03 2.08e+03 1.14e+03 6.55e+02 4.03e+02 2.75e+02 2.14e+02 1.85e+02 1.73e+02 1.70e+02 1.69e+02
+AAE-Cox 1.32e+13 3.68e+11 1.03e+10 2.98e+08 8.67e+06 2.82e+05 1.17e+04 8.14e+02 1.98e+02 1.68e+02 1.67e+02 1.67e+02
+SAWAR 6.10e+02 4.69e+02 3.69e+02 3.00e+02 2.51e+02 2.18e+02 1.96e+02 1.82e+02 1.74e+02 1.69e+02 1.68e+02 1.68e+02
+stagec DRAFT 1.17e+04 5.05e+03 2.20e+03 9.79e+02 4.44e+02 2.10e+02 1.07e+02 6.47e+01 4.80e+01 4.24e+01 4.15e+01 4.13e+01
+Noise 6.21e+04 2.17e+04 7.71e+03 2.72e+03 1.01e+03 3.86e+02 1.63e+02 8.11e+01 5.29e+01 4.44e+01 4.31e+01 4.27e+01
+FGSM 1.02e+04 4.31e+03 1.80e+03 7.90e+02 3.54e+02 1.69e+02 9.18e+01 5.99e+01 4.77e+01 4.40e+01 4.35e+01 4.36e+01
+PGD 8.55e+03 3.69e+03 1.60e+03 7.18e+02 3.32e+02 1.62e+02 8.89e+01 5.88e+01 4.73e+01 4.38e+01 4.34e+01 4.35e+01
+AAE-Cox 1.32e+20 6.14e+17 2.58e+15 1.21e+13 5.37e+10 2.81e+08 1.53e+06 1.01e+04 1.53e+02 4.20e+01 4.10e+01 4.09e+01
+SAWAR 2.52e+02 1.75e+02 1.26e+02 9.40e+01 7.32e+01 5.96e+01 5.14e+01 4.61e+01 4.28e+01 4.09e+01 4.04e+01 4.01e+01
+zinc DRAFT 1.34e+06 3.43e+05 8.95e+04 2.36e+04 6.49e+03 1.84e+03 5.72e+02 2.12e+02 1.11e+02 8.33e+01 7.84e+01 7.64e+01
+Noise 1.01e+07 2.02e+06 4.01e+05 8.52e+04 1.86e+04 4.26e+03 1.05e+03 3.06e+02 1.28e+02 8.63e+01 8.05e+01 7.86e+01
+FGSM 3.49e+05 9.39e+04 2.69e+04 7.98e+03 2.42e+03 8.00e+02 3.01e+02 1.41e+02 9.24e+01 7.95e+01 7.82e+01 7.82e+01
+PGD 4.28e+05 1.09e+05 2.89e+04 7.84e+03 2.37e+03 7.72e+02 2.87e+02 1.36e+02 9.09e+01 7.94e+01 7.83e+01 7.85e+01
+AAE-Cox 1.15e+24 1.54e+21 4.97e+18 8.90e+15 1.88e+13 2.84e+10 5.64e+07 1.63e+05 7.39e+02 8.21e+01 7.79e+01 7.76e+01
+SAWAR 5.06e+02 3.35e+02 2.31e+02 1.67e+02 1.28e+02 1.04e+02 9.02e+01 8.23e+01 7.82e+01 7.68e+01 7.68e+01 7.72e+01
 TABLE VIII: Negative Log Likelihood metric for SurvSet datasets (lower is better) for each adversarial training method against the worst-case adversarial attack.
+We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1 for the FGSM adversarial attack, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller.
+The relative percent changes in Concordance Index metric from the DRAFT training objective to SAWAR training objective is shown in Table IX (where higher percentage change is better).
+We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
+$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
+%$\%\Delta$ 39.82 59.68 72.41 86.66 105.06 136.98 168.31 178.41 116.94 26.54 12.48 1.6
+TABLE IX: The relative percent change in Concordance Index metric from the DRAFT model to the SAWAR training objective averaged across the SurvSet datasets for the FGSM adversarial attack.
+A lower relative percent change is better.
+$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
+Dataset Algorithm
+Aids2 DRAFT 0.232 0.233 0.235 0.238 0.243 0.25 0.259 0.274 0.3 0.376 0.458 0.57
+Noise 0.23 0.231 0.233 0.236 0.24 0.249 0.26 0.277 0.306 0.386 0.465 0.569
+FGSM 0.356 0.363 0.375 0.389 0.403 0.421 0.439 0.459 0.477 0.508 0.531 0.566
+PGD 0.326 0.336 0.35 0.367 0.385 0.409 0.434 0.46 0.486 0.515 0.534 0.567
+AAE-Cox 0.241 0.246 0.253 0.261 0.271 0.282 0.299 0.332 0.39 0.474 0.523 0.573
+SAWAR 0.259 0.265 0.276 0.289 0.306 0.327 0.352 0.385 0.43 0.493 0.53 0.569
+Framingham DRAFT 0.142 0.143 0.143 0.144 0.144 0.145 0.148 0.168 0.257 0.468 0.6 0.722
+Noise 0.143 0.144 0.144 0.145 0.145 0.146 0.15 0.173 0.265 0.473 0.601 0.719
+FGSM 0.149 0.152 0.156 0.164 0.18 0.208 0.257 0.332 0.437 0.57 0.643 0.715
+PGD 0.181 0.189 0.2 0.218 0.245 0.285 0.34 0.411 0.5 0.603 0.66 0.717
+AAE-Cox 0.148 0.152 0.162 0.181 0.216 0.269 0.34 0.427 0.53 0.636 0.687 0.733
+SAWAR 0.149 0.158 0.178 0.213 0.268 0.34 0.417 0.5 0.584 0.665 0.702 0.737
+LeukSurv DRAFT 0.378 0.38 0.381 0.383 0.387 0.394 0.407 0.432 0.475 0.544 0.587 0.633
+Noise 0.39 0.393 0.397 0.401 0.406 0.411 0.417 0.428 0.453 0.507 0.556 0.624
+FGSM 0.397 0.401 0.405 0.41 0.418 0.427 0.443 0.466 0.499 0.554 0.59 0.635
+PGD 0.399 0.402 0.407 0.411 0.419 0.431 0.447 0.473 0.508 0.562 0.597 0.638
+AAE-Cox 0.376 0.377 0.379 0.382 0.388 0.4 0.42 0.453 0.506 0.575 0.615 0.656
+SAWAR 0.392 0.399 0.409 0.423 0.442 0.465 0.494 0.529 0.574 0.623 0.649 0.673
+TRACE DRAFT 0.219 0.221 0.225 0.232 0.247 0.279 0.334 0.415 0.524 0.642 0.696 0.745
+Noise 0.234 0.236 0.24 0.246 0.259 0.286 0.336 0.414 0.52 0.639 0.695 0.745
+FGSM 0.29 0.311 0.336 0.369 0.408 0.454 0.507 0.565 0.625 0.685 0.715 0.744
+PGD 0.304 0.326 0.353 0.388 0.428 0.475 0.525 0.58 0.635 0.691 0.718 0.745
+AAE-Cox 0.251 0.273 0.302 0.339 0.384 0.438 0.498 0.563 0.63 0.692 0.721 0.747
+SAWAR 0.287 0.32 0.363 0.411 0.464 0.519 0.57 0.62 0.665 0.708 0.729 0.747
+dataDIVAT1 DRAFT 0.097 0.098 0.099 0.1 0.101 0.103 0.105 0.122 0.214 0.414 0.53 0.655
+Noise 0.094 0.094 0.095 0.096 0.097 0.099 0.101 0.111 0.178 0.366 0.494 0.635
+FGSM 0.114 0.116 0.12 0.126 0.138 0.161 0.203 0.27 0.362 0.482 0.559 0.646
+PGD 0.137 0.143 0.152 0.167 0.188 0.22 0.265 0.33 0.409 0.511 0.574 0.648
+AAE-Cox 0.111 0.114 0.122 0.139 0.167 0.209 0.268 0.347 0.443 0.551 0.607 0.663
+SAWAR 0.156 0.201 0.253 0.311 0.368 0.422 0.461 0.506 0.555 0.609 0.637 0.67
+flchain DRAFT 0.152 0.153 0.154 0.155 0.156 0.157 0.159 0.164 0.197 0.895 0.903 0.921
+Noise 0.166 0.167 0.169 0.171 0.175 0.182 0.213 0.333 0.809 0.9 0.904 0.918
+FGSM 0.507 0.628 0.727 0.8 0.854 0.883 0.898 0.902 0.906 0.911 0.915 0.922
+PGD 0.539 0.717 0.824 0.876 0.896 0.9 0.903 0.905 0.908 0.913 0.917 0.923
+AAE-Cox 0.172 0.188 0.218 0.275 0.372 0.508 0.665 0.81 0.899 0.918 0.922 0.926
+SAWAR 0.628 0.831 0.894 0.9 0.904 0.907 0.909 0.912 0.916 0.921 0.924 0.927
+prostate DRAFT 0.308 0.311 0.314 0.318 0.321 0.326 0.333 0.339 0.357 0.435 0.531 0.668
+Noise 0.305 0.305 0.306 0.307 0.309 0.314 0.32 0.328 0.344 0.394 0.457 0.562
+FGSM 0.293 0.295 0.297 0.3 0.304 0.308 0.312 0.32 0.343 0.407 0.474 0.571
+PGD 0.299 0.301 0.303 0.305 0.308 0.312 0.318 0.331 0.36 0.431 0.502 0.588
+AAE-Cox 0.308 0.308 0.311 0.315 0.325 0.342 0.375 0.426 0.499 0.597 0.645 0.691
+SAWAR 0.288 0.292 0.299 0.31 0.326 0.349 0.384 0.432 0.499 0.579 0.618 0.657
+retinopathy DRAFT 0.139 0.141 0.145 0.15 0.152 0.152 0.153 0.158 0.209 0.425 0.554 0.669
+Noise 0.134 0.137 0.138 0.138 0.138 0.138 0.138 0.15 0.245 0.456 0.57 0.668
+FGSM 0.131 0.133 0.135 0.135 0.135 0.135 0.137 0.155 0.255 0.456 0.561 0.659
+PGD 0.13 0.13 0.131 0.131 0.131 0.131 0.134 0.154 0.254 0.456 0.56 0.657
+AAE-Cox 0.129 0.129 0.13 0.137 0.157 0.196 0.247 0.337 0.444 0.541 0.595 0.648
+SAWAR 0.131 0.13 0.13 0.131 0.135 0.149 0.178 0.251 0.371 0.51 0.584 0.647
+stagec DRAFT 0.137 0.136 0.136 0.132 0.136 0.135 0.151 0.174 0.23 0.34 0.407 0.512
+Noise 0.133 0.133 0.132 0.132 0.136 0.144 0.16 0.187 0.261 0.362 0.442 0.555
+FGSM 0.12 0.124 0.127 0.134 0.146 0.147 0.157 0.184 0.255 0.33 0.417 0.51
+PGD 0.115 0.117 0.122 0.128 0.141 0.144 0.154 0.181 0.259 0.327 0.412 0.505
+AAE-Cox 0.125 0.121 0.131 0.141 0.163 0.202 0.241 0.296 0.352 0.426 0.485 0.54
+SAWAR 0.123 0.124 0.127 0.139 0.149 0.179 0.239 0.281 0.347 0.425 0.486 0.558
+zinc DRAFT 0.057 0.057 0.057 0.057 0.057 0.057 0.057 0.068 0.155 0.485 0.646 0.77
+Noise 0.053 0.053 0.054 0.053 0.054 0.057 0.062 0.095 0.248 0.547 0.668 0.776
+FGSM 0.055 0.056 0.056 0.058 0.06 0.07 0.095 0.199 0.404 0.61 0.705 0.783
+PGD 0.056 0.056 0.057 0.058 0.064 0.078 0.113 0.22 0.42 0.615 0.706 0.781
+AAE-Cox 0.062 0.065 0.079 0.101 0.162 0.227 0.291 0.409 0.551 0.655 0.708 0.759
+SAWAR 0.055 0.056 0.058 0.073 0.111 0.213 0.325 0.446 0.556 0.679 0.729 0.774
+TABLE X: Concordance Index metric for SurvSet datasets (higher is better) for each adversarial training method against the FGSM adversarial attack.
+We find that as the $\epsilon$-perturbation magnitude increases from 0 to 1 for the FGSM adversarial attack, the relative percentage change from DRAFT to the adversarial training methods becomes larger and then smaller.
+The relative percent changes in Integrated Brier Scores metric from the DRAFT training objective to SAWAR training objective is shown in Table XI (where lower percentage change is better).
+We note that for very large $\epsilon$, since our data is standard normalized all methods begin to fail.
+$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
+%$\%\Delta$44.57 -46.37 -47.93 -49.06 -49.43 -48.69 -46.35 -41.82 -33.56 -20.63 -11.37 -0.65
+TABLE XI: The relative percent change in Integrated Brier Score metric from the DRAFT model to the SAWAR training objective averaged across the SurvSet datasets for the FGSM adversarial attack.
+A lower relative percent change is better.
+$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
+Dataset Algorithm
+Aids2 DRAFT 0.232 0.233 0.235 0.238 0.243 0.25 0.259 0.274 0.3 0.376 0.458 0.57
+Noise 0.23 0.231 0.233 0.236 0.24 0.249 0.26 0.277 0.306 0.386 0.465 0.569
+FGSM 0.356 0.363 0.375 0.389 0.403 0.421 0.439 0.459 0.477 0.508 0.531 0.566
+PGD 0.326 0.336 0.35 0.367 0.385 0.409 0.434 0.46 0.486 0.515 0.534 0.567
+AAE-Cox 0.241 0.246 0.253 0.261 0.271 0.282 0.299 0.332 0.39 0.474 0.523 0.573
+SAWAR 0.259 0.265 0.276 0.289 0.306 0.327 0.352 0.385 0.43 0.493 0.53 0.569
+Framingham DRAFT 0.142 0.143 0.143 0.144 0.144 0.145 0.148 0.168 0.257 0.468 0.6 0.722
+Noise 0.143 0.144 0.144 0.145 0.145 0.146 0.15 0.173 0.265 0.473 0.601 0.719
+FGSM 0.149 0.152 0.156 0.164 0.18 0.208 0.257 0.332 0.437 0.57 0.643 0.715
+PGD 0.181 0.189 0.2 0.218 0.245 0.285 0.34 0.411 0.5 0.603 0.66 0.717
+AAE-Cox 0.148 0.152 0.162 0.181 0.216 0.269 0.34 0.427 0.53 0.636 0.687 0.733
+SAWAR 0.149 0.158 0.178 0.213 0.268 0.34 0.417 0.5 0.584 0.665 0.702 0.737
+LeukSurv DRAFT 0.378 0.38 0.381 0.383 0.387 0.394 0.407 0.432 0.475 0.544 0.587 0.633
+Noise 0.39 0.393 0.397 0.401 0.406 0.411 0.417 0.428 0.453 0.507 0.556 0.624
+FGSM 0.397 0.401 0.405 0.41 0.418 0.427 0.443 0.466 0.499 0.554 0.59 0.635
+PGD 0.399 0.402 0.407 0.411 0.419 0.431 0.447 0.473 0.508 0.562 0.597 0.638
+AAE-Cox 0.376 0.377 0.379 0.382 0.388 0.4 0.42 0.453 0.506 0.575 0.615 0.656
+SAWAR 0.392 0.399 0.409 0.423 0.442 0.465 0.494 0.529 0.574 0.623 0.649 0.673
+TRACE DRAFT 0.219 0.221 0.225 0.232 0.247 0.279 0.334 0.415 0.524 0.642 0.696 0.745
+Noise 0.234 0.236 0.24 0.246 0.259 0.286 0.336 0.414 0.52 0.639 0.695 0.745
+FGSM 0.29 0.311 0.336 0.369 0.408 0.454 0.507 0.565 0.625 0.685 0.715 0.744
+PGD 0.304 0.326 0.353 0.388 0.428 0.475 0.525 0.58 0.635 0.691 0.718 0.745
+AAE-Cox 0.251 0.273 0.302 0.339 0.384 0.438 0.498 0.563 0.63 0.692 0.721 0.747
+SAWAR 0.287 0.32 0.363 0.411 0.464 0.519 0.57 0.62 0.665 0.708 0.729 0.747
+dataDIVAT1 DRAFT 0.097 0.098 0.099 0.1 0.101 0.103 0.105 0.122 0.214 0.414 0.53 0.655
+Noise 0.094 0.094 0.095 0.096 0.097 0.099 0.101 0.111 0.178 0.366 0.494 0.635
+FGSM 0.114 0.116 0.12 0.126 0.138 0.161 0.203 0.27 0.362 0.482 0.559 0.646
+PGD 0.137 0.143 0.152 0.167 0.188 0.22 0.265 0.33 0.409 0.511 0.574 0.648
+AAE-Cox 0.111 0.114 0.122 0.139 0.167 0.209 0.268 0.347 0.443 0.551 0.607 0.663
+SAWAR 0.156 0.201 0.253 0.311 0.368 0.422 0.461 0.506 0.555 0.609 0.637 0.67
+flchain DRAFT 0.152 0.153 0.154 0.155 0.156 0.157 0.159 0.164 0.197 0.895 0.903 0.921
+Noise 0.166 0.167 0.169 0.171 0.175 0.182 0.213 0.333 0.809 0.9 0.904 0.918
+FGSM 0.507 0.628 0.727 0.8 0.854 0.883 0.898 0.902 0.906 0.911 0.915 0.922
+PGD 0.539 0.717 0.824 0.876 0.896 0.9 0.903 0.905 0.908 0.913 0.917 0.923
+AAE-Cox 0.172 0.188 0.218 0.275 0.372 0.508 0.665 0.81 0.899 0.918 0.922 0.926
+SAWAR 0.628 0.831 0.894 0.9 0.904 0.907 0.909 0.912 0.916 0.921 0.924 0.927
+prostate DRAFT 0.308 0.311 0.314 0.318 0.321 0.326 0.333 0.339 0.357 0.435 0.531 0.668
+Noise 0.305 0.305 0.306 0.307 0.309 0.314 0.32 0.328 0.344 0.394 0.457 0.562
+FGSM 0.293 0.295 0.297 0.3 0.304 0.308 0.312 0.32 0.343 0.407 0.474 0.571
+PGD 0.299 0.301 0.303 0.305 0.308 0.312 0.318 0.331 0.36 0.431 0.502 0.588
+AAE-Cox 0.308 0.308 0.311 0.315 0.325 0.342 0.375 0.426 0.499 0.597 0.645 0.691
+SAWAR 0.288 0.292 0.299 0.31 0.326 0.349 0.384 0.432 0.499 0.579 0.618 0.657
+retinopathy DRAFT 0.139 0.141 0.145 0.15 0.152 0.152 0.153 0.158 0.209 0.425 0.554 0.669
+Noise 0.134 0.137 0.138 0.138 0.138 0.138 0.138 0.15 0.245 0.456 0.57 0.668
+FGSM 0.131 0.133 0.135 0.135 0.135 0.135 0.137 0.155 0.255 0.456 0.561 0.659
+PGD 0.13 0.13 0.131 0.131 0.131 0.131 0.134 0.154 0.254 0.456 0.56 0.657
+AAE-Cox 0.129 0.129 0.13 0.137 0.157 0.196 0.247 0.337 0.444 0.541 0.595 0.648
+SAWAR 0.131 0.13 0.13 0.131 0.135 0.149 0.178 0.251 0.371 0.51 0.584 0.647
+stagec DRAFT 0.137 0.136 0.136 0.132 0.136 0.135 0.151 0.174 0.23 0.34 0.407 0.512
+Noise 0.133 0.133 0.132 0.132 0.136 0.144 0.16 0.187 0.261 0.362 0.442 0.555
+FGSM 0.12 0.124 0.127 0.134 0.146 0.147 0.157 0.184 0.255 0.33 0.417 0.51
+PGD 0.115 0.117 0.122 0.128 0.141 0.144 0.154 0.181 0.259 0.327 0.412 0.505
+AAE-Cox 0.125 0.121 0.131 0.141 0.163 0.202 0.241 0.296 0.352 0.426 0.485 0.54
+SAWAR 0.123 0.124 0.127 0.139 0.149 0.179 0.239 0.281 0.347 0.425 0.486 0.558
+zinc DRAFT 0.057 0.057 0.057 0.057 0.057 0.057 0.057 0.068 0.155 0.485 0.646 0.77
+Noise 0.053 0.053 0.054 0.053 0.054 0.057 0.062 0.095 0.248 0.547 0.668 0.776
+FGSM 0.055 0.056 0.056 0.058 0.06 0.07 0.095 0.199 0.404 0.61 0.705 0.783
+PGD 0.056 0.056 0.057 0.058 0.064 0.078 0.113 0.22 0.42 0.615 0.706 0.781
+AAE-Cox 0.062 0.065 0.079 0.101 0.162 0.227 0.291 0.409 0.551 0.655 0.708 0.759
+SAWAR 0.055 0.056 0.058 0.073 0.111 0.213 0.325 0.446 0.556 0.679 0.729 0.774
+TABLE XII: Concordance Index metric for SurvSet datasets (higher is better) for each adversarial training method against the FGSM adversarial attack.
+$\epsilon$ 1.00 0.90 0.80 0.70 0.60 0.50 0.40 0.30 0.20 0.10 0.05 0.00
+Dataset Algorithm
+Aids2 DRAFT 1.04e+03 9.95e+02 9.48e+02 9.01e+02 8.54e+02 8.06e+02 7.56e+02 7.04e+02 6.49e+02 5.92e+02 5.66e+02 5.41e+02
+Noise 9.94e+02 9.53e+02 9.11e+02 8.69e+02 8.26e+02 7.82e+02 7.37e+02 6.89e+02 6.39e+02 5.88e+02 5.64e+02 5.41e+02
+FGSM 6.16e+02 6.10e+02 6.04e+02 5.97e+02 5.91e+02 5.84e+02 5.77e+02 5.70e+02 5.63e+02 5.54e+02 5.48e+02 5.41e+02
+PGD 6.20e+02 6.13e+02 6.05e+02 5.98e+02 5.91e+02 5.83e+02 5.76e+02 5.68e+02 5.60e+02 5.52e+02 5.47e+02 5.41e+02
+AAE-Cox 7.00e+02 6.83e+02 6.67e+02 6.51e+02 6.35e+02 6.19e+02 6.03e+02 5.87e+02 5.71e+02 5.55e+02 5.47e+02 5.39e+02
+SAWAR 5.84e+02 5.80e+02 5.76e+02 5.72e+02 5.68e+02 5.63e+02 5.59e+02 5.55e+02 5.50e+02 5.45e+02 5.43e+02 5.40e+02
+Framingham DRAFT 4.23e+03 4.10e+03 3.91e+03 3.65e+03 3.33e+03 2.97e+03 2.61e+03 2.26e+03 1.95e+03 1.68e+03 1.57e+03 1.48e+03
+Noise 4.00e+03 3.89e+03 3.73e+03 3.51e+03 3.23e+03 2.91e+03 2.57e+03 2.24e+03 1.94e+03 1.69e+03 1.58e+03 1.48e+03
+FGSM 2.11e+03 2.07e+03 2.03e+03 1.98e+03 1.93e+03 1.88e+03 1.82e+03 1.75e+03 1.67e+03 1.58e+03 1.53e+03 1.48e+03
+PGD 1.99e+03 1.95e+03 1.91e+03 1.87e+03 1.83e+03 1.79e+03 1.74e+03 1.68e+03 1.62e+03 1.56e+03 1.52e+03 1.48e+03
+AAE-Cox 2.38e+03 2.27e+03 2.16e+03 2.06e+03 1.96e+03 1.87e+03 1.78e+03 1.70e+03 1.62e+03 1.54e+03 1.51e+03 1.47e+03
+SAWAR 1.96e+03 1.90e+03 1.84e+03 1.79e+03 1.73e+03 1.68e+03 1.64e+03 1.59e+03 1.55e+03 1.51e+03 1.49e+03 1.48e+03
+LeukSurv DRAFT 2.66e+03 2.16e+03 1.77e+03 1.45e+03 1.19e+03 9.73e+02 7.93e+02 6.41e+02 5.10e+02 3.99e+02 3.50e+02 3.07e+02
+Noise 1.13e+06 5.84e+05 2.96e+05 1.52e+05 7.88e+04 4.05e+04 1.94e+04 9.21e+03 4.25e+03 1.89e+03 1.23e+03 7.75e+02
+FGSM 2.76e+03 2.33e+03 1.96e+03 1.65e+03 1.39e+03 1.16e+03 9.64e+02 7.82e+02 6.18e+02 4.64e+02 3.92e+02 3.25e+02
+PGD 1.79e+03 1.57e+03 1.38e+03 1.21e+03 1.05e+03 9.10e+02 7.76e+02 6.49e+02 5.30e+02 4.15e+02 3.60e+02 3.07e+02
+AAE-Cox 1.09e+03 9.83e+02 8.87e+02 7.97e+02 7.12e+02 6.32e+02 5.57e+02 4.84e+02 4.12e+02 3.43e+02 3.12e+02 2.84e+02
+SAWAR 5.40e+02 5.12e+02 4.85e+02 4.57e+02 4.29e+02 4.01e+02 3.73e+02 3.46e+02 3.20e+02 2.94e+02 2.82e+02 2.70e+02
+TRACE DRAFT 2.31e+03 2.09e+03 1.87e+03 1.64e+03 1.43e+03 1.22e+03 1.03e+03 8.69e+02 7.29e+02 6.14e+02 5.66e+02 5.24e+02
+Noise 4.55e+03 3.85e+03 3.20e+03 2.63e+03 2.14e+03 1.72e+03 1.36e+03 1.06e+03 8.29e+02 6.60e+02 5.93e+02 5.35e+02
+FGSM 1.23e+03 1.15e+03 1.07e+03 9.98e+02 9.26e+02 8.54e+02 7.82e+02 7.13e+02 6.47e+02 5.85e+02 5.56e+02 5.28e+02
+PGD 1.17e+03 1.10e+03 1.03e+03 9.59e+02 8.90e+02 8.21e+02 7.54e+02 6.93e+02 6.35e+02 5.80e+02 5.53e+02 5.27e+02
+AAE-Cox 1.07e+03 1.00e+03 9.38e+02 8.77e+02 8.19e+02 7.63e+02 7.09e+02 6.59e+02 6.11e+02 5.68e+02 5.48e+02 5.29e+02
+SAWAR 1.01e+03 9.32e+02 8.63e+02 8.01e+02 7.47e+02 6.99e+02 6.57e+02 6.19e+02 5.84e+02 5.53e+02 5.38e+02 5.24e+02
+dataDIVAT1 DRAFT 1.78e+03 1.72e+03 1.64e+03 1.55e+03 1.44e+03 1.32e+03 1.20e+03 1.07e+03 9.55e+02 8.46e+02 7.96e+02 7.50e+02
+Noise 2.02e+03 1.96e+03 1.88e+03 1.78e+03 1.65e+03 1.50e+03 1.35e+03 1.19e+03 1.03e+03 8.87e+02 8.20e+02 7.59e+02
+FGSM 1.09e+03 1.07e+03 1.05e+03 1.02e+03 1.00e+03 9.72e+02 9.40e+02 9.03e+02 8.59e+02 8.09e+02 7.81e+02 7.52e+02
+PGD 1.01e+03 9.95e+02 9.78e+02 9.60e+02 9.41e+02 9.19e+02 8.94e+02 8.65e+02 8.32e+02 7.95e+02 7.74e+02 7.51e+02
+AAE-Cox 1.25e+03 1.18e+03 1.12e+03 1.06e+03 1.01e+03 9.55e+02 9.09e+02 8.65e+02 8.23e+02 7.83e+02 7.64e+02 7.45e+02
+SAWAR 9.24e+02 9.01e+02 8.79e+02 8.59e+02 8.39e+02 8.20e+02 8.04e+02 7.89e+02 7.74e+02 7.60e+02 7.53e+02 7.46e+02
+flchain DRAFT 2.56e+04 2.36e+04 2.15e+04 1.95e+04 1.73e+04 1.47e+04 1.11e+04 6.38e+03 2.91e+03 1.59e+03 1.27e+03 1.10e+03
+Noise 1.62e+05 1.27e+05 8.24e+04 4.85e+04 2.97e+04 1.81e+04 9.84e+03 5.14e+03 2.95e+03 1.80e+03 1.45e+03 1.24e+03
+FGSM 2.49e+03 2.17e+03 1.87e+03 1.62e+03 1.44e+03 1.34e+03 1.28e+03 1.23e+03 1.19e+03 1.14e+03 1.12e+03 1.10e+03
+PGD 2.07e+03 1.76e+03 1.53e+03 1.37e+03 1.28e+03 1.22e+03 1.19e+03 1.16e+03 1.14e+03 1.11e+03 1.10e+03 1.09e+03
+AAE-Cox 2.98e+03 2.85e+03 2.73e+03 2.62e+03 2.50e+03 2.38e+03 2.23e+03 1.97e+03 1.47e+03 1.12e+03 1.09e+03 1.08e+03
+SAWAR 2.18e+03 1.83e+03 1.55e+03 1.33e+03 1.20e+03 1.14e+03 1.13e+03 1.11e+03 1.10e+03 1.09e+03 1.09e+03 1.09e+03
+prostate DRAFT 1.50e+03 1.44e+03 1.36e+03 1.24e+03 1.10e+03 9.38e+02 7.70e+02 6.16e+02 4.90e+02 3.97e+02 3.63e+02 3.37e+02
+Noise 1.88e+06 1.21e+06 6.72e+05 3.64e+05 1.97e+05 9.86e+04 4.32e+04 1.69e+04 6.34e+03 2.38e+03 1.46e+03 9.01e+02
+FGSM 8.51e+02 8.48e+02 8.41e+02 8.25e+02 7.97e+02 7.47e+02 6.78e+02 5.92e+02 5.05e+02 4.25e+02 3.92e+02 3.63e+02
+PGD 7.46e+02 7.32e+02 7.17e+02 6.97e+02 6.69e+02 6.30e+02 5.79e+02 5.22e+02 4.63e+02 4.05e+02 3.79e+02 3.57e+02
+AAE-Cox 4.64e+02 4.47e+02 4.31e+02 4.16e+02 4.02e+02 3.89e+02 3.77e+02 3.64e+02 3.52e+02 3.41e+02 3.36e+02 3.31e+02
+SAWAR 4.40e+02 4.28e+02 4.15e+02 4.04e+02 3.92e+02 3.81e+02 3.70e+02 3.60e+02 3.51e+02 3.42e+02 3.37e+02 3.33e+02
+retinopathy DRAFT 5.04e+02 4.93e+02 4.70e+02 4.35e+02 3.93e+02 3.48e+02 3.02e+02 2.59e+02 2.22e+02 1.92e+02 1.80e+02 1.69e+02
+Noise 5.47e+02 5.29e+02 4.96e+02 4.54e+02 4.04e+02 3.54e+02 3.05e+02 2.62e+02 2.24e+02 1.94e+02 1.81e+02 1.69e+02
+FGSM 4.16e+02 4.09e+02 3.93e+02 3.69e+02 3.39e+02 3.06e+02 2.73e+02 2.41e+02 2.13e+02 1.89e+02 1.79e+02 1.69e+02
+PGD 4.11e+02 4.03e+02 3.88e+02 3.65e+02 3.37e+02 3.05e+02 2.71e+02 2.40e+02 2.12e+02 1.89e+02 1.78e+02 1.69e+02
+AAE-Cox 2.27e+02 2.21e+02 2.14e+02 2.09e+02 2.03e+02 1.97e+02 1.91e+02 1.85e+02 1.79e+02 1.73e+02 1.70e+02 1.67e+02
+SAWAR 3.27e+02 3.14e+02 2.99e+02 2.82e+02 2.65e+02 2.47e+02 2.29e+02 2.12e+02 1.96e+02 1.81e+02 1.74e+02 1.68e+02
+stagec DRAFT 1.38e+02 1.32e+02 1.24e+02 1.15e+02 1.04e+02 9.29e+01 8.09e+01 6.91e+01 5.82e+01 4.90e+01 4.49e+01 4.13e+01
+Noise 1.52e+02 1.44e+02 1.35e+02 1.24e+02 1.12e+02 9.97e+01 8.67e+01 7.39e+01 6.18e+01 5.13e+01 4.68e+01 4.27e+01
+FGSM 1.04e+02 1.01e+02 9.63e+01 9.11e+01 8.52e+01 7.85e+01 7.12e+01 6.37e+01 5.64e+01 4.96e+01 4.65e+01 4.36e+01
+PGD 1.01e+02 9.73e+01 9.31e+01 8.82e+01 8.27e+01 7.65e+01 6.97e+01 6.27e+01 5.58e+01 4.93e+01 4.63e+01 4.35e+01
+AAE-Cox 8.00e+01 7.47e+01 6.99e+01 6.55e+01 6.15e+01 5.77e+01 5.42e+01 5.07e+01 4.74e+01 4.41e+01 4.25e+01 4.09e+01
+SAWAR 9.46e+01 8.98e+01 8.41e+01 7.80e+01 7.17e+01 6.57e+01 6.03e+01 5.48e+01 4.95e+01 4.45e+01 4.23e+01 4.01e+01
+zinc DRAFT 5.01e+02 4.94e+02 4.77e+02 4.45e+02 3.93e+02 3.19e+02 2.39e+02 1.73e+02 1.27e+02 9.62e+01 8.52e+01 7.64e+01
+Noise 5.54e+02 5.34e+02 5.00e+02 4.52e+02 3.87e+02 3.10e+02 2.35e+02 1.75e+02 1.32e+02 1.01e+02 8.85e+01 7.86e+01
+FGSM 2.89e+02 2.79e+02 2.63e+02 2.39e+02 2.10e+02 1.79e+02 1.52e+02 1.28e+02 1.09e+02 9.21e+01 8.48e+01 7.82e+01
+PGD 2.77e+02 2.67e+02 2.51e+02 2.29e+02 2.02e+02 1.74e+02 1.48e+02 1.26e+02 1.07e+02 9.18e+01 8.48e+01 7.85e+01
+AAE-Cox 1.66e+02 1.57e+02 1.48e+02 1.38e+02 1.27e+02 1.17e+02 1.08e+02 9.94e+01 9.16e+01 8.44e+01 8.10e+01 7.76e+01
+SAWAR 2.11e+02 1.92e+02 1.73e+02 1.57e+02 1.42e+02 1.28e+02 1.15e+02 1.04e+02 9.42e+01 8.52e+01 8.11e+01 7.72e+01
+TABLE XIII: Negative Log Likelihood metric for SurvSet datasets (lower is better) for each adversarial training method against the FGSM adversarial attack.